### v1

# assumeAssumption

Abstract
What does it mean to be flat? We propose to define it by measuring the maximal variation around a point, or from a dual perspective, the distance to neighboring level sets. After developing some calculus rules, we show how flat minima, conservation laws, and symmetries are intertwined. Gradient flows of conserved quantities are of particular interest, due to their flattening properties. Keywords: dynamical systems, semi-algebraic geometry, variational analysis.
MSC 2020: 14P10, 34A60, 49-XX.
1 Introduction
Flat minima were informally introduced by Hochreiter and Schmidhuber [19, 20] in the context of deep learning as connected regions of minima where the objective is nearly constant. In a desire to formalize this idea, it was later suggested to use the volume of connected components of sublevel sets [13, Definition 1]. Around the same time, an empirical measure of sharpness [26, Metric 2.1] was proposed to analyze the role of batch size in the stochastic subgradient method for training neural networks. For nonnegative functions, it was redefined as the maximal ratio of the function variation over one plus the function value, over a ball of fixed radius [13, Definition 2].
More recently, flat minima of deep matrix factorization were defined as global minima which minimize the trace [16] or the maximum eigenvalue [33, 29, 32] of the Hessian of the objective function. A scaled trace of the Hessian tailored for matrix factorization [12] was also proposed. Others still use the trace of the Hessian combined with gradient dynamics for $C^{4}$ functions [1, Definition 3].
Evidently, there is no commonly agreed upon definition of flatness. Yet, it has been reported that, when training deep neural networks, algorithms tend to find flat minima [26, 42]. These may have good generalization properties [20, 2]. Due to their possible role in deep learning theory, it seems enviable to have a definition that is not problem specific and somehow captures the previous ones. It should also fill in the gaps left by prior work.
As it stands, being a flat global minimum of an overparametrized ReLU neural network has no meaning. To make things concrete, consider
$f(x)=(x_{2}\mathrm{ReLU}(x_{1})+x_{3}-1)^{2}$
where $\mathrm{ReLU}(t)=\max\{0,t\}$. The objective function is differentiable at its global minima, so its gradient is defined and equal to zero (details in Example˜6). But it is not twice differentiable there, so its Hessian is not defined. As for the volumes of sublevel sets, they are generally not finite on ReLU networks [13, Theorem 2], nor on the most simple neural network, i.e., $f(x)=(x_{1}x_{2}-1)^{2}$.
Even when the Hessian is available, flatness is currently ill-defined. Minimizing the maximal eigenvalue $\lambda_{1}(\nabla^{2}f(x))$ over the solution set arbitrarily discards higher-order variation. This is problematic. For example,
$f(x)=x_{2}^{2}+x_{1}^{2}x_{2}^{4}$
obeys $\nabla f(x_{1},0)=(0,0)$ and $\lambda_{1}(\nabla^{2}f(x_{1},0))=2$ for all $x_{1}\in\mathbb{R}$, according to which all the global minima $(x_{1},0)$ are allegedly flat. But factoring in the fourth-order growth actually suggests that $(0,0)$ is the unique flat minimum (see Example˜8). This is not merely a theoretical matter, as gradient descent with sufficiently slowly diminishing step lengths converges to the origin. This can be seen in Fig.˜1, and is proved in the sequel [22].
Figure 1: 1000 iterations of gradient descent applied to $f(x_{1},x_{2})=x_{2}^{2}+x_{1}^{2}x_{2}^{4}$ with step length $(k+1)^{-1/6}$ initialized at $(3.2,0.6)$.
Leaving algorithmic aspects aside, in this work we propose to define flatness by measuring the maximal variation around a point, or equivalently, for a large function class [^1] The class of locally Lipschitz definable functions with nowhere dense level sets. If one is solely interested in the flatness of global minima, it suffices for the set of global minima to be nowhere dense., the distance to neighboring level sets. It is inspired by topographic maps used in mountain hiking: concentrated contour lines signal significant elevation change, while diffuse ones suggest a flat region like a valley. We actually define a preorder on $\mathbb{R}^{n}$, which is a total preorder if the objective is definable in an o-minimal structure on the real field [40]. This gives a precise meaning to the adjectives flatter and sharper.
At a glance, our results are as follows. We first cover basic aspects: the definition of flatness, the properties of maximizing curves, and several calculus rules. The rules rely on the subgradient, gradient, Hessian, and possibly higher-order derivatives. Notably, for a local minimum $\overline{x}$ of a smooth function $f$, $\overline{x}$ is flat only if it is a local minimum of $\lambda_{1}(\nabla^{2}f(x))$ subject to $f(x)=f(\overline{x})$. The converse holds if it is a strict local minimum of the constrained problem.
Second, we show how conserved quantities $c(x)$ in subgradient dynamics
$\dot{x}\in-\overline{\partial}f(x)$
provide a useful tool for analyzing flatness, a theme of recent interest [43]. On the one hand, they enable one to detect situations where level sets are expanding in a certain direction. In that case, gradient trajectories of the conserved quantity, modeled by
$\dot{x}=-\nabla c(x),$
flatten over time. In particular, if $x(0)$ is a local minimum of $f$ and $f$ is smooth, then
$\lambda_{1}(\nabla^{2}f(x(t)))\leq e^{-\omega t}\lambda_{1}(\nabla^{2}f(x(0)))$
for some constant $\omega>0$. On the other hand, quadratic conserved quantities $C(x)$ arising from linear symmetries of $f$ can help characterize flatness. For instance, if $x$ is a flat minimum of a smooth function $f$, then there exists a maximal eigenvector $v$ of $\nabla^{2}f(x)$ such that $C(v)=0$.
This clarifies the picture in matrix factorization
$\begin{array}[]{cccc}f:&\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}&\longrightarrow&\mathbb{R}\\ &(X,Y)&\longmapsto&\|XY-M\|_{F}^{2}\end{array}$
where $M\in\mathbb{R}^{m\times n}$, $\|\cdot\|_{F}$ is the Frobenius norm, and $C(X,Y)=X^{T}X=YY^{T}$. Some authors [12] report the flat minima are “balanced”, meaning that $X^{T}X=YY^{T}$, while others [33] say that they are “nearly balanced”. With our definition, we find that while $X^{T}X-YY^{T}$ can be nonzero, there does exist a balanced maximal eigenvector of the Hessian $\nabla^{2}f(X,Y)$. In passing, we note that flat minima admit a simple characterization. Namely, if $XY=M$, then
$(X,Y)~\text{is flat}~~~\Longleftrightarrow~~~\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}$
where $\|\cdot\|_{2}$ is the spectral norm. Our analysis also reveals a new local-global property of matrix factorization. While it is known the every local minimum is a global minimum [5, 38], it was not known that locally flat minima are globally flat.
Third, we determine flat minima in a series of examples that were out of reach with previous definitions and tools. They demonstrate that the definition of flatness proposed in this work is both versatile and workable. We also illustrate the connection with symmetry and conservation. In one example,
$f(x)=(x_{1}^{\upsilon_{1}}\cdots x_{n}^{\upsilon_{n}}-1)^{2}$
where $\upsilon_{1},\ldots,\upsilon_{n}$ are positive integers, any gradient trajectory of the conserved quantity initialized at a global minimum converges to a flat global minimum. It does so while remaining in the (n1)(n-1)-dimensional solution manifold.
The paper is organized as follows. Section˜2 contains background material on several branches of mathematics. Section˜3 treats basic aspects of flat minima. Section˜4 builds on them to forge a link with conservation laws and symmetries. Finally Section˜5 provides several examples of flat minima.
2 Background
This works relies on several branches of mathematics. We include a vignette of each one after introducing some standard notations. As usual,
$\mathbb{N}=\{0,1,\ldots\},~~~\mathbb{N}^{*}=\mathbb{N}\setminus\{0\},~~~\llbracket 1,k\rrbracket=\{1,\ldots,k\},$
$\overline{\mathbb{R}}=\mathbb{R}\cup\{\infty\},~~~\mathbb{R}_{-}=[0,\infty),~~~\mathbb{R}_{+}=[0,\infty).$
The sign of a real number $t$ is defined by
$\mathrm{sign}(t)=\left\{\begin{array}[]{cl}t/|t|&\text{if}~t\neq 0,\\ \left[-1,1\right]&\text{else.}\end{array}\right$.
The symbol $\land$ means ‘and’, $\lor$ means ‘or’, and $\neg$ means ‘negation’. Let $\langle\cdot,\cdot\rangle$ and $|\cdot|$ respectively denote the dot product and the Euclidean norm on $\mathbb{R}^{n}$. Let $B_{r}(x)$, $\overline{B}_{r}(x)$, and $S_{r}(x)$ respectively denote the open ball, closed ball, and sphere of center $x\in\mathbb{R}^{n}$ and radius $r\geq 0$. In particular, $B^{n}=\overline{B}_{1}(0)$ and $S^{n-1}=S_{1}(0)$.
Given a matrix $M\in\mathbb{R}^{m\times n}$, $M^{T}$ denotes the transpose. If $A:U\to V$ is a linear map between two finite dimensional inner product spaces, then $A^{*}:V\to U$ denotes the adjoint. Let $\|\cdot\|_{F}$, $\|\cdot\|_{1}$, $\|\cdot\|_{2}$, $\|\cdot\|_{*}$, $\|\cdot\|_{p,q}$, and $\rho(\cdot)$ respectively denote the Frobenius norm, entrywise $\ell_{1}$-norm, spectral norm, nuclear norm, $(p,q)$-induced norm, and spectral radius. Also, $\langle\cdot,\cdot\rangle_{F}$ denotes the Frobenius norm. Given a symmetric matrix $M\in\mathbb{R}^{n\times n}$, $\lambda_{1}(M)$ and $E_{1}(M)$ respectively denote its maximal eigenvalue and its associated eigenspace, whose nonzero elements are referred to as maximal eigenvectors.
2.1 Ordered sets
A binary relation $R$ on a set $X$ is a subset of $X\times X$. An element $x\in X$ is related to $y\in X$, denoted $xRy$, if $(x,y)\in R$. A relation $R$ is
  1. (i)
    reflexive if $\forall x\in X,~xRx$;
  2. (ii)
    irreflexive if $\forall x\in X,~\neg xRx$;
  3. (iii)
    transitive if $\forall x,y,z\in X,~xRy\land yRz\Longrightarrow xRz$;
  4. (iv)
    antisymmetric if $\forall x,y\in X,~xRy\land yRx\Longrightarrow x=y$;
  5. (v)
    total if $\forall x,y\in X,x\neq y\Longrightarrow x\leq y\lor y\leq x$.
[…]
2.2 Differential calculus
Given some vector spaces $V_{1},\ldots,V_{k}$ and $W$, a map $F:V_{1}\times\cdots\times V_{k}\rightarrow W$ is multilinear if it is linear in each variable taken separately when the others are held fixed. It is symmetric if $V_{1}=\cdots=V_{k}$ and $F(x_{\sigma(1)},\ldots,x_{\sigma(k)})=F(x_{1},\ldots,x_{k})$ for any permutation $\sigma$ of $\llbracket 1,k\rrbracket$. When it is real-valued, i.e., $W=\mathbb{R}$, $F$ is called a multilinear form. Given a symmetric multilinear map $F:V\times\cdots\times V\to W$ and norms $\|\cdot\|_{V},\|\cdot\|_{W}$ on $V,W$, consider the norm [7, Theorem A]
$\|F\|=\sup_{\|v\|_{V}=1}\|F(v,\ldots,v)\|_{W}.$
A function $f:U\to W$ where $U\subseteq V$ is open is Fréchet differentiable [10] at $\overline{x}\in U$, or $D^{1}$ at $\overline{x}$, if there exists a bounded linear map, denoted $f^{\prime}(\overline{x}):V\to W$, such that
$f(x)=f(\overline{x})+f^{\prime}(\overline{x})(x-\overline{x})+o(|x-\overline{x}|)$
where $o:\mathbb{R}_{+}\to W$ is a function such that $o(t)/t\to 0$ at $t\searrow 0$. We say that $f$ is $C^{1}$ at $\overline{x}$ if $f$ is $D^{1}$ near $\overline{x}$ and $f^{\prime}$ is continuous at $\overline{x}$. By induction, for any integer $k\geq 2$, we may define $f$ to be $D^{k}$ (resp. $C^{k}$) at $\overline{x}$ if it is $D^{k-1}$ (resp. $C^{k-1}$) near $\overline{x}$ and $f^{(k-1)}$ is $D^{1}$ (resp. $C^{1}$) at $\overline{x}$. We only define $D^{k}$ and $C^{k}$ for $k\in\mathbb{N}^{*}$. We say $f$ is $D^{k}$ (resp. $C^{k}$) if it is so at every point in $U$. Also, $f$ is $C^{k,\ell}$ (resp. $C^{k,\ell}_{L}$) if it is $C^{k}$ and $f^{(\ell)}$ is locally Lipschitz continuous (resp. $L$-Lipschitz continuous). If $f$ is $D^{k}$ at $\overline{x}$, then $f^{(k)}$ is a symmetric multilinear map [17, Proposition C.16]. When evaluated on the diagonal, we use the shorthand $f^{(k)}(x)v^{k}=f^{(k)}(x)(v,\ldots,v)$.
If $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is $D^{1}$, then we identify $f^{\prime}$ with the Jacobian $(\partial f_{i}/\partial x_{j})_{i,j}$. When $f:\mathbb{R}^{n}\to\mathbb{R}$ is $D^{1}$ at $\overline{x}$, then there exists a unique vector $\nabla f(\overline{x})\in\mathbb{R}^{n}$ such that $f^{\prime}(\overline{x})(v)=\langle\nabla f(\overline{x}),v\rangle$ by the Riesz-Fréchet representation theorem. Likewise, there exists a unique matrix $\nabla^{2}f(\overline{x})\in\mathbb{R}^{n\times n}$ such that $f^{\prime\prime}(\overline{x})(v_{1},v_{2})=\langle\nabla^{2}f(\overline{x})v_{1},v_{2}\rangle$ for all $v_{1},v_{2}\in\mathbb{R}^{n}$. Given $v_{2},\ldots,v_{n}\in\mathbb{R}^{n}$, we let $\nabla^{k}f(\overline{x})(v_{2},\ldots,v_{n})\in\mathbb{R}^{n}$ denote the unique vector such that $f^{(k)}(\overline{x})(v_{1},v_{2},\ldots,v_{n})=\langle v_{1},\nabla^{k}f(\overline{x})(v_{2},\ldots,v_{n})\rangle$ for all $v_{1}\in\mathbb{R}^{n}$. When $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is $D^{2}$ at $\overline{x}$, $f^{\prime\prime}(\overline{x})$ is a symmetric bilinear form and
$\|f^{\prime\prime}(\overline{x})\|=\max_{|v|=1}|\langle\nabla^{2}f(\overline{x})v,v\rangle|=\max_{|v|=1}|\nabla^{2}f(\overline{x})v|=\|\nabla^{2}f(\overline{x})\|_{2}=\rho(\nabla^{2}f(\overline{x})).$
If $\overline{x}$ is a local minimum of $f$, then $\|f^{\prime\prime}(\overline{x})\|=\lambda_{1}(\nabla^{2}f(\overline{x}))$ since $\nabla^{2}f(\overline{x})$ is positive semidefinite.
2.3 Differential geometry
An action of a group $G$ with identity $e$ on a set $M$ [27, p. 161] is a map $\theta:G\times M\rightarrow M$ such that
  1. (i)
    $\forall g,h\in G,~\forall x\in M,~\theta(g,\theta(h,x))=\theta(gh,x)$,
  2. (ii)
    $\forall x\in M,~\theta(e,x)=x$.
When an action exists, $M$ is referred to as a $G$-space. It is homogeneous if for all $x,y\in M$, there exists $g\in G$ such that $\theta(g,x)=y$. A function $f:M\rightarrow N$ between sets $M$ and $N$ is invariant under an action of a group $G$ on $M$ if $f(gx)=f(x)$ for all $g\in G$ and $x\in M$.
A Lie group $G$ is a smooth manifold and a group whose operations are smooth. A Lie group $G$ acts smoothly on a smooth manifold $M$ if there exists a smooth action $\theta:G\times M\rightarrow M$. A Lie subgroup of a Lie group $G$ is a subgroup of $G$ is endowed with a topology and smooth structure making into a Lie group and an immersed submanifold. Topologically closed subgroups of Lie groups are Lie subgroups by the closed subgroup theorem [27, Theorem 20.12]. Let $\mathfrak{g}$ denote the Lie algebra of a Lie group $G$, which we identify with its tangent space at $e$.
Let $I_{n}$ denote the identity matrix of order $n$. The set of invertible matrices with real coefficients of order $n$, denoted $\mathrm{GL}(n,\mathbb{R})$, is a Lie group. The natural action of a Lie subgroup $G$ of $\mathrm{GL}(n,\mathbb{R})$ on $\mathbb{R}^{n}$ is defined by the matrix vector multiplication $G\times\mathbb{R}^{n}\ni(g,x)\mapsto gx\in\mathbb{R}^{n}$. The orthogonal group $\mathrm{O}(n)=\{Q\in\mathbb{R}^{n\times n}:Q^{T}Q=I_{n}\}$ is a Lie subgroup of $\mathrm{GL}(n,\mathbb{R})$.
Given a Lipschitz continuous function $F:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $x_{0}\in\mathbb{R}^{n}$, consider the ODE
$\left\{\begin{array}[]{ccc}\dot{x}&=&F(x)\\ x(0)&=&x_{0}.\end{array}\right$.
Suppose it has a unique global solution $x:\mathbb{R}\to\mathbb{R}^{n}$ for every initial point $x_{0}$. Then one can define the global flow $\theta:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ by $\theta(t,x_{0})=x(t)$. If $F$ is $C^{k}$, then $\theta$ is a $C^{k}$ smooth action of $\mathbb{R}$ on $\mathbb{R}^{n}$. In that case, $\theta_{t}=\theta(t,\cdot)$ is a $C^{k}$ diffeomorphism with inverse $\theta_{-t}$ for any $t\in\mathbb{R}$.
2.4 Variational analysis
Given $X\subseteq\mathbb{R}^{n}$ and $f:X\to\overline{\mathbb{R}}$, a point $\overline{x}\in X$ where $f(\overline{x})$ is finite is a local minimum (resp. strict local minimum) of $f$ if there exists a neighborhood $U$ of $\overline{x}$ in $X$ such that $f(\overline{x})\leq f(x)$ (resp. $f(\overline{x})<f(x)$) for all $x\in U\setminus\{\overline{x}\}$. Let
$\arg\mathrm{loc}\min_{X}f=\{\overline{x}\in X:\overline{x}~\text{is a local minimum of}~f\}.$
In amounts to an abuse of notation, we will alternatively write this as $\arg\mathrm{loc}\min\{f(x):x\in X\}$. A point $\overline{x}\in X$ where $f(\overline{x})$ is finite is a global minimum of $f$ if there exists a neighborhood $U$ of $\overline{x}$ in $\mathbb{R}^{n}$ such that $f(\overline{x})\leq f(x)$ for all $x\in\mathbb{R}^{n}$. Accordingly, $\min_{X}f=\min\{f(x):x\in X\}$ and
$\arg\min_{X}f=\{x\in X:f(x)=\min_{X}f\},$
which we will also denote by $\arg\min\{f(x):x\in X\}$. A local minimum is spurious if it is not a global minimum.
It will be convenient to use the generalization of local optiamlity from points to sets proposed in [25, 24]. A nonempty set $S\subseteq\mathbb{R}^{n}$ is a local minimum (resp. strict) of $f:\mathbb{R}^{n}\to\mathbb{R}$ if there exists a neighborhood $U$ of $S$ such that $f(x)\leq f(y)$ (resp. $f(x)<f(y)$) for all $x\in S$ and $y\in U\setminus S$. In constrast to [24], we do not assume $S$ to be closed.
Given $x\in\mathbb{R}^{n}$ and $S\subseteq\mathbb{R}^{n}$, let
$d(x,S)=\inf\{|x-y|:y\in S\}~~~\text{and}~~~P_{S}(x)=\arg\min\{|x-y|:y\in S\}.$
Given $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ and $\ell\in\mathbb{R}$, let
$f=\ell]=\{x\in\mathbb{R}^{n}:f(x)=\ell\}$
and define other expressions like $f\leq\ell]$ similarly. Let
$\mathrm{dom}f=\{x\in\mathbb{R}^{n}:f(x)<\infty\}~~~\text{and}~~~\mathrm{gph}f=\{(x,t)\in\mathbb{R}^{n}\times\mathbb{R}:f(x)=t\}.$
The indicator of $S\subseteq\mathbb{R}^{n}$ is defined by
$\delta_{S}(x)=\left\{\begin{array}[]{cl}0&\text{if}~x\in S,\\ \infty&\text{if}~x\notin S.\end{array}\right$.
A function $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is lower semicontinuous at $\overline{x}\in\mathrm{dom}f$ if $\liminf_{x\rightarrow\overline{x}}f(x)\geq f(\overline{x})$ [36, Definition 1.5]. It is lower semicontinuous if it is so at every point in its domain. The regular normal cone and normal cone [36, Definition 6.3] are defined by
$\widehat{N}_{C}(\overline{x})=\{v\in\mathbb{R}^{n}:\langle v,x-\overline{x}\rangle\leq o(|x-\overline{x}|)~\text{for}~x\in C~\text{near}~\overline{x}\},$
$N_{C}(\overline{x})=\{v\in\mathbb{R}^{n}:\exists x_{k}\xrightarrow[C]{}\overline{x}~\text{and}~\exists v_{k}\rightarrow v~\text{with}~v_{k}\in\widehat{N}_{C}(x_{k})\},$
where $x_{k}\xrightarrow[C]{}\overline{x}$ is a shorthand for $x_{k}\rightarrow\overline{x}$ and $x_{k}\in C$. Explicitly, the $o$ means that
$\limsup_{\scriptsize\begin{array}[]{c}x\xrightarrow[C]{}\overline{x}\\ x\neq\overline{x}\end{array}}\frac{\langle v,x-\overline{x}\rangle}{|x-\overline{x}|}\leqslant 0.$
A set $C\subseteq\mathbb{R}^{n}$ is regular [36, Definition 6.4] at one of its points $\overline{x}$ if it is locally closed and $\widehat{N}_{C}(\overline{x})=N_{C}(\overline{x})$.
Given $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ and a point $\overline{x}\in\mathbb{R}^{n}$ where $f(\overline{x})$ is finite, the regular subdifferential, subdifferential [36, Definition 8.3], and Clarke subdifferential of $f$ at $\overline{x}$ [14, Definition 4.1] are
$\widehat{\partial}f(\overline{x})=\{v\in\mathbb{R}^{n}:f(x)\geq f(\overline{x})+\langle v,x-\overline{x}\rangle+o(|x-\overline{x}|)~\text{near}~\overline{x}\},$
$\partial f(\overline{x})=\{v\in\mathbb{R}^{n}:\exists(x_{k},v_{k})\in\mathrm{gph}\hskip 1.42262pt\widehat{\partial}f:(x_{k},f(x_{k}),v_{k})\rightarrow(\overline{x},f(\overline{x}),v)\},$
$\overline{\partial}f(\overline{x})=\overline{\mathrm{co}}[\partial f(\overline{x})+\partial^{\infty}f(\overline{x})],$
where $\mathrm{co}$} denotes the convex hull, and $\overline{\mathrm{co}}$ its closure. The $o$ means that $\liminf[f(x)-f(\overline{x})-\langle v,x-\overline{x}\rangle]/|x-\overline{x}|\geq 0$ where $\overline{x}\neq x\to\overline{x}$. A point $x\in\mathbb{R}^{n}$ is critical (resp. Clarke critical) if $0\in\partial f(x)$ (resp. $0\in\overline{\partial}f(x)$). A function $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is regular [36, Definition 7.25] at $\overline{x}$ if $f(\overline{x})$ is finite and $\mathrm{epi}f$ is regular at $(\overline{x},f(\overline{x}))$ as a subset of $\mathbb{R}^{n+1}$. The Lipschitz modulus of a function $f:\mathbb{R}^{n}\to\mathbb{R}$ is defined by [36, p. 354]
$\mathrm{lip}f(\overline{x})=\limsup_{\scriptsize\begin{array}[]{c}x,y\rightarrow\overline{x}\\ x\neq y\end{array}}\frac{|f(x)-f(y)|}{|x-y|}.$
If $f$ is Lipschitz continuous near $\overline{x}$, i.e., $\mathrm{lip}f(\overline{x})<\infty$, then let
$\overline{\nabla}f(\overline{x})=\{v\in\mathbb{R}^{n}:\exists x_{k}\xrightarrow[D]{}\overline{x}~\text{with}~\nabla f(x_{k})\rightarrow v\}$
where $D$ are the differentiable points of $f$. In what amounts to a slight abuse of notation, we may replace $D$ by some $D^{\prime}\subseteq D$ such that $D\setminus D^{\prime}$ has zero Lebesgue measure. Still assuming $\mathrm{lip}f(\overline{x})<\infty$, by [36, Theorem 9.13] $\partial f(\overline{x})$ is nonempty and compact, and $\mathrm{lip}f(\overline{x})=\max\{|v|:v\in\partial f(\overline{x})\}$. By [36, Theorem 9.61], $\overline{\nabla}f(\overline{x})$ is a nonempty compact subset of $\partial f(\overline{x})$ and $\mathrm{co}\overline{\nabla}f(\overline{x})=\mathrm{co}\hskip 0.85358pt\partial f(\overline{x})=\overline{\partial}f(\overline{x})$. Thus $\arg\max\{|v|:v\in\partial f(\overline{x})\}=\arg\max\{|v|:v\in\overline{\nabla}f(\overline{x})\}$, using the following fact.
Fact 1.
For any set $X\subseteq\mathbb{R}^{n}$, $\arg\max\{|x|:x\in X\}=\arg\max\{|x|:x\in\mathrm{co}\hskip 0.85358ptX\}$.
Proof.
Let $\overline{x}\in\arg\max\{|x|:x\in X\}$ and $x\in\mathrm{co}\hskip 0.85358ptX$. There exist finitely many $t_{i}\geq 0$ with $\sum_{i}t_{i}=1$ and $x_{i}\in X$ such that $x=\sum_{i}t_{i}x_{i}$. Thus $|x|=\left|\sum_{i}t_{i}x_{i}\right|\leq\sum_{i}t_{i}|x_{i}|\leq\sum_{i}t_{i}|\overline{x}|=|\overline{x}|$ and $\overline{x}\in\arg\max\{|x|:x\in\mathrm{co}\hskip 0.85358ptX\}$. Conversely, let $\overline{x}\in\arg\max\{|x|:x\in\mathrm{co}\hskip 0.85358ptX\}$. There exist finitely many $t_{i}\geq 0$ with $\sum_{i}t_{i}=1$ and $x_{i}\in X\subseteq\mathrm{co}\hskip 0.85358ptX$ such that $\overline{x}=\sum_{i}t_{i}x_{i}$. If $\overline{x}\neq x_{j}$ for some $j$, then $t_{j}<1$ and one obtains the contradiction
$|\overline{x}|^{2}$ $=\left|(1-t_{j})\sum_{i\neq j}\frac{t_{i}}{1-t_{j}}x_{i}+t_{j}x_{j}\right|^{2}<(1-t_{j})\left|\sum_{i\neq j}\frac{t_{i}}{1-t_{j}}x_{i}\right|^{2}+t_{j}|x_{j}|^{2}$
$\leq(1-t_{j})\sum_{i\neq j}\frac{t_{i}}{1-t_{j}}|x_{i}|^{2}+t_{j}|x_{j}|^{2}=\sum_{i}t_{i}|x_{i}|^{2}\leq\sum_{i}t_{i}|\overline{x}|^{2}=|\overline{x}|^{2}$
using the strict convexity of $|\cdot|^{2}$. Thus $\overline{x}=x_{j}\in X$ for some $j$. ∎
2.5 O-minimal structures
O-minimal structures (short for order-minimal) were originally considered by van den Dries, Pillay, and Steinhorn [39, 35]. They are founded on the observation that many properties of semi-algebraic sets can be deduced from a few simple axioms [40]. Recall that a subset $A$ of $\mathbb{R}^{n}$ is semi-algebraic [6] if it is a finite union of basic semi-algebraic sets, which are of the form
$\{x\in\mathbb{R}^{n}:f_{1}(x)>0,\ldots,f_{p}(x)>0,f_{p+1}(x)=0,\ldots,f_{q}(x)=0\}$
where $f_{1},\ldots,f_{q}$ are polynomials with real coefficients. We adopt [41, Definition p. 503-506] below.
Definition 1.
An o-minimal structure on the real field is a sequence $S=(S_{k})_{k\in\mathbb{N}}$ such that for all $k\in\mathbb{N}$:
  1. 1.
    $S_{k}$ is a boolean algebra of subsets of $\mathbb{R}^{k}$, with $\mathbb{R}^{k}\in S_{k}$;
  2. 2.
    $S_{k}$ contains the diagonal $\{(x_{1},\ldots,x_{k})\in\mathbb{R}^{k}:x_{i}=x_{j}\}$ for $1\leqslant i<j\leqslant k$;
  3. 3.
    If $A\in S_{k}$, then $A\times\mathbb{R}$ and $\mathbb{R}\times A$ belong to $S_{k+1}$;
  4. 4.
    If $A\in S_{k+1}$ and $\pi:\mathbb{R}^{k+1}\rightarrow\mathbb{R}^{k}$ is the projection onto the first $k$ coordinates, then $\pi(A)\in S_{k}$;
  5. 5.
    $S_{3}$ contains the graphs of addition and multiplication;
  6. 6.
    $S_{1}$ consists exactly of the finite unions of open intervals and singletons.
A subset $A$ of $\mathbb{R}^{n}$ is definable in an o-minimal structure $(S_{k})_{k\in\mathbb{N}}$ if $A\in S_{k}$ for some $k\in\mathbb{N}$. A function $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is definable in an o-minimal structure if $\mathrm{gph}f$ is definable in that structure. Throughout this paper, we fix an arbitrary o-minimal structure $(S_{k})_{k\in\mathbb{N}}$.
A key property of univariate definable functions is that they satisfy the monotonicity theorem [41, 4.1]. It states that on bounded open intervals, for any $p\in\mathbb{N}$ there exist finitely many open subintervals where the function is $C^{p}$ and either constant or strictly monotone. The extension of the monotonicity theorem to multivariate functions is the cell decomposition theorem [41, 4.2], which is used to prove the definable Morse-Sard theorem [8, Corollary 9]. It asserts that that lower semi-continuous definable functions have finitely many Clarke critical values.
3 Basic aspects
We investigate two dual viewpoints on flatness. To each one naturally corresponds an optimal curve emanating from the point of interest. We study its properties in preparation of future sections and develop calculus rules.
3.1 Definition of flatness
Below are two ways to measure the variation of a function around a point.
Definition 2.
Given $f:\mathbb{R}^{n}\to\mathbb{R}$, let $\accentset{\circ}{f},\overline{f}:\mathbb{R}^{n}\times\mathbb{R}_{+}\to\overline{\mathbb{R}}$ be defined by
$\accentset{\circ}{f}(x,r)=\sup_{\overline{B}_{r}(x)}|f-f(x)|~~~\text{and}~~~\overline{f}(x,\ell)=d(x,[|f-f(x)|\geq\ell]).$
The following definition is the central tenet of this paper.
Definition 3.
Consider the binary relation on $\mathbb{R}^{n}$ defined by
$x\preceq y~~~\Longleftrightarrow~~~\exists\overline{r}>0:~\forall r\in(0,\overline{r}],~~~\accentset{\circ}{f}(x,r)\leq\accentset{\circ}{f}(y,r).$
We say $x$ is flatter than $y$, or $y$ is sharper than $x$ if $x\prec y$. We say $\overline{x}$ is flat (resp. strictly flat) if there exists a neighborhood $U$ of $\overline{x}$ in $f=f(\overline{x})]$ such that $\overline{x}\preceq x$ (resp. $\overline{x}\prec x$) for all $x\in U\setminus\{\overline{x}\}$. We say $\overline{x}$ is globally flat if $\overline{x}\preceq x$ for all $x\in[f=f(\overline{x})]$.
A function $f:\mathbb{R}^{n}\to\mathbb{R}$ is constant near $x\in\mathbb{R}^{n}$ if it is constant on a neighborhood of $x$. Otherwise, $f$ is nonconstant near $x$, which happens iff $\accentset{\circ}{f}(x,r)>0$ for all $r>0$ iff $x\notin\mathrm{int}[f=f(x)]$. Given $x$, $f$ is constant iff $\overline{f}(x,\ell)=\infty$ for all $\ell>0$, and $f$ is continuous at $x$ iff $\overline{f}(x,\ell)>0$ for all $\ell>0$. A set $X\subset\mathbb{R}^{n}$ is nowhere dense if its closure has empty interior. If $f$ is continuous and $\alpha\in\mathbb{R}$, then $f$ is nonconstant near every point in $f=\alpha]$ iff $f=\alpha]$ is nowhere dense.
A function $\varphi:S\to\overline{\mathbb{R}}$ where $S\subset\mathbb{R}$ is increasing (resp. strictly increasing) if $\varphi(s)\leq\varphi(t)$ (resp. $\varphi(s)<\varphi(t)$) for all $s,t\in S$ such that $s<t$. It is monotone (resp. strictly monotone) if either $\varphi$ or $\varphi$ is increasing (resp. strictly increasing). Given $x\in\mathbb{R}^{n}$, let $\accentset{\circ}{f}_{x}=\accentset{\circ}{f}(x,\cdot)$ and $\overline{f}_{x}=\overline{f}(x,\cdot)$, both of which are increasing.
Fact 2.
$\preceq$ is a preorder. If $f$ is definable, then it is a total preorder and
$\forall x,y\in\mathbb{R}^{n},~~~x\prec y~~\Longleftrightarrow~~\exists\overline{r}>0:~\forall r\in(0,\overline{r}],~~~\accentset{\circ}{f}(x,r)<\accentset{\circ}{f}(y,r).$
Proof.
[…]
We now gather several useful facts when $f$ is continuous (i.e., Facts 3-8).
Fact 3.
$\forall(x,\ell)\in\mathbb{R}^{n}\times\mathbb{R}_{+},~P_{[|f-f(x)|\geq\ell]}(x)=P_{[|f-f(x)|=\ell]}(x)~\land~\overline{f}(x,\ell)=d(x,[|f-f(x)|=\ell])$.
Proof.
Let $y\in P_{[|f-f(x)|\geq\ell]}(x)$. If $|f(y)-f(x)|>\ell$, then by intermediary value theorem there exists $z\in(x,y)$ such that $|f(z)-f(x)|=\ell$. Since |xz|<|xy||x-z|<|x-y| and $z\in[|f-f(x)|\geq\ell]$, this contradicts the optimality of $y$. Thus $|f(y)-f(x)|=\ell$ and $y\in P_{[|f-f(x)|=\ell]}(x)$. As for the second equality, $\overline{f}(x,\ell)\leq d(x,[|f-f(x)|=\ell])$ is obvious. Let $y_{k}\in\mathbb{R}^{n}$ be such that $|y_{k}-x|\to\overline{f}(x,\ell)$ and $|f(y_{k})-f(x)|\geq\ell$. By the intermediary value theorem, there exists $z_{k}\in[x,y_{k}]$ such that $|f(z_{k})-f(x)|=\ell$. Hence $d(x,[|f-f(x)|=\ell])\leq|z_{k}-x|\leq|y_{k}-x|$ and so $d(x,[|f-f(x)|=\ell])\leq\overline{f}(x,\ell)$. ∎
There exists a natural duality between $\accentset{\circ}{f}$ and $\overline{f}$.
Fact 4.
$\forall(x,\ell)\in\mathbb{R}^{n}\times\mathbb{R}^{+}$, $\overline{f}(x,\ell)=\inf\{r\geq 0:\accentset{\circ}{f}(x,r)\geq\ell\}$.
Proof.
$\overline{f}(x,\ell)=d(x,[|f-f(x)|\geq\ell])=\inf\{r\geq 0:\exists y\in\overline{B}_{r}(x):|f(y)-f(x)|\geq\ell\}=\inf\{r\geq 0:\sup_{y\in\overline{B}_{r}(x)}|f(y)-f(x)|\geq\ell\}=\inf\{r\geq 0:\accentset{\circ}{f}(x,r)\geq\ell\}$. ∎
Fact 5.
$\forall(x,r)\in\mathbb{R}^{n}\times\mathbb{R}_{+}$, $\accentset{\circ}{f}(x,r)=\sup\{\ell\geq 0:\overline{f}(x,\ell)\leq r\}$.
Proof.
$\accentset{\circ}{f}(x,r)=\sup_{\overline{B}_{r}(x)}|f-f(x)|=\sup\{\ell\geq 0:\exists y\in\overline{B}_{r}(x),|f(y)-f(x)|\geq\ell\}=\sup\{\ell\geq 0:d(x,[|f-f(x)|\geq\ell])\leq r\}=\sup\{\ell\geq 0:\overline{f}(x,\ell)\leq r\}$. ∎
˜4 and ˜5 actually hold without continuity if one uses strict relations, including in the definition of $\accentset{\circ}{f}$ and $\overline{f}$, with the convention that $\accentset{\circ}{f}(x,0)=\overline{f}(x,0)=0$.
Fact 6.
If $f$ is nonconstant near $x$, then the following are equivalent:
  1. (i)
    $\exists\overline{r}>0:~\forall r\in(0,\overline{r}],~~\accentset{\circ}{f}(x,r)\leq\accentset{\circ}{f}(y,r)$;
  2. (ii)
    $\exists\overline{\ell}>0:~\forall\ell\in(0,\overline{\ell}],~~\overline{f}(x,\ell)\geq\overline{f}(y,\ell)$.
Proof.
(i) $\Longrightarrow$ (ii) Let $\overline{\ell}=\accentset{\circ}{f}(x,\overline{r})>0$ and $\ell\in(0,\overline{\ell}]$. By ˜4,
$\overline{f}(x,\ell)$ $=\inf\{r\geq 0:\accentset{\circ}{f}(x,r)\geq\ell\}$
$=\inf\{r\in[0,\overline{r}]:\accentset{\circ}{f}(x,r)\geq\ell\}$
$\geq\inf\{r\in[0,\overline{r}]:\accentset{\circ}{f}(y,r)\geq\ell\}$
$\geq\inf\{r\geq 0:\accentset{\circ}{f}(y,r)\geq\ell\}=\overline{f}(y,\ell).$
(ii) $\Longrightarrow$ (i) Since $f$ is nonconstant, $\infty>\overline{f}(x,\overline{\ell})\geq\overline{f}(y,\overline{\ell})>0$ after possibly reducing $\overline{\ell}$. Let $\overline{r}=\accentset{\circ}{f}(y,\overline{\ell})/2>0$ and $r\in(0,\overline{r}]$. Since $\overline{f}(x,\ell)\geq\overline{f}(x,\overline{\ell})\geq\overline{f}(y,\overline{\ell})=2\overline{r}>r$ for all $\ell\in(\overline{\ell},\infty)$, by ˜5,
$\accentset{\circ}{f}(x,r)$ $=\sup\{\ell\geq 0:\overline{f}(x,\ell)\leq r\}$
$=\sup\{\ell\in[0,\overline{\ell}]:\overline{f}(x,\ell)\leq r\}$
$\leq\sup\{\ell\in[0,\overline{\ell}]:\overline{f}(y,\ell)\leq r\}$
$\leq\sup\{\ell\geq 0:\overline{f}(y,\ell)\leq r\}=\accentset{\circ}{f}(y,r).$
Fact 7.
$\accentset{\circ}{f}$ is continuous.
Proof.
Let $\mathbb{R}^{n}\times\mathbb{R}_{+}\ni(x_{k},r_{k})\to(x,r)$. By continuity, there exists $y_{k}\in\overline{B}_{r_{k}}(x_{k})$ such that $\accentset{\circ}{f}(x_{k},r_{k})=|f(y_{k})-f(x_{k})|$, namely, $|f(y_{k})-f(x_{k})|\geq|f(z)-f(x_{k})|$ for all $z\in\overline{B}_{r_{k}}(x_{k})$. By compactness, $y_{k}\to y$ after taking a subsequence. For all $\epsilon>0$, we eventually have $B_{r-\epsilon}(x)\subseteq\overline{B}_{r_{k}}(x_{k})$ and so $|f(y_{k})-f(x_{k})|\geq|f(z)-f(x_{k})|$ for all $z\in B_{r-\epsilon}(x)$. Passing to the limit yields $|f(y)-f(x)|\geq|f(z)-f(x)|$ for all $z\in B_{r-\epsilon}(x)$. As $\epsilon$ is arbitrary, we have $|f(y)-f(x)|\geq|f(z)-f(x)|$ for all $z\in B_{r}(x)$. By continuity of $f$, this holds for all $z\in\overline{B}_{r}(x)$ and thus $\accentset{\circ}{f}(x_{k},r_{k})=|f(y_{k})-f(x_{k})|\to|f(y)-f(x)|=\accentset{\circ}{f}(x,r)$. ∎
Fact 8.
$\overline{f}$ is lower semicontinuous.
Proof.
Let $(x_{k},\ell_{k})\to(x,\ell)\in\mathbb{R}^{n}\times\mathbb{R}_{+}$ be such that $r=\lim\inf\overline{f}(x_{k},\ell_{k})<\infty$. There exists $y_{k}\in\mathbb{R}^{n}$ such that $|x_{k}-y_{k}|\to r$ and $|f(x_{k})-f(y_{k})|\geq\ell_{k}$ up to a subsequence. By compactness, $y_{k}\to y$ up to another subsequence. Since $f$ is continuous, we may pass to the limit: |xy|=r|x-y|=r and $|f(x)-f(y)|\geq\ell$. Thus $\lim\inf\overline{f}(x_{k},\ell_{k})=r\geq\overline{f}(x,\ell)$. ∎
Since $\accentset{\circ}{f}$ is continuous, one has equality in ˜4, namely, $\overline{f}(x,\ell)=\inf\{r\geq 0:\accentset{\circ}{f}(x,r)=\ell\}$. This holds by the intermediate value theorem as $\accentset{\circ}{f}(x,0)=0$. In contrast, since $\overline{f}$ is merely lower semicontinuous, equality does not necessarily hold in ˜5 (think of the Cantor function). Likewise, $\accentset{\circ}{f}_{x}$ need not be strictly increasing. To obtain such properties and others, more assumptions are needed.
Assumption 1.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz and nonconstant near $x\in\mathbb{R}^{n}$. Suppose $\mathbb{R}\setminus\{f(x)\}$ contains no Clarke critical value of $f$ reached in $\overline{B}_{\overline{r}}(x)$ for some $\overline{r}>0$. Let $\overline{\ell}=\accentset{\circ}{f}(x,\overline{r})$.
Proposition 1.
Under ˜1,
  1. (i)
    $\accentset{\circ}{f}_{x}$ and $\overline{f}_{x}$ are continuous and strictly increasing on $0,\overline{r}]$ and $0,\overline{\ell}]$ respectively.
  2. (ii)
    $\forall(r,\ell)\in[0,\overline{r}]\times[0,\overline{\ell}],~(\accentset{\circ}{f}_{x}\circ\overline{f}_{x})(r)=r$ and $(\overline{f}_{x}\circ\accentset{\circ}{f}_{x})(\ell)=\ell$.
  3. (iii)
    $\forall(r,\ell)\in\mathrm{gph}\accentset{\circ}{f}_{x}|_{[0,\overline{r}]},~\arg\max_{\overline{B}_{r}(x)}|f-f(x)|=\arg\max_{S_{r}(x)}|f-f(x)|=P_{[|f-f(x)|=\ell]}(x)$.
Proof.
(i) & (ii) Continuity of $\accentset{\circ}{f}_{x}$ follows from ˜7. Clearly $\accentset{\circ}{f}_{x}$ is increasing on $0,\overline{r}]$. It is in fact strictly increasing. Indeed, suppose $\accentset{\circ}{f}_{x}$ is constant near $r\in(0,\overline{r})$. If $\accentset{\circ}{f}(x,r)=0$, then $f$ is constant near $x$, violating our assumption. Otherwise, by continuity there exists $y\in\overline{B}_{r}(x)$ such that $\accentset{\circ}{f}(x,r)=|f(y)-f(x)|$, which must be a local maximum of |ff(x)||f-f(x)|. Since $\accentset{\circ}{f}(x,r)>0$, $y$ is either a local minimum or a local maximum of $f$, and hence $0\in\partial f(y)=\mathrm{co}\hskip 0.85358pt\partial f(y)=\overline{\partial}f(y)$ or $0\in\partial(-f)(y)\subseteq\overline{\partial}(-f)(y)=-\overline{\partial}f(y)$ by Fermat’s rule [27, Theorem 10.1]. Since $\mathbb{R}\setminus\{f(x)\}$ contains no Clarke critical value of $f$ reached in $\overline{B}_{\overline{r}}(x)$, we have $f(y)=f(x)$. It follows that $\accentset{\circ}{f}(x,r)=0$, a contradiction.
By ˜4 and the discussion below ˜8, $\accentset{\circ}{f}_{x}(\overline{f}_{x}(\ell))=\ell$ for all $\ell\in[0,\overline{\ell}]$. Thus for all $r\in[0,\overline{r}]$, $\accentset{\circ}{f}_{x}((\overline{f}_{x}\circ\accentset{\circ}{f}_{x})(r))=(\accentset{\circ}{f}_{x}\circ\overline{f}_{x})(\accentset{\circ}{f}_{x}(r))=\accentset{\circ}{f}_{x}(r)$ and $(\overline{f}_{x}\circ\accentset{\circ}{f}_{x})(r)=r$. It follows that $\overline{f}_{x}$ is continuous and strictly increasing.
(iii) Let $r\in[0,\overline{r}]$ and $y\in\arg\max_{\overline{B}_{r}(x)}|f-f(x)|$. As $\accentset{\circ}{f}_{x}(r)=|f(y)-f(x)|\leq\accentset{\circ}{f}(x,|y-x|)$ and $\accentset{\circ}{f}_{x}$ is strictly increasing, $r\leq|x-y|$. Hence $y\in\arg\max_{S_{r}(x)}|f-f(x)|$. Let $\ell=\accentset{\circ}{f}_{x}(r)$. $y\in\arg\max_{S_{r}(x)}|f-f(x)|$ iff $\accentset{\circ}{f}_{x}(r)=|f(y)-f(x)|$ and |xy|=r|x-y|=r iff $\accentset{\circ}{f}_{x}(|x-y|)=\ell=|f(y)-f(x)|=\accentset{\circ}{f}_{x}(r)$ iff $\overline{f}_{x}(\ell)=|x-y|$ and $|f(y)-f(x)|=\ell$ iff $y\in P_{[|f-f(x)|=\ell]}(x)$. ∎
Flatness can be equivalently defined using $\overline{f}$ under minimal assumptions.
Proposition 2.
Suppose $f:\mathbb{R}^{n}\to\mathbb{R}$ is continuous and $f=\alpha]$ is nowhere dense for some $\alpha\in\mathbb{R}$. Then
$\forall x,y\in[f=\alpha],~~~x\preceq y~~\Longleftrightarrow~~\exists\overline{\ell}>0:~\forall\ell\in(0,\overline{\ell}],~~~\overline{f}(x,\ell)\geq\overline{f}(y,\ell).$
If in addition, $f$ is locally Lipschitz definable, then
$\forall x,y\in[f=\alpha],~~~x\prec y~~\Longleftrightarrow~~\exists\overline{\ell}>0:~\forall\ell\in(0,\overline{\ell}],~~~\overline{f}(x,\ell)>\overline{f}(y,\ell).$
Proof.
The first point follows from ˜6, while the second follows from Facts 2-7, Proposition˜1, and the definable Morse-Sard theorem [8, Corollary 9]. ∎
3.2 Curve selection
With the above definitions in place, it is natural to try and select optimal curves. This will be useful later.
Lemma 1.
Under ˜1,
  1. (i)
    There exist $\gamma:[0,\overline{r}]\to\mathbb{R}^{n}$, $\lambda:[0,\overline{r}]\to\mathbb{R}$, and $v:[0,\overline{r}]\to\mathbb{R}^{n}$ such that for all $r\in[0,\overline{r}]$, $\gamma(r)=x+\lambda(r)v(r)\in\arg\max_{\overline{B}_{r}(x)}|f-f(x)|$ and $v(r)\in\overline{\partial}f(\gamma(r))$.
  2. (ii)
    If $x$ is a local minimum of $f$, then $\lambda(r)>0$ and $v(r)\in-\partial(-f)(\gamma(r))$ for all $r\in(0,\overline{r}]$ after possibly reducing $\overline{r}$. If in addition, $f$ is regular near $x$, then $v(r)\in\partial f(\gamma(r))$ for all $r\in(0,\overline{r}]$.
  3. (iii)
    If $f$ is definable, then $\gamma,\lambda,v$ can be made $C^{k}$ definable on $(0,\overline{r}]$ for any $k\in\mathbb{N}$, such that $\gamma$ is $C^{1}$ on $0,\overline{r}]$, and $v$ continuous on $0,\overline{r}]$, after possibly reducing $\overline{r}$.
  4. (iv)
    If $x$ is a local minimum of $f$ and $f$ is definable and differentiable, then $\lambda(r)>0$, $\accentset{\circ}{f}_{x}$ is differentiable on $0,\overline{r}]$, and $(\accentset{\circ}{f}_{x})^{\prime}(r)=r/\lambda(r)$ for all $r\in(0,\overline{r}]$.
Proof.
(i) Let $(r,\ell)\in\mathrm{gph}\accentset{\circ}{f}_{x}$. There exists $\gamma(r)\in\overline{B}_{r}(x)$ such that $\accentset{\circ}{f}(x,r)=|f(\gamma(r))-f(x)|$. By [36, Theorem 8.15], $0\in\partial|f-f(x)|(\gamma(r))+N_{\overline{B}_{r}(x)}(\gamma(r))$. If $r=0$, then let $\lambda(0)=0$ and $v(0)=0$. Otherwise, there exist $w(r)\in\partial(s(r)f)(\gamma(r))\subseteq\mathrm{co}\hskip 0.85358pt\partial(s(r)f)(\gamma(r))=\overline{\partial}(s(r)f)(\gamma(r))=s(r)\overline{\partial}f(\gamma(r))$ with $s(r)=\mathrm{sign}(f(x)-f(\gamma(r)))$ and $\mu(r)\geq 0$ such that $w(r)+\mu(r)(\gamma(r)-x)=0$. If $\mu(r)=0$, then $w(r)=0$ and so $f(\gamma(r))$ is a Clarke critical value of $f$. By assumption $f(\gamma(r))=f(x)$ and thus $\accentset{\circ}{f}_{x}(r)=0$, i.e., $r=0$, a contradiction. Thus $\mu(r)>0$. We may hence define $\lambda(r)=-s(r)/\mu(r)$ and $v(r)=s(r)w(r)$.
(ii) When $x$ is a local minimum of $f$, s(r)=1s(r)=-1. By [36, Theorem 9.16],
$\lim_{\tau\searrow 0}\frac{f(\gamma(r)+\tau w)-f(\gamma(r))}{\tau}=\lim_{\scriptsize\begin{array}[]{c}\tau\searrow 0\\ w^{\prime}\to w\end{array}}\frac{f(\gamma(r)+\tau w^{\prime})-f(\gamma(r))}{\tau}=\max_{v\in\partial f(\gamma(r))}\langle w,v\rangle.$
Since the directional derivative is nonpositive for any direction $w$ such that $\langle w,\gamma(r)-x\rangle<0$, this also holds for any direction $w$ such that $\langle w,\gamma(r)-x\rangle=0$. Given $v\in\partial f(\gamma(r))$, let $w=v|\gamma(r)-x|^{2}-\langle v,\gamma(r)-x\rangle(\gamma(r)-x)$. Since $\langle w,\gamma(r)-x\rangle=0$, it follows that $\langle w,v\rangle=|v|^{2}|\gamma(r)-x|^{2}-\langle v,\gamma(r)-x\rangle^{2}\leq 0$. There is equality in the Cauchy-Schwarz inequality, so there exists $\mu\in\mathbb{R}$ such that $v=\mu(\gamma(r)-x)$. Since $\langle x-\gamma(r),\gamma(r)-x\rangle<0$, we have $\langle x-\gamma(r),v\rangle=-\mu|\gamma(r)-x|^{2}\leq 0$ and so $\mu\geq 0$. Since $\partial f(\gamma(r))\neq\emptyset$ by [36, Theorem 9.13] and $\mathbb{R}\setminus\{f(x)\}$ contains no Clarke critical value of $f$ reached in $\overline{B}_{\overline{r}(x)}$, there exists $v(r)\in\partial f(\gamma(r))\setminus\{0\}$ and it satisfies $\gamma(r)-x=\lambda(r)v(r)$ for some $\lambda(r)>0$.
(iii) When $f$ is definable, $0,\overline{r}]\ni r\rightrightarrows\arg\max_{\overline{B}_{r}(x)}|f-f(x)|$ is definable, there exists a definable selection $\gamma$ [41, 4.5]. Consequently $0,\overline{r}]\ni r\rightrightarrows\{v\in\overline{\partial}f(\gamma(t)):\exists\lambda\in\mathbb{R}:\gamma(t)=x+\lambda v\}$ is definable, yielding a definable selection $v$. Finally, $\lambda(r)=\langle\gamma(r)-x,v(r)\rangle/|v(r)|^{2}$ is definable. The monotonicity theorem ensures $\gamma,\lambda,v$ are $C^{k}$ for any $k\in\mathbb{N}$ on $(0,\overline{r}]$ after possibly reducing $\overline{r}$. Also $|\gamma_{i}(r)|\leq|\gamma(r)|=r$ so $|\gamma^{\prime}_{i}(r)|\leq 1$ on $(0,\overline{r}]$ after possibly reducing $\overline{r}$ and $|\gamma^{\prime}(r)|=\sqrt{|\gamma^{\prime}_{1}(r)|^{2}+\cdots+|\gamma^{\prime}_{n}(r)|^{2}}\leq\sqrt{n}$. Hence $\lim_{r\searrow 0}\gamma^{\prime}(r)$ exists. By the mean value theorem, $\gamma_{i}^{\prime}(0)=\lim_{r\searrow 0}(\gamma_{i}(r)-x_{i})/r=\lim_{r\searrow 0}\gamma_{i}^{\prime}(r)$. Finally, one can take $v(0)=\lim_{r\searrow 0}v(r)\in\overline{\partial}f(\gamma(0))$, where the limit exists because $v(r)$ is bounded by [36, Theorem 9.13] and [4, Proposition 3 p. 42].
(iv) With $\nabla f(\gamma(r))+\mu(r)(\gamma(r)-x)=0$, we have
$\accentset{\circ}{f}_{x}(r+\epsilon)-\accentset{\circ}{f}_{x}(r)$ $=f(\gamma(r+\epsilon))-f(\gamma(r))$
$=\langle\nabla f(\gamma(r)),\gamma(r+\epsilon)-\gamma(r)\rangle+o(|\gamma(r+\epsilon)-\gamma(r)|)$
$=\mu(r)\langle\gamma(r)-x,\gamma(r+\epsilon)-\gamma(r)\rangle+o(|\gamma(r+\epsilon)-\gamma(r)|)$
$=\mu(r)(|\gamma(r+\epsilon)|^{2}-|\gamma(r)|^{2})/2+o((\mu(r)+1)|\gamma(r+\epsilon)-\gamma(r)|)$
$=\mu(r)((r+\epsilon)^{2}-r^{2})/2+o((\mu(r)+1)|\gamma(r+\epsilon)-\gamma(r)|)$
$=\mu(r)(r\epsilon+\epsilon^{2}/2)+o((\mu(r)+1)|\gamma(r+\epsilon)-\gamma(r)|)$
$=r\mu(r)\epsilon+o(\epsilon).$
3.3 Calculus rules
While the definition of flatness is quite general, determining flat minima directly from the definition is not always easy. We thus develop some simple calculus rules. We begin with the nonsmooth case.
Lemma 2.
If $f:\mathbb{R}^{n}\to\mathbb{R}$ is Lipschitz continuous and regular near $x\in\mathbb{R}^{n}$, then
$\mathrm{lip}f(x)=\max_{|u|=1}~\lim_{\tau\searrow 0}\frac{|f(x+\tau u)-f(x)|}{\tau}=\limsup_{\scriptsize\begin{array}[]{c}y\rightarrow x\\ y\neq x\end{array}}\frac{|f(y)-f(x)|}{|y-x|}.$
Proof.
Since $f$ is Lipschitz continuous near $x$, by [36, Theorem 9.13] $\partial f(x)$ is nonempty and compact, and $\mathrm{lip}f(x)=\max\{|v|:v\in\partial f(x)\}$. Since $f$ is Lipschitz continuous and regular near $x$, by [36, Theorem 9.16] we have $\lim_{\tau\searrow 0}|f(x+\tau u)-f(x)|/\tau=\max\{\langle u,v\rangle:v\in\partial f(x)\}$. Thus
$\max_{|u|=1}\lim_{\tau\searrow 0}\frac{|f(x+\tau u)-f(x)|}{\tau}=\max_{|u|=1}\max_{v\in\partial f(x)}\langle u,v\rangle=\max_{v\in\partial f(x)}\max_{|u|=1}\langle u,v\rangle=\max_{v\in\partial f(x)}|v|=\mathrm{lip}f(x).$
Proposition 3.
If $f:\mathbb{R}^{n}\to\mathbb{R}$ is regular near $x\in\mathbb{R}^{n}$ and $\mathrm{lip}f(x)<\infty$, then $\accentset{\circ}{f}(x,r)=\mathrm{lip}f(x)r+o(r)$.
Proof.
On the one hand,
$\frac{\accentset{\circ}{f}(x,r)}{r}$ $=\sup_{y\in B_{r}(x)\setminus\{x\}}\frac{|f(y)-f(x)|}{r}\leq\sup_{y\in B_{r}(x)\setminus\{x\}}\frac{|f(y)-f(x)|}{|y-x|}$
$\leq\sup_{y,z\in B_{r}(x),y\neq z}\frac{|f(y)-f(z)|}{|y-z|}=\mathrm{lip}f(x)+o(1)$
for $r>0$ near 0. On the other hand, let $\overline{u}\in\arg\max_{|u|=1}\lim_{\tau\searrow 0}|f(x+\tau u)-f(x)|/\tau$. By Lemma˜2, we have $\accentset{\circ}{f}(x,r)\geq|f(x+r\overline{u})-f(x)|=\mathrm{lip}f(x)r+o(r)$ for $r\geq 0$ near 0. ∎
Without regularity, Proposition˜3 may fail.
Example 1.
If $f(x,y)=\min\{x^{2},|y|\}$, then $\mathrm{lip}f(0,0)=1$ and yet $\accentset{\circ}{f}((0,0),r)=(\sqrt{1+4r^{2}}-1)/2=r^{2}+o(r^{2})\neq r+o(r)$.
We next turn to the differentiable case.
Proposition 4.
If $f:\mathbb{R}^{n}\to\mathbb{R}$ is $D^{k}$ at $x\in\mathbb{R}^{n}$ and $f^{(i)}(x)=0$ for all $i\in\llbracket 1,k-1\rrbracket$, then $\accentset{\circ}{f}(x,r)=\|f^{(k)}(x)\|r^{k}/k!+o(r^{k})$.
Proof.
Since $f(y)-f(x)=f^{(k)}(x)(y-x)^{k}/k!+o(|y-x|^{k})$, we have
$|f(y)-f(x)|=\frac{1}{k!}|f^{(k)}(x)(y-x)^{k}|+o(|y-x|^{k})\leq\frac{1}{k!}\|f^{(k)}(x)\||y-x|^{k}(1+o(1)).$
On the one hand,
$\frac{\accentset{\circ}{f}(x,r)}{r^{k}}=\sup_{y\in B_{r}(x)\setminus\{x\}}\frac{|f(y)-f(x)|}{r^{k}}\leq\sup_{y\in B_{r}(x)\setminus\{x\}}\frac{|f(y)-f(x)|}{|y-x|^{k}}\leq\frac{1}{k!}\|f^{(k)}(x)\|+o(1).$
On the other hand, with $\overline{u}\in\arg\max\{|f^{(k)}(x)u^{k}|:|u|=1\}$, we have $f(x+r\overline{u})=f(x)+r^{k}f^{(k)}(x)\overline{u}^{k}+o(r^{k})$ for $r\geq 0$ near 0 and thus $\accentset{\circ}{f}(x,r)\geq|f(x+r\overline{u})-f(x)|=r^{k}\|f^{(k)}(x)\|(1+o(1))/k!$. ∎
Proposition˜3 and Proposition˜4 admit the following corollaries.
Corollary 1.
Suppose $f:\mathbb{R}^{n}\to\mathbb{R}$ is regular near $\overline{x}\in\mathbb{R}^{n}$ and $\mathrm{lip}f(x)<\infty$ for all $x\in[f=f(\overline{x})]$ near $\overline{x}$.
  1. (i)
    If $\overline{x}$ is flat, then $\overline{x}$ is a local minimum of $\mathrm{lip}f+\delta_{[f=f(\overline{x})]}$.
  2. (ii)
    If $\overline{x}$ is a strict local minimum of $\mathrm{lip}f+\delta_{[f=f(\overline{x})]}$, then $\overline{x}$ is strictly flat.
Proof.
(i) There exists a neighborhood $U$ of $\overline{x}$ in $f=f(\overline{x})]$ such that
$\forall x\in U,~\exists\overline{r}>0:~\forall r\in(0,\overline{r}],~~\accentset{\circ}{f}(\overline{x},r)\leq\accentset{\circ}{f}(x,r),$
and in particular $\accentset{\circ}{f}(\overline{x},r)/r\leq\accentset{\circ}{f}(x,r)/r$ for all $r\in(0,\overline{r}]$. Passing to the limit as $r\searrow 0$ yields $\mathrm{lip}f(\overline{x})\leq\mathrm{lip}f(x)$ by Proposition˜3.
(ii) There exists a neighborhood $U$ of $\overline{x}$ in $f=f(\overline{x})]$ such that $\mathrm{lip}f(\overline{x})<\mathrm{lip}f(x)$ for all $x\in U\setminus\{\overline{x}\}$. By Proposition˜3, $\accentset{\circ}{f}(x,r)=\mathrm{lip}f(x)r+o(r)$ for all $x\in U$ after possibly reducing $U$. Thus, for all $x\in U\setminus\{\overline{x}\}$, there exists $\overline{r}>0$ such that $\accentset{\circ}{f}(\overline{x},r)<\accentset{\circ}{f}(x,r)$ for all $r\in(0,\overline{r}]$. ∎
Corollary 2.
Suppose $f:\mathbb{R}^{n}\to\mathbb{R}$ is $D^{k}$ near $\overline{x}\in\mathbb{R}^{n}$ and for all $x\in[f=f(\overline{x})]$ near $\overline{x}$, $f^{(i)}(x)=0$ for all $i\in\llbracket 1,k-1\rrbracket$.
  1. (i)
    If $\overline{x}$ is flat, then $\overline{x}$ is a local minimum of $\|f^{(k)}\|+\delta_{[f=f(\overline{x})]}$.
  2. (ii)
    If $\overline{x}$ is a strict local minimum of $\|f^{(k)}\|+\delta_{[f=f(\overline{x})]}$, then $\overline{x}$ is strictly flat.
Proof.
One argues as for Corollary˜2 using Proposition˜4. ∎
When $\|f^{(i)}\|$ are merely constant on $f=f(\overline{x})]$, it is neither necessary nor sufficient for $\overline{x}$ to be flat that it be a local or strict local minimum of $\|f^{(k)}\|+\delta_{[f=f(\overline{x})]}$.
Example 2.
All the minima of $f(x)=x_{1}^{2}+x_{2}^{4}(1+x_{3}^{2})$ are flat even though $\|f^{(1)}(0,0,x_{3})\|=0$, $\|f^{(2)}(0,0,x_{3})\|=2$, $\|f^{(3)}(0,0,x_{3})\|=0$, and $\|f^{(4)}(0,0,x_{3})\|=24(1+x_{3}^{2})$.
Proof.
For all $x_{3}\in\mathbb{R}$ and $r\in[0,1/(1+x_{3}^{2})]$, $\accentset{\circ}{f}(0,0,x_{3},r)=r^{2}$. ∎
Example 3.
$(0,0,1)$ is the unique flat minimum of $f(x)=x_{1}^{2}+x_{2}^{4}(1+x_{3}^{2})+x_{1}^{6}(1+(x_{3}-1)^{2})$ yet $\|f^{(1)}(0,0,x_{3})\|=0$, $\|f^{(2)}(0,0,x_{3})\|=2$, $\|f^{(3)}(0,0,x_{3})\|=0$, and $\|f^{(4)}(0,0,x_{3})\|=24(1+x_{3}^{2})$.
Proof.
For all $x_{3}\in\mathbb{R}$, there exists $\overline{r}>0$ such that $\accentset{\circ}{f}(0,0,x_{3},r)=r^{2}+(1+(x_{3}-1)^{2})r^{6}$ for all $r\in[0,\overline{r}]$. ∎
Notwithstanding, there is one situation where local optimality can be sufficient for flatness.
Fact 9.
If $f:\mathbb{R}^{n}\to\mathbb{R}$ is invariant under the natural action of a Lie subgroup $G$ of $\mathrm{O}(n)$, then $\accentset{\circ}{f}(x,r)=\accentset{\circ}{f}(gx,r)$ for $(x,r,g)\in\mathbb{R}^{n}\times\mathbb{R}_{+}\times G$.
Proof.
$\accentset{\circ}{f}(x,r)=\sup_{y\in\overline{B}_{r}(x)}|f(y)-f(x)|=\sup_{gy\in\overline{B}_{r}(gx)}|f(gy)-f(gx)|=\sup_{z\in\overline{B}_{r}(gx)}|f(z)-f(gx)|=\accentset{\circ}{f}(gx,r)$. ∎
Corollary 3.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz regular and invariant under the natural action of a Lie subgroup $G$ of $\mathrm{O}(n)$. Let $\overline{x}\in\mathbb{R}^{n}$. If $G\overline{x}$ is a strict local minimum of $\mathrm{lip}f+\delta_{[f=f(\overline{x})]}$, then every point in $G\overline{x}$ is flat.
Proof.
˜9 implies that $x\preceq y$ and $y\preceq x$ for all $x,y\in G\overline{x}$. By Proposition˜3, $\accentset{\circ}{f}(x,r)=\mathrm{lip}f(x)r+o(r)$ for all $x\in\mathbb{R}^{n}$. There exists a neighborhood $U$ of $G\overline{x}$ in $f=f(\overline{x})]$ such that $\mathrm{lip}f(x)<\mathrm{lip}f(y)$ for all $x\in G\overline{x}$ and $y\in U\setminus G\overline{x}$. Hence, for any such $x$ and $y$, there exists $\overline{r}>0$ such that $\accentset{\circ}{f}(x,r)<\accentset{\circ}{f}(y,r)$ for all $r\in(0,\overline{r}]$. ∎
An analogous result holds in the differentiable case.
Corollary 4.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be $D^{k}$ and invariant under the natural action of a Lie subgroup $G$ of $\mathrm{O}(n)$. Let $\overline{x}\in\mathbb{R}^{n}$. Suppose that for all $x\in[f=f(\overline{x})]$, $f^{(i)}(x)=0$ for all $i\in\llbracket 1,k-1\rrbracket$. If $G\overline{x}$ is a strict local minimum of $\|f^{(k)}\|+\delta_{[f=f(\overline{x})]}$, then every point in $G\overline{x}$ is flat.
When applying the calculus rules, a composite structure can help.
Fact 10.
[36, Theorem 10.6] Let $f=g\circ F$ where $g:\mathbb{R}^{m}\to\mathbb{R}$ is locally Lipschitz regular and $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is $C^{1}$. Then $f$ is regular and
$\forall x\in\mathbb{R}^{n},~~~\partial f(x)=F^{\prime}(x)^{*}\partial g(F(x)).$
In particular, if $F(x)=0$ and $g=|\cdot|_{1}$, then $\mathrm{lip}f(x)=\|F^{\prime}(x)^{*}\|_{\infty,2}$.
Fact 11.
Let $f=g\circ F$ where $g:\mathbb{R}^{m}\to\mathbb{R}$ and $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$ are $D^{k}$, and $g^{(i)}(0)=0$ for all $i\in\llbracket 1,k-1\rrbracket$. Let $x\in\mathbb{R}^{n}$ be such that $F(x)=0$. Then
$\forall i\in\llbracket 1,k-1\rrbracket,~f^{(i)}(x)=0~~\land~~f^{(k)}(x)v^{k}=g^{(k)}(x)(F^{\prime}(x)v)^{k}.$
In particular, if $g=|\cdot|^{k}/k$, then $\|f^{(k)}(x)\|=\|F^{\prime}(x)\|_{2}^{k}$.
Sometimes the composite structure helps with subdifferentiation, but not with determining flat minima.
Remark 1.
Let $f=g\circ F$ where $g:\mathbb{R}^{m}\to\mathbb{R}$ is $C^{1}$ such that $\nabla g(0)=0$ and $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is locally Lipschitz. The Jacobian chain rule [11, Theorem 2.6.6] yields $\overline{\partial}f(x)=\overline{\partial}F(x)^{T}\nabla g(F(x))$. Hence, if $F(x)=0$, then $\overline{\partial}f(x)=\{0\}$. Since the Clarke subdifferential is a singleton, by [36, Theorem 9.18] $f$ is strictly differentiable [36, Definition 9.17] at $x$, $\nabla f(x)=0$, and $\mathrm{lip}f(x)=0$.
Two additional rules are useful in the presence of symmetries.
Fact 12.
Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be lower semicontinuous and invariant under the natural action of a Lie subgroup of $\mathrm{GL}(n,\mathbb{R})$. For all $\overline{x}\in\mathrm{dom}f$ and $g\in G$, $\partial f(g^{-1}\overline{x})=g^{T}\partial f(\overline{x})$.
Proof.
Applying the chain rule [36, Exercise 10.7] to $f(x)=f(gx)$ at $g^{-1}\overline{x}$ for any $g\in G\subseteq\mathrm{GL}(n,\mathbb{R})$ yields the result. ∎
Fact 13.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be $D^{k}$ and invariant under the natural action of a Lie subgroup of $\mathrm{GL}(n,\mathbb{R})$. For all $\overline{x},v_{1},\ldots,v_{k}\in\mathbb{R}^{n}$, $f^{(k)}(g^{-1}\overline{x})(v_{1},\ldots,v_{k})=f^{(k)}(\overline{x})(gv_{1},\ldots,gv_{k})$.
Proof.
Differentiating $f(x)=f(gx)$ at $g^{-1}\overline{x}$ in the direction $v$ yields $f^{\prime}(g^{-1}\overline{x})(v)=f^{\prime}(\overline{x})(gv)$. Assuming that $f^{(i)}(g^{-1}\overline{x})(v_{1},\ldots,v_{i})=f^{(i)}(\overline{x})(gv_{1},\ldots,gv_{i})$, deriving with respect to $\overline{x}$ in the direction $gv_{i+1}$ yields $f^{(i+1)}(g^{-1}\overline{x})(v_{1},\ldots,v_{i+1})=f^{(i+1)}(\overline{x})(gv_{1},\ldots,gv_{i+1})$. ∎
4 Flatness and conservation
Having established basic properties of flat minima, we now show how conserved quantities in subgradient dynamics provide a useful tool for analyzing them in more detail.
4.1 Flattening trajectories
A curve is a function $x:I\to\mathbb{R}$ where $I$ is an interval of $\mathbb{R}$. We refer to trajectories as curves that solve an ODE. We say that a curve $x:I\to\mathbb{R}$ is flattening, or flattens over time, if $x(t)\prec x(s)$ for all $s,t\in I$ such that $s<t$. A curve $x:I\to\mathbb{R}$ sharpens over time if $I\ni t\mapsto x(-t)\in\mathbb{R}^{n}$ flattens over time. We seek to construct such curves using the following assumption.
Assumption 2.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz definable and $c:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be $C^{4}$ on a bounded convex open set $U\subseteq\mathbb{R}^{n}$ such that
$\exists\omega>0:~\forall x\in U,~\forall v\in\overline{\partial}f(x),~~~\langle\nabla c(x),v\rangle=0~~\text{and}~~~\langle\nabla^{2}c(x)v,v\rangle\leq-\omega|v|^{2}.$
The first-order condition in ˜2 induces conservation laws.
Proposition 5.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz definable and $c:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be $D^{1}$ on an open set $U\subseteq\mathbb{R}^{n}$ such that $\langle\nabla c(x),v\rangle=0$ for all $x\in U$ and $v\in\overline{\partial}f(x)$.
  1. (i)
    If $x:I\to U$ is a solution to $\dot{x}\in-\overline{\partial}f(x)$ almost everywhere, then $c$ is conserved.
  2. (ii)
    If $x:I\to U$ is a solution to $\dot{x}=\nabla c(x)$, then $f$ is conserved.
Proof.
(i) Since $x(\cdot)$ is absolutely continuous, it is differentiable almost everywhere. Thus, for almost every $t\in I$, $(c\circ x)^{\prime}(t)=\langle\nabla c(x),x^{\prime}(t)\rangle=0$ and $I\ni t\to c(x(t))$ is constant.
(ii) Since $f$ is definable and Lipschitz continuous, by [9, Corollary 2, Proposition 2] $f$ is path differentiable. Thus, for almost every $t\in I$,
$\forall v\in\overline{\partial}f(x(t)),~~~(f\circ x)^{\prime}(t)=\langle v,x^{\prime}(t)\rangle=\langle v,\nabla c(x(t))\rangle=0$
and $I\ni t\to f(x(t))$ is constant. ∎
Under ˜2, gradient trajectories of the conserved quantity $c$ sharpen over time. This is because the level sets are contracting. To see why, we begin with a simple lemma.
Lemma 3.
If $c:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is $C^{4}$ on a bounded convex open set $U\subseteq\mathbb{R}^{n}$, then there exists $M>0$ such that
$\forall x,y\in U,~~~\left|\langle\nabla c(x)-\nabla c(y),x-y\rangle-\langle\nabla^{2}c(x)(x-y),x-y\rangle\right|\leq M|x-y|^{3}.$
Proof.
Consider the function $g:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ defined by $g(x,y)=\langle\nabla c(x)-\nabla c(y),x-y\rangle$. Since $c$ is $C^{4}$ on $U$, $g$ is $C^{3}$ on $U$. By [34, Lemma 1.2.4], there exists $M>0$ such that for all $x,y\in U$, $u\in U+\{-x\}$, and $v\in U+\{-y\}$ we have
$\left|g(x+u,y+v)-g(x,y)-\langle\nabla g(x,y),(u,v)\rangle-\langle\nabla^{2}g(x,y)(u,v),(u,v)\rangle/2\right|\leq M(|u|^{3}+|v|^{3}).$
In particular,
$\left|g(x,x+v)-g(x,x)-\langle\nabla g(x,x),(0,v)\rangle-\langle\nabla^{2}g(x,x)(0,v),(0,v)\rangle/2\right|\leq M|v|^{3}.$
Fix $x\in U$ and consider the function $h:\mathbb{R}^{n}\to\mathbb{R}$ be defined by $h(y)=g(x,y)=c^{\prime}(x)(x-y)-c^{\prime}(y)(x-y)$. We have
$\forall u\in\mathbb{R}^{n},~~~h^{\prime}(y)u=-c^{\prime}(x)u-c^{\prime\prime}(y)(x-y,u)+c^{\prime}(y)u,$
$\forall u,v\in\mathbb{R}^{n},~~~h^{\prime\prime}(y)(u,v)=-c^{\prime\prime\prime}(y)(x-y,u,v)+c^{\prime\prime}(y)(v,u)+c^{\prime\prime}(y)(u,v).$
In particular, $\nabla h(x)=0$ and $\nabla^{2}h(x)=\nabla^{2}c(x)$. For all $x\in U$ and $v\in U+\{-x\}$, it follows that $|\langle\nabla c(x+v)-\nabla c(x),v\rangle-\langle\nabla^{2}c(x)v,v\rangle|\leq M|v|^{3}$, yielding the desired inequality. ∎
The following result is one of the main findings of this paper.
Theorem 1.
Under ˜2, for all $x_{0}\in U$ near which $f$ is nonconstant, there exists a local solution $x:[0,\overline{t})\to\mathbb{R}^{n}$ to
$\left\{\begin{array}[]{ccc}\dot{x}&=&\nabla c(x)\\ x(0)&=&x_{0}\end{array}\right$.
such that the two equivalent conditions hold:
  1. (i)
    $\exists\overline{r}>0,~\forall r\in[0,\overline{r}),~\forall t\in[0,\overline{t}),~~~\accentset{\circ}{f}(x_{0},r)\leq\accentset{\circ}{f}(x(t),e^{-\omega t/2}r)$,
  2. (ii)
    $\exists\overline{\ell}>0,~\forall\ell\in[0,\overline{\ell}),~\forall t\in[0,\overline{t}),~~~\overline{f}(x(t),\ell)\leq e^{-\omega t/2}\overline{f}(x_{0},\ell)$.
Proof.
Since $\nabla c$ is Lipschitz continuous on $U$, by [27, Theorem D.1], there exist $\overline{t},\overline{r}>0$ such that the ODE
$\left\{\begin{array}[]{ccc}\dot{z}&=&\nabla c(z)\\ z(0)&=&z_{0}\end{array}\right$.
admits a unique $C^{1}$ solution $z:[0,\overline{t})\to U$ for all $z_{0}\in B_{\overline{r}}(x_{0})$. Since $\nabla c$ is $C^{2}$, by [27, Theorem D.1], the semiflow $\theta:[0,\overline{t})\times B_{\overline{r}}(x_{0})\to\mathbb{R}^{n}$ defined by $\theta(t,z_{0})=z(t)$ is $C^{2}$. By [34, Lemma 1.2.4], there exists $C>0$ such that for all $t,\widetilde{t}\in[0,\overline{t})$ and $z_{0},\widetilde{z}_{0}\in B_{\overline{r}}(x_{0})$, we have
$|\theta(t,z_{0})-\theta(\widetilde{t},\widetilde{z}_{0})-\theta^{\prime}(\widetilde{t},\widetilde{z}_{0})(t-\widetilde{t},z_{0}-\widetilde{z}_{0})|\leq C\left[(t-\widetilde{t})^{2}+|z_{0}-\widetilde{z}_{0}|^{2}\right].$ (1)
Since $f$ is locally Lipschitz definable and nonconstant near $x_{0}$, by Lemma˜1 there exist definable curves $\gamma:[0,\overline{r}]\to\mathbb{R}^{n}$, $\lambda:[0,\overline{r}]\to\mathbb{R}$, and $v:[0,\overline{r}]\to\mathbb{R}^{n}$ such that $\gamma$ is $C^{1}$ on $0,\overline{r}]$ and
$\forall r\in[0,\overline{r}],~~~\gamma(r)=x_{0}+\lambda(r)v(r)\in\arg\max_{\overline{B}_{r}(x_{0})}|f-f(x_{0})|~~~\text{and}~~~v(r)\in\overline{\partial}f(\gamma(r)),$
after possibly reducing $\overline{r}$. As $c$ is $C^{4}$ on $U$, there exists $L>0$ such that
$\forall x,y\in U,~~~|\nabla^{2}c(x)-\nabla^{2}c(y)|\leq L|x-y|.$
Let $M>0$ be given by Lemma˜3. For all $(t,r)\in[0,\overline{t})\times[0,\overline{r})$, let $x_{r}(t)=\theta(t,\gamma(r))$ and $x(t)=x_{0}(t)=\theta(t,\overline{x})$. As $\gamma$ is continuous, so is the composition $0,\overline{t})\times[0,\overline{r})\ni(t,r)\to x_{r}(t)\in\mathbb{R}^{n}$. In particular, $|x_{r}(t)-x_{0}|\leq\omega/(8\max\{L,M\})$ for all $t\in[0,\overline{t})$ and $r\in[0,\overline{r})$ after possibly reducing $\overline{t}$ and $\overline{r}$. It follows that $|x_{r}(t)-\gamma(r)|\leq|x_{r}(t)-x_{0}|+|x_{0}-\gamma(r)|\leq\omega/(8L)+\omega/(8L)=\omega/(4L)$ and $|x_{r}(t)-x(t)|\leq|x_{r}(t)-x_{0}|+|x_{0}-x_{0}(t)|\leq\omega/(4M)$.
The main idea of the proof is now as follows. Using Proposition˜5, we have
$\accentset{\circ}{f}(x_{0},r)=\max_{\overline{B}_{r}(x_{0})}|f-f(x_{0})|=|f(x_{r}(0))-f(x(0))|=|f(x_{r}(t))-f(x(t))|\leq\accentset{\circ}{f}(x(t),e^{-\omega t/2}r)$
where the last inequality is due to $|x_{r}(t)-x(t)|\leq e^{-\omega t/2}|x_{r}(0)-x(0)|=e^{-\omega t/2}r$. The step we need to justify is the contraction. To that avail, let $y_{r}:[0,\overline{t})\to\mathbb{R}^{n}$ be defined by $y_{r}(t)=x_{r}(t)-x(t)$. Initially $y_{r}(0)=x_{r}(0)-x(0)=\gamma(r)-x_{0}=\lambda(r)v(r)\neq 0$ for all $r\in(0,\overline{r})$. Since $v(r)\in\overline{\partial}f(\gamma(r))$, we have $\langle\nabla^{2}c(\gamma(r))v(r),v(r)\rangle\leq-\omega|v(r)|^{2}$. Together with $\lambda(r)\neq 0$, we find $\langle\nabla^{2}c(\gamma(r))\lambda(r)v(r),\lambda(r)v(r)\rangle\leq-\omega|\lambda(r)v(r)|^{2}$. In other words,
$\forall r\in(0,\overline{r}),~~~\left\langle\nabla^{2}c(\gamma(r))\frac{y_{r}(0)}{|y_{r}(0)|},\frac{y_{r}(0)}{|y_{r}(0)|}\right\rangle\leq-\omega.$ (2)
By (1), for all $(t,r)\in[0,\overline{t})\times(0,\overline{r})$ we have $\left|\theta(t,\gamma(r))-\theta(t,x_{0})-\theta^{\prime}(t,x_{0})(0,\gamma(r)-x_{0})\right|\leq C|\gamma(r)-x_{0}|^{2}$ and thus
$\left|\frac{\theta(t,\gamma(r))-\theta(t,x_{0})}{|\gamma(r)-x_{0}|}-\frac{\partial\theta}{\partial z_{0}}(t,x_{0})\frac{\gamma(r)-x_{0}}{|\gamma(r)-x_{0}|}\right|\leq C|\gamma(r)-x_{0}|.$ (3)
Each entry of $(\gamma-x_{0})/|\gamma-x_{0}|$ is definable and bounded. By the monotonicity theorem [40, 4.1], it is monotone near 0 and thus convergent. Thus there exists $u\in S^{n-1}$ such that $(\gamma(r)-x_{0})/|\gamma(r)-x_{0}|\to u$ as $r\searrow 0$. Since $\theta$ is $C^{1}$, it follows that
$\frac{\partial\theta}{\partial z_{0}}(t,x_{0})\frac{\gamma(r)-x_{0}}{|\gamma(r)-x_{0}|}\xrightarrow[(t,r)\searrow(0,0)]{}\frac{\partial\theta}{\partial z_{0}}(0,x_{0})u=u,$
where the equality holds because $\theta(0,z_{0})=z_{0}$ for all $z_{0}\in B_{\overline{r}}(x_{0})$. From (3), we deduce
$\frac{y_{r}(t)}{|\gamma(r)-x_{0}|}=\frac{x_{r}(t)-x(t)}{|\gamma(r)-x_{0}|}=\frac{\theta(t,\gamma(r))-\theta(t,x_{0})}{|\gamma(r)-x_{0}|}\xrightarrow[(t,r)\searrow(0,0)]{}u$
and
$\frac{y_{r}(t)}{|y_{r}(t)|}=\frac{y_{r}(t)/|\gamma(r)-x_{0}|}{|y_{r}(t)|/|\gamma(r)-x_{0}|}\xrightarrow[(t,r)\searrow(0,0)]{}u.$
Since $\gamma$ and $\nabla c$ are continuous, passing to the limit in (2) yields $\langle\nabla^{2}c(x_{0})u,u\rangle\leq-\omega$. As a result,
$\forall t\in[0,\overline{t}),~\forall r\in(0,\overline{r}),~~~\left\langle\nabla^{2}c(\gamma(r))\frac{y_{r}(t)}{|y_{r}(t)|},\frac{y_{r}(t)}{|y_{r}(t)|}\right\rangle\leq-\frac{\omega}{2}$
after possibly reducing $\overline{t}$ and $\overline{r}$. Now
$\frac{d|y_{r}|^{2}}{dt}$ $=2\langle\dot{y}_{r},y_{r}\rangle$
$=2\langle\dot{x}_{r}-\dot{x},y_{r}\rangle$
$=2\langle\nabla c(x_{r})-\nabla c(x),y_{r}\rangle$
$\leq 2\langle\nabla^{2}c(x_{r})y_{r},y_{r}\rangle+2M|y_{r}|^{3}$
$=2\langle\nabla^{2}c(\gamma(r))y_{r},y_{r}\rangle+2\langle[\nabla^{2}c(x_{r})-\nabla^{2}c(\gamma(r))]y_{r},y_{r}\rangle+2M|y_{r}|^{3}$
$\leq 2\langle\nabla^{2}c(\gamma(r))y_{r},y_{r}\rangle+2|\nabla^{2}c(x_{r})-\nabla^{2}c(\gamma(r))||y_{r}|^{2}+2M|y_{r}|^{3}$
$\leq 2\langle\nabla^{2}c(\gamma(r))y_{r},y_{r}\rangle+2L|x_{r}-\gamma(r)||y_{r}|^{2}+2M|y_{r}|^{3}$
$\leq-\omega|y_{r}|^{2}+2L|x_{r}-\gamma(r)||y_{r}|^{2}+2M|y_{r}|^{3}$
$=-(2\omega-2L|x_{r}-\gamma(r)|-2M|y_{r}|)|y_{r}|^{2}$
$\leq-(2\omega-\omega/2-\omega/2)|y_{r}|^{2}$
$\leq-\omega|y_{r}|^{2}.$
The ODE comparison theorem [27, Theorem D.2] implies that $|y_{r}(t)|^{2}\leq e^{-\omega t}|y_{r}(0)|^{2}$, namely, $|x_{r}(t)-x(t)|\leq e^{-\omega t/2}|x_{r}(0)-x(0)|$ for all $t\in[0,\overline{t})$ and $r\in[0,\overline{r})$. As for the equivalence, it is obtained via Lemma˜1 and the change of variables $\ell=\accentset{\circ}{f}(x_{0},r)$: $\accentset{\circ}{f}(x_{0},r)\leq\accentset{\circ}{f}(x(t),e^{-\omega t/2}r)$ iff $\overline{f}(x(t),\accentset{\circ}{f}(x_{0},r))\leq\overline{f}(x(t),\accentset{\circ}{f}(x(t),e^{-\omega t/2}r))$ iff $\overline{f}(x(t),\ell)\leq e^{-\omega t/2}\overline{f}(x_{0},\ell)$. ∎
Due to the local nature of Theorem˜1, it does not guarantee that reversing time yields flattening trajectories from a given initial point. In order to do so, we will add some assumptions and rely on the following fact.
Fact 14.
Let $g:(a,b)\to\mathbb{R}$ be continuous such that for all $s\in(a,b)\subseteq\mathbb{R}$, there exists $s^{\prime}\in(s,b)$ such that $g(t)\geq e^{t-s}g(s)$ for all $t\in(s,s^{\prime})$. Then $g(t)\geq e^{t-s}g(s)$ for all $s,t\in(a,b)$.
Proof.
Let $a<\overline{s}<\overline{t}<b$ and suppose $t_{0}=\inf\{t\in[\overline{s},\overline{t}]:g(t)<e^{t-\overline{s}}g(\overline{s})\}<\infty$. By assumption $t_{0}>\overline{s}$. By definition of $t_{0}$, we have $g(t)\geq e^{t-\overline{s}}g(\overline{s})$ for all $t\in[\overline{s},t_{0})$. Since $g$ is continuous, $g(t_{0})\geq e^{t_{0}-\overline{s}}g(\overline{s})$. By assumption, there exists $t_{1}\in(t_{0},b)$ such that $g(t)\geq e^{t-t_{0}}g(t_{0})$ for all $t\in(t_{0},t_{1})$, and so $g(t)\geq e^{t-t_{0}}e^{t_{0}-\overline{s}}g(\overline{s})=e^{t-\overline{s}}g(\overline{s})$. Hence $g(t)\geq e^{t-t_{0}}g(t_{0})$ for all $t\in[\overline{s},t_{1}]$, contradicting the optimality of $t_{0}$. ∎
Corollary 5.
Under ˜2, let $x:I\to U$ be such that $\dot{x}=-\nabla c(x)$ on a compact interval $I$ of $\mathbb{R}$ containing $t_{0}$. Let $s,t\in I$ be such that $s<t$.
  1. (i)
    If $f$ is locally Lipschitz regular, then $\mathrm{lip}f(x(t))\leq e^{-\omega(t-s)/2}\mathrm{lip}f(x(s))$.
  2. (ii)
    If $c$ is $C^{k+1}$, $f$ is $D^{k}$, and $f^{(i)}(x(t_{0}))=0$ for all $i\in\llbracket 1,k-1\rrbracket$, then $\|f^{(k)}(x(t))\|\leq e^{-k\omega(t-s)/2}\|f^{(k)}(x(s))\|$.
If $\mathrm{lip}f(x(t_{0}))>0$ in (i) or $\|f^{(k)}(x(t_{0}))\|>0$ in (ii), then $x(t)\prec x(s)$.
Proof.
(i) Let $t\in I$. By Theorem˜1 and Proposition˜3, $\mathrm{lip}f(x(t))+o(1)=\accentset{\circ}{f}(x(t),r)/r\leq\accentset{\circ}{f}(x(s),e^{-\omega(t-s)/2}r)/r=e^{-\omega(t-s)/2}\mathrm{lip}f(x(s))+o(1)$ for all $s\in I\cap(-\infty,t]$ sufficiently close to $t$. ˜14 enables one to extend it to any $s\in I\cap(-\infty,t]$. (ii) Assuming the existence of a solution $x:I\to U$ to the ODE passing through $x_{0}=x(t_{0})$ at time $t_{0}$ implies the existence of solutions on an open interval $J\supset I$ for initial points in a neighborhood $U_{0}$ of $x_{0}$. This holds because $\nabla c$ is Lipschitz continuous on $U$, which also implies uniqueness of solutions (using the ODE comparison theorem [27, Theorem D.2]). Thus the flow $\theta:J\times U_{0}\to U$ is well defined and $C^{k}$. For any $t\in J$, $\theta_{t}=\theta(t,\cdot)$ defines a diffeomorphism from $U_{0}$ to $\theta_{t}(U_{0})$ [27, p. 209], which is a neighborhood of $x(t)$. Conservation implies that $f=f\circ\theta_{t}$. By the chain rule, $(f\circ\theta_{t})^{\prime}(x)=f^{\prime}(\theta_{t}(x))\circ\theta_{t}^{\prime}(x)$. In particular, $0=f^{\prime}(x_{0})(v)=(f\circ\theta_{t})^{\prime}(x_{0})(v)=f^{\prime}(\theta_{t}(x_{0}))(\theta_{t}^{\prime}(x_{0})v)$ for all $v\in\mathbb{R}^{n}$, so that $f^{\prime}(\theta_{t}(x_{0}))=0$. By induction, $(f\circ\theta_{t})^{(i)}(x_{0})v^{i}=f^{(i)}(\theta_{t}(x_{0}))(\theta_{t}^{\prime}(x_{0})v)^{i}$ for all $i\in\llbracket 1,k\rrbracket$ and thus $f^{(i)}(x(t))=0$ for all $i\in\llbracket 1,k-1\rrbracket$. It follows that $\|f^{(k)}(x(t))\|/k!+o(1)=\accentset{\circ}{f}(x(t),r)/r^{k}\leq\accentset{\circ}{f}(x(s),e^{-\omega(t-s)/2}r)/r^{k}=e^{-k\omega(t-s)/2}\|f^{(k)}(x(s))\|/k!+o(1)$ for all $s\in I\cap(-\infty,t]$ sufficiently close to $t$ by Theorem˜1 and Proposition˜4. One again concludes by ˜14. ∎
Remark 2.
One can actually weaken the equality $\langle\nabla c(x),v\rangle=0$ for all $v\in\overline{\partial}f(x)$ in ˜2 to an inequality $\langle\nabla c(x),v\rangle\geq 0$. In exchange, one must assume that $x(\cdot)$ is defined on $0,\overline{t}]$ (at no cost) and $x(\overline{t})\in\arg\min\{f(x):x\in U\}$ in Theorem˜1. In the proof, $f$ instead increases along the trajectories of $\nabla c$, but is still constant on $x(\cdot)$, so that $|f(x_{r}(0))-f(x(0))|\leq|f(x_{r}(t))-f(x(t))|$.
Similarly, one must assume that $x(t_{0})\in\arg\min\{f(x):x\in U\}$ in Corollary˜5. In the proof, instead of the chain rule, one uses Taylor expansions: $f(y)=f(x_{0})+\langle\nabla f(x_{0}),y-x_{0}\rangle+o(|y-x_{0}|)$ and $f(\theta_{t}(y))=f(\theta_{t}(x_{0}))+\langle\nabla f(\theta_{t}(x_{0})),\theta_{t}(y)-\theta_{t}(x_{0})\rangle+o(|\theta_{t}(y)-\theta_{t}(x_{0})|)$. Since $f(x_{0})=f(\theta_{t}(x_{0}))\leq f(\theta_{t}(y))\leq f(y)$, this yields $\langle\nabla f(\theta_{t}(x_{0})),\theta_{t}(y)-\theta_{t}(x_{0})\rangle+o(|\theta_{t}(y)-\theta_{t}(x_{0})|)=o(|y-x_{0}|)$ and so $\langle\nabla f(\theta_{t}(x_{0})),\theta_{t}^{\prime}(x_{0})v\rangle=0$ for all $v\in\mathbb{R}^{n}$, namely $\nabla f(x(t))=0$. The rest follows by induction.
With these new assumptions, it is sufficient to satisfy the second inequality in ˜2, i.e., $\langle\nabla^{2}c(x)v,v\rangle\leq-\omega|v|^{2}$ for all $v\in\overline{\partial}f(x)$ and $x\in U$, merely for all $x\in U\setminus\arg\min\{f(x):x\in U\}$.
4.2 Linear symmetries
Linear symmetries give rise to a conservation law in subgradient dynamics, as shown in [23]. We next recall it and compute the Hessian of the conserved quantity in some directions, since it appears in ˜2. Note that a conservation law in gradient dynamics was proposed earlier in [43, Proposition 5.1]
Assumption 3.
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be locally Lipschitz and invariant under the natural action of a Lie subgroup $G$ of $\mathrm{GL}(n,\mathbb{R})$, $C(x)=P_{\mathrm{s}(\mathfrak{g})}(xx^{T})$, and $\overline{x}\in\mathbb{R}^{n}$.
Proposition 6.
Let ˜3 hold and $c(x)=\|C(x)-C(\overline{x})\|_{F}^{2}/4$. Then
$\forall x\in\mathbb{R}^{n},~\forall v\in\overline{\partial}f(x),~~~\langle\nabla c(x),v\rangle=0~~~\text{and}~~~\langle\nabla^{2}c(x)v,v\rangle=\langle C(x)-C(\overline{x}),C(v)\rangle_{F}.$
Proof.
By [23, Corollary 1], for all $x\in\mathbb{R}^{n}$, $v\in\overline{\partial}f(x)$, and $\alpha\in\mathbb{R}$, we have $C(x+\alpha v)=C(x)+\alpha^{2}C(v)$ and thus
$c(x+\alpha v)$ $=\|C(x+\alpha v)-C(\overline{x})\|_{F}^{2}/4=\|C(x)+\alpha^{2}C(v)-C(\overline{x})\|_{F}^{2}/4$
$=\|C(x)-C(\overline{x})\|_{F}^{2}/4+\langle C(x)-C(\overline{x}),C(v)\rangle_{F}\alpha^{2}/2+\alpha^{4}\|C(v)\|_{F}^{2}/4$
$=c(x)+\langle\nabla c(x),v\rangle\alpha+\langle\nabla^{2}c(x)v,v\rangle\alpha^{2}/2+o(\alpha^{3}).\qed$
Our focus is now on deriving necessary conditions for flatness using the conserved quantity. The following result can be seen as a warm up. It will be useful later.
Proposition 7.
Under ˜3, $\overline{x}\in\arg\mathrm{loc}\min\{|x|:x\in G\overline{x}\}\Longrightarrow C(\overline{x})=0$.
Proof.
Let $\gamma:(-\epsilon,\epsilon)\rightarrow G$ be a smooth curve such that $\gamma(0)=I_{n}$ where $\epsilon>0$. Since 0 is a local minimum of $(-\epsilon,\epsilon)\ni t\to|\gamma(t)\overline{x}|^{2}$, $\langle\gamma(0)\overline{x},\gamma^{\prime}(0)\overline{x}\rangle=\langle\overline{x},\gamma^{\prime}(0)\overline{x}\rangle=\langle\gamma^{\prime}(0),\overline{x}\hskip 1.42262pt\overline{x}^{T}\rangle=0$. Thus $\langle v,\overline{x}\hskip 1.42262pt\overline{x}^{T}\rangle=0$ for all $v\in\mathfrak{g}$ and $C(\overline{x})=P_{\mathrm{s}(\mathfrak{g})}(\overline{x}\hskip 1.42262pt\overline{x}^{T})=0$. ∎
We have arrived at our second main result.
Theorem 2.
Under ˜3,
$\overline{x}\in\arg\mathrm{loc}\min\{\mathrm{lip}f(x):x\in G\overline{x}\}~\Longrightarrow~\exists\overline{v}\in\arg\max\{|v|:v\in\partial f(\overline{x})\}:~C(\overline{v})=0.$
Proof.
By assumption, $I_{n}$ is a local solution to
$\inf_{g\in G}~\mathrm{lip}f(g^{-1}\overline{x})^{2}=\inf_{g\in G}\sup_{v\in\partial f(g^{-1}\overline{x})}|v|^{2}=\inf_{g\in G}\sup_{v\in\partial f(\overline{x})}|g^{T}v|^{2}=\inf_{g\in G}-\inf_{v\in\mathbb{R}^{n}}\phi(v,g)=-\sup_{g\in G}\varphi(g)$
where the second equality holds by ˜12. Also, $\phi:\mathbb{R}^{n}\times G\rightarrow\overline{\mathbb{R}}$ and $\varphi:G\rightarrow\mathbb{R}_{-}$ are defined by
$\phi(v,g)=\delta_{\partial f(\overline{x})}(v)-|g^{T}v|^{2}~~~\text{and}~~~\varphi(g)=\inf_{v\in\mathbb{R}^{n}}\phi(v,g).$
Let $\gamma:(-\epsilon,\epsilon)\rightarrow G$ be a smooth curve such that $\gamma(0)=I_{n}$ where $\epsilon>0$. Consider the function $\tau:\mathbb{R}^{n}\times\mathbb{R}\rightarrow\overline{\mathbb{R}}$ defined by
$\tau(v,t)=\phi(v,\gamma(t))+\delta_{[-\epsilon/2,\epsilon/2]}(t).$
We have
$\forall t\in[-\epsilon/2,\epsilon/2],~~~(\varphi\circ\gamma)(t)=\inf_{v\in\mathbb{R}^{n}}\tau(v,t).$
The function $\tau$ is lower semicontinuous and continuous on its compact domain $\partial f(\overline{x})\times[-\epsilon/2,\epsilon/2]$ since $f$ is Lipschitz continuous near $\overline{x}$. Thus $\varphi\circ\gamma$ is real-valued on $\epsilon/2,\epsilon/2]$. Since 0 is a local maximum of $\varphi\circ\gamma$, we have $0\in\partial(\varphi\circ\gamma)(0)$ by Fermat’s rule [36, Theorem 10.1]. As the sum of the indicator of a closed convex set and a smooth function, $\tau$ is regular [36, Example 7.28]. [36, Corollary 10.11] then implies that for all $(v,t)\in\mathrm{dom}\tau$, one has
$\partial\tau(v,t)\subseteq\begin{pmatrix}N_{\partial f(\overline{x})}(v)-2\gamma(t)\gamma(t)^{T}v\\ N_{[-\epsilon/2,\epsilon/2]}(t)-2\langle\gamma(t)^{T}v,\gamma^{\prime}(t)^{T}v\rangle\end{pmatrix}$.
Due to its bounded domain, the function $\tau(v,t)$ is level-bounded in $v$ locally uniformly in $t$ [36, 1.16 Definition]. We may thus apply [36, Theorem 10.13] on parametric subdifferentiation. For all $\overline{t}\in[-\epsilon/2,\epsilon/2]$, it yields
$\partial(\varphi\circ\gamma)(\overline{t})\subseteq\bigcup_{\overline{v}\in\arg\min\limits_{v\in\mathbb{R}^{n}}\tau(v,\overline{t})}M(\overline{v},\overline{t})~~~\text{where}~~~M(\overline{v},\overline{t})=\{y\in\mathbb{R}:(0,y)\in\partial\tau(\overline{v},\overline{t})\}.$
In particular, since $0\in\partial(\varphi\circ\gamma)(0)$, there exists
$\overline{v}\in\arg\min\{\tau(v,0):v\in\mathbb{R}^{n}\}=\arg\max\{|v|:v\in\partial f(\overline{x})\}$
such that
$(0,0)\in\partial\tau(\overline{v},0)\subseteq\begin{pmatrix}N_{\partial f(\overline{x})}(\overline{v})-2\overline{v}\\ -2\langle\overline{v},\gamma^{\prime}(0)\overline{v}\rangle\end{pmatrix}$.
Hence $\langle\overline{v},u\overline{v}\rangle=\langle\overline{v}\hskip 0.85358pt\overline{v}^{T},u\rangle=0$ for all $u\in\mathfrak{g}$, and so $C(\overline{v})=P_{\mathrm{s}(\mathfrak{g})}(\overline{v}\hskip 0.85358pt\overline{v}^{T})=0$. ∎
In Theorem˜2, one can also take a maximal Bouligand subdifferential by ˜1. The converse of Theorem˜2 do not always hold.
Example 4.
The function $f(x)=|2x_{1}^{2}-x_{2}^{2}|$ is invariant under the natural action of $G=\{A\in\mathrm{GL}(2,\mathbb{R}):A^{T}DA=D\}$ where $D=\mathrm{diag}(2,-1)$. The only flat minimum is $(0,0)$ even though $C(x)=P_{\mathrm{s}(\mathfrak{g})}(xx^{T})=0$ for all $x\in\mathbb{R}^{2}$.
Proof.
Observe that $f(x)=\langle x,Dx\rangle$. For all $A\in\mathrm{GL}(2,\mathbb{R})$, $f(Ax)=\langle Ax,DAx\rangle=\langle x,A^{T}DAx\rangle=\langle x,Dx\rangle=f(x)$. Thus $f$ is invariant under the natural action of $G$, whose Lie algebra is
$\mathfrak{g}=\left\{B\in\mathbb{R}^{2\times 2}:~B^{T}D+DB=0\right\}=\left\{\begin{pmatrix}0&2t\\ -t&0\end{pmatrix}:~t\in\mathbb{R}\right\}.$
Thus $\mathrm{s}(\mathfrak{g})=\{0\}$ and $C(x)=P_{\mathrm{s}(\mathfrak{g})}(xx^{T})=0$ for all $x\in\mathbb{R}^{2}$. By ˜10, $\mathrm{lip}f(x)=2\sqrt{4x_{1}^{2}+x_{2}^{2}}$ and thus $(0,0)$ is the only flat minimum by Corollary˜1. ∎
The converse of Theorem˜2 does hold in $\ell_{1}$-matrix factorization, as we show in Section˜4.3.
Corollary 6.
Under ˜3, if $\overline{x}$ is flat and $f$ is regular near $\overline{x}$, then there exists $\overline{v}\in\arg\max\{|v|:v\in\partial f(\overline{x})\}$ such that $C(\overline{v})=0$.
Proof.
This a consequence of ˜12, Corollary˜1, and Theorem˜2. ∎
Theorem˜2 can be generalized to higher orders, as follows.
Theorem 3.
Suppose $f$ is $D^{k}$ near $\overline{x}\in\mathbb{R}^{n}$ and $f^{(k)}(\overline{x})\neq 0$. Then
$\overline{x}\in\arg\mathrm{loc}\min\{\|f^{(k)}(x)\|:x\in G\overline{x}\}$
$\Longrightarrow$
$\exists\overline{v}\in\arg\max\{|f^{(k)}(\overline{x})(v,\ldots,v)|:|v|=1\}:~C(\overline{v})=0.$
Proof.
The proof naturally mirrors that of Theorem˜2. By assumption, $I_{n}$ is a local solution to
$\inf_{g\in G}\|f^{(k)}(g^{-1}\overline{x})\|^{2}$ $=\inf_{g\in G}\sup_{|v|=1}(f^{(k)}(g^{-1}\overline{x})(v,\ldots,v))^{2}=\inf_{g\in G}\sup_{|v|=1}(f^{(k)}(\overline{x})(gv,\ldots,gv))^{2}$
$=\inf_{g\in G}-\inf_{v\in\mathbb{R}^{n}}\phi(v,g)=-\sup_{g\in G}\varphi(g)$
where the second equality holds by ˜13. Also, $\phi:\mathbb{R}^{n}\times G\rightarrow\overline{\mathbb{R}}$ and $\varphi:G\rightarrow\mathbb{R}_{-}$ are defined by
$\phi(v,g)=\delta_{B^{n}}(v)-(f^{(k)}(\overline{x})(gv,\ldots,gv))^{2}~~~\text{and}~~~\varphi(g)=\inf_{v\in\mathbb{R}^{n}}\phi(v,g).$
Let $\gamma:(-\epsilon,\epsilon)\rightarrow G$ be a smooth curve such that $\gamma(0)=I_{n}$ where $\epsilon>0$. Consider the function $\tau:\mathbb{R}^{n}\times\mathbb{R}\rightarrow\overline{\mathbb{R}}$ defined by
$\tau(v,t)=\phi(v,\gamma(t))+\delta_{[-\epsilon/2,\epsilon/2]}(t).$
It holds that
$\forall t\in[-\epsilon/2,\epsilon/2],~~~(\varphi\circ\gamma)(t)=\inf_{v\in\mathbb{R}^{n}}\tau(v,t).$
Thus for all $(v,t)\in\mathrm{dom}\tau$, one has $\partial\tau(v,t)\subseteq$
$\begin{pmatrix}N_{B^{n}}(v)-2kf^{(k)}(\overline{x})(\gamma(t)v,\ldots,\gamma(t)v)\gamma(t)^{T}\nabla^{k}f(\overline{x})(\gamma(t)v,\ldots,\gamma(t)v)\\ N_{[-\epsilon/2,\epsilon/2]}(t)-2kf^{(k)}(\overline{x})(\gamma(t)v,\ldots,\gamma(t)v)f^{(k)}(\overline{x})(\gamma^{\prime}(t)v,\gamma(t)v,\ldots,\gamma(t)v)\end{pmatrix}$
and for all $\overline{t}\in[-\epsilon/2,\epsilon/2]\subseteq\mathrm{dom}(\varphi\circ\gamma)$,
$\partial(\varphi\circ\gamma)(\overline{t})\subseteq\bigcup_{\overline{v}\in\arg\min\limits_{v\in\mathbb{R}^{n}}\tau(v,\overline{t})}M(\overline{v},\overline{t})~~~\text{where}~~~M(\overline{v},\overline{t})=\{y\in\mathbb{R}:(0,y)\in\partial\tau(\overline{v},\overline{t})\}.$
In particular, since $0\in\partial(\varphi\circ\gamma)(0)$, there exists
$\overline{v}\in\arg\min\{\tau(v,0):v\in\mathbb{R}^{n}\}=\arg\max\{(f^{(k)}(\overline{x})(v,\ldots,v))^{2}:|v|=1\}$
such that
$(0,0)\in\partial\tau(\overline{v},0)\subseteq\begin{pmatrix}N_{B^{n}}(\overline{v})-2kf^{(k)}(\overline{x})(\overline{v},\ldots,\overline{v})\nabla^{k}f(\overline{x})(\overline{v},\ldots,\overline{v})\\ -2kf^{(k)}(\overline{x})(\overline{v},\ldots,\overline{v})f^{(k)}(\overline{x})(\gamma^{\prime}(0)\overline{v},\overline{v},\ldots,\overline{v})\end{pmatrix}$. (4)
Consider the Lagrangian
$L(v,\lambda)=(f^{(k)}(\overline{x})(v,\ldots,v))^{2}-\lambda(|x|^{2k}-1),$
following Lim [28, Section 4]. Since the constraint is qualified by gradient independence, the first-order optimality condition ensures the existence of $\overline{\lambda}\in\mathbb{R}$ such that
$\nabla_{v}L(\overline{v},\overline{\lambda})=2kf^{(k)}(\overline{x})(\overline{v},\ldots,\overline{v})\nabla^{k}f(\overline{x})(\overline{v},\ldots,\overline{v})-2k\overline{\lambda}\overline{v}=0,$
that is to say,
$f^{(k)}(\overline{x})(\overline{v},\ldots,\overline{v})\nabla^{k}f(\overline{x})(\overline{v},\ldots,\overline{v})=\overline{\lambda}\overline{v}.$
Taking the inner product with $\overline{v}$ yields
$0\neq|f^{(k)}(\overline{x})|^{2}=(f^{(k)}(\overline{x})(\overline{v},\ldots,\overline{v}))^{2}=\overline{\lambda}|\overline{v}|^{2}=\overline{\lambda}.$
The inclusion in (4) then implies that for all $u\in\mathfrak{g}$, we have
$0$ $=f^{(k)}(\overline{x})(\overline{v},\ldots,\overline{v})f^{(k)}(\overline{x})(u\overline{v},\overline{v},\ldots,\overline{v})$
$=f^{(k)}(\overline{x})(\overline{v},\ldots,\overline{v})\langle u\overline{v},\nabla^{k}f(\overline{x})(\overline{v},\ldots,\overline{v})\rangle$
$=\langle u\overline{v},\overline{\lambda}\overline{v}\rangle=\overline{\lambda}\langle u,\overline{v}\hskip 1.13809pt\overline{v}^{T}\rangle.$
As in the proof of Theorem˜2, we conclude that $C(\overline{v})=0$. ∎
Corollary 7.
Suppose $f$ is $D^{k}$ near $\overline{x}\in\mathbb{R}^{n}$ with $k\in\mathbb{N}^{*}$, $f^{(i)}(\overline{x})=0$ for all $i\in\llbracket 1,k-1\rrbracket$, and $f^{(k)}(\overline{x})\neq 0$. If $\overline{x}$ is flat, then there exists $\overline{v}\in\arg\max\{|f^{(k)}(\overline{x})(v,\ldots,v)|:|v|=1\}$ such that $C(\overline{v})=0$.
Proof.
This a consequence of ˜13, Corollary˜2, and Theorem˜3. ∎
4.3 Matrix factorization
We seek to establish converse results to those developed in the previous section, in the case of matrix factorization, in a desire to characterize flat minima. Given $m,n,r\in\mathbb{N}^{*}$ and $M\in\mathbb{R}^{m\times n}$, the map $F:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\ni(X,Y)\mapsto XY-M\in\mathbb{R}^{m\times n}$ is invariant under the action of $\mathrm{GL}(r,\mathbb{R})$ on $\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}$ defined by $(X,Y,A)\mapsto(XA,A^{-1}Y)$. As shown in [23, Example 3], for any locally Lipschitz function $g:\mathbb{R}^{m\times n}\to\mathbb{R}$, $g\circ F$ then admits the conserved quantity $C(X,Y)=X^{T}X-YY^{T}$ (see [3, 15, 21, 31] for other derivations). We successively treat the cases where $g=\|\cdot\|_{1}$ and $g=\|\cdot\|_{F}$. We begin with a simple fact.
Fact 15.
Let $D=\mathrm{diag}(d_{1}I_{k_{1}},\dots,d_{\ell}I_{k_{\ell}})$ with $d_{1}>\cdots>d_{\ell}$ and $k_{1}+\cdots+k_{\ell}=n$, and let $P\in\mathrm{O}(n)$. Then
$D=PDP^{T}~~~\Longleftrightarrow~~~\forall i\in\llbracket 1,\ell\rrbracket,~\exists P_{k_{i}}\in\mathrm{O}(k_{i}):~P=\mathrm{diag}(P_{k_{1}},\dots,P_{k_{\ell}}).$
Proof.
One direction is obvious. Suppose $D=PDP^{T}$ and let $p_{i}=(p_{ij})_{j}$ denote the $i$th column of $P$. Then $(d_{j}p_{ij})_{j}=Dp_{i}=PDP^{T}p_{i}=\sum_{j}d_{j}p_{j}p_{j}^{T}p_{i}=d_{i}p_{i}$, i.e., $(d_{j}-d_{i})p_{ij}=0$. ∎
˜15 implies that for any diagonal matrix $D$ of order $n$ with nonnegative entries and $P\in\mathrm{O}(n)$, one has $DP=PD$ iff $D^{1/2}P=PD^{1/2}$. The optimal value in the lemma below is already known [37, Lemma 1], but the local-global property appears to be new. Also, the characterization of global minima via the balance condition on $X$ and $Y$ was stated in [12, Lemma 2.2], although the proof seems invalid.
Lemma 4.
$2\|M\|_{*}=\min\{\|X\|_{F}^{2}+\|Y\|_{F}^{2}:XY=M\}$ and a feasible point $(X,Y)$ is a global minimum iff it is a local minimum iff $X^{T}X=YY^{T}$.
Proof.
Consider a singular value decomposition $M=U\Sigma V^{T}$ and $p=\mathrm{rank}\hskip 0.85358ptM$. Let $x_{i}^{T}$ denote the rows of $X$ and $y_{j}$ the columns of $Y$. If $XY=\Sigma$, then
$\|\Sigma\|_{*}=\sum_{i=1}^{p}\langle x_{i},y_{i}\rangle\leq\sum_{i=1}^{p}|x_{i}||y_{i}|\leq\sqrt{\sum_{i=1}^{p}|x_{i}|^{2}}\sqrt{\sum_{i=1}^{p}|y_{i}|^{2}}\leq\|X\|_{F}\|Y\|_{F}\leq(\|X\|_{F}^{2}+\|Y\|_{F}^{2})/2.$
If $XY=M$, the $U^{T}XYV=\Sigma$ and $2\|M\|_{*}=2\|\Sigma\|_{*}\leq\|U^{T}X\|_{F}^{2}+\|YV\|_{F}^{2}=\|X\|_{F}^{2}+\|Y\|_{F}^{2}$, with equality when $(X,Y)=(U\Sigma^{1/2},\Sigma^{1/2}V^{T})$.
If $(X,Y)$ is a local minimum then $XY=M$ and $X^{T}X=YY^{T}$ by Proposition˜7. Conversely, consider some compact singular value decompositions $X=U_{X}\Sigma_{X}V_{X}^{T}$ and $Y=U_{Y}\Sigma_{Y}V_{Y}^{T}$. Then $V_{X}\Sigma_{X}^{2}V_{X}^{T}=U_{Y}\Sigma_{Y}^{2}U_{Y}^{T}$, so that $\Sigma_{X}=\Sigma_{Y}$ and $\Sigma_{X}^{2}=(V_{X}^{T}U_{Y})\Sigma_{X}^{2}(V_{X}^{T}U_{Y})^{T}$. In the proof of [12, Lemma 2.2], one concludes that $V_{X}^{T}U_{Y}=I_{r}$, but that seems untrue (think of $\Sigma_{X}=I_{r}$). In fact, $\Sigma_{X}^{2}(V_{X}^{T}U_{Y})=(V_{X}^{T}U_{Y})\Sigma_{X}^{2}$ and thus $\Sigma_{X}(V_{X}^{T}U_{Y})=(V_{X}^{T}U_{Y})\Sigma_{X}$ by ˜15. Using a compact singular value decomposition $M=U\Sigma V^{T}$, this yields $XY=U_{X}\Sigma_{X}V_{X}^{T}U_{Y}\Sigma_{X}V_{Y}^{T}=U_{X}V_{X}^{T}U_{Y}\Sigma_{X}^{2}V_{Y}^{T}=U\Sigma V^{T}$, and so $\Sigma_{X}=\Sigma_{Y}=\sqrt{\Sigma}$. Thus $\|X\|_{F}^{2}+\|Y\|_{F}^{2}=2\|M\|_{*}$ and $(X,Y)$ is globally optimal. ∎
Using Lemma˜4, we can show that the converse of Theorem˜2 holds in $\ell_{1}$-matrix factorization.
Proposition 8.
Given $M\in\mathbb{R}^{m\times n}$, let $f:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\to\mathbb{R}$ be defined by $f(X,Y)=\|XY-M\|_{1}$. Let $\overline{X}\hskip 1.13809pt\overline{Y}=M$. The following are equivalent:
  1. (i)
    $(\overline{X},\overline{Y})\in\arg\min\{\mathrm{lip}f(X,Y):XY=M\}$;
  2. (ii)
    $(\overline{X},\overline{Y})\in\arg\mathrm{loc}\min\{\mathrm{lip}f(X,Y):XY=M\}$;
  3. (iii)
    $\exists(\overline{H},\overline{K})\in\arg\max\{\|H\|_{F}^{2}+\|K\|_{F}^{2}:(H,K)\in\partial f(\overline{X},\overline{Y})\}:~\overline{H}^{T}\overline{H}=\overline{K}\hskip 1.13809pt\overline{K}^{T}$.
There exist global minima since for all $XY=M$, $\mathrm{lip}f(X,Y)\geq\sqrt{\|X\|_{F}^{2}+\|Y\|_{F}^{2}}/\sqrt{m+n}$.
Proof.
(i) $\Longrightarrow$ (ii) is obvious. (ii) $\Longrightarrow$ (iii) is due to Theorem˜2. (iii) $\Longrightarrow$ (i) Observe that $f(X,Y)=\|F(X,Y)\|_{1}$ where F(X,Y)=XYMF(X,Y)=XY-M. We thus compute $F(X,Y)(H,K)=XK+HY$ and $F^{\prime}(X,Y)^{*}(\Lambda)=(X^{T}\Lambda,\Lambda Y^{T})$. By ˜10,
$\partial f(X,Y)=\left\{\begin{pmatrix}\Lambda Y^{T}\\ X^{T}\Lambda\end{pmatrix}:\Lambda\in\mathrm{sign}(XY-M)\right\}.$
By assumption, there exists $\overline{\Lambda}\in[-1,1]^{m\times n}$ such that $(\overline{H},\overline{K})=(\overline{\Lambda}\hskip 1.13809pt\overline{Y}^{T},\overline{X}^{T}\overline{\Lambda})$. Hence $\overline{H}\hskip 1.13809pt\overline{K}=\overline{\Lambda}\hskip 1.13809pt\overline{Y}^{T}\overline{X}^{T}\overline{\Lambda}=\overline{\Lambda}\hskip 1.13809ptM^{T}\overline{\Lambda}$. Since $\overline{H}^{T}\overline{H}=\overline{K}\hskip 1.13809pt\overline{K}^{T}$, by Lemma˜4,
$(\overline{H},\overline{K})\in\arg\min\{\|H\|_{F}^{2}+\|K\|_{F}^{2}:HK=\overline{\Lambda}\hskip 1.13809ptM^{T}\overline{\Lambda}\}.$
Thus, whenever $XY=M$, we have
$\mathrm{lip}f(X,Y)\geq\sqrt{\|\overline{\Lambda}Y^{T}\|_{F}^{2}+\|X^{T}\overline{\Lambda}\|_{F}^{2}}\geq\sqrt{\|\overline{H}\|_{F}^{2}+\|\overline{K}\|_{F}^{2}}=\mathrm{lip}f(\overline{X},\overline{Y}).$
Finally, let $x_{i}^{T}$ denote the rows of $X$ and $y_{j}$ denote the columns of $Y$. We have
$\mathrm{lip}f(X,Y)$ $=\max_{\Lambda\in[-1,1]^{m\times n}}\sqrt{\|\Lambda Y^{T}\|_{F}^{2}+\|X^{T}\Lambda\|_{F}^{2}}\geq\max\{|x_{1}|,\ldots,|x_{m}|,|y_{1}|,\ldots,|y_{n}|\}$
$\geq\sqrt{|x_{1}|^{2}+\cdots+|x_{m}|^{2}+|y_{1}|^{2}+\cdots+|y_{n}|^{2}}/\sqrt{m+n}$
$=\sqrt{\|X\|_{F}^{2}+\|Y\|_{F}^{2}}/\sqrt{m+n}.$
To see this, successively take $\Lambda$ with a single nonzero entry, equal to one, to generate the rows of $X$ and the columns of $Y$. ∎
We now consider matrix factorization with the Frobenius norm. It is known to have no spurious local minima, as proved in [5] when $m=n$, and in [38] for any $m,n$. We will show that it also has no spurious flat minima, i.e., flat minima that are not globally flat. To see why, it is useful to recall when equality holds in $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ where $(X,Y)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}$.
Given a matrix $A\in\mathbb{R}^{m\times n}$ with singular value decomposition $A=U\Sigma V^{T}$, let $S_{A}=\{v\in\mathbb{R}^{n}:|Av|=\|A\|_{2}|v|\}$ and $d_{A}=\dim S_{A}$, i.e., the multiplicity of the maximal singular value of $A$. Let $\widehat{U}$ (resp. $\widehat{V}$) denote the first $d_{A}$ columns of $U$ (resp. $V$), in other words, the maximal left (resp. right) singular vectors of $A$.
Fact 16.
$\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}\Longleftrightarrow S_{X}\cap YS_{Y}\neq\{0\}\Longleftrightarrow\mathrm{Im}\hskip 0.85358pt\widehat{V}_{X}\cap\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}\neq\{0\}$.
Proof.
For all $v\in\mathbb{R}^{n}$, $|XYv|\leq\|XY\|_{2}|v|\leq\|X\|_{2}\|Y\|_{2}|v|$ and $|XYv|\leq\|X\|_{2}|Yv|\leq\|X\|_{2}\|Y\|_{2}|v|$. Thus $S_{XY}\supset S_{X}\cap YS_{Y}$ and, if $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$, then $S_{XY}\subset S_{X}\cap YS_{Y}$. Lastly, $S_{X}=\mathrm{Im}\hskip 0.85358pt\widehat{V}_{X}$ and $YS_{Y}=\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}$. ∎
The optimal value in the next result is known in the multilayer case [33], a fortiori in the two layer case. On the other hand, the local-global property again appears to be new.
Lemma 5.
$2\|M\|_{2}=\min\{\|X\|_{2}^{2}+\|Y\|_{2}^{2}:XY=M\}$ and a feasible point $(X,Y)$ is a global minimum iff it is a local minimum iff $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ and $\|X\|_{2}=\|Y\|_{2}$ iff $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}$.
Proof.
Given a singular value decomposition $M=U\Sigma V^{T}$, one has
$2\|M\|_{2}=2\|XY\|_{2}\leq 2\|X\|_{2}\|Y\|_{2}\leq\|X\|_{2}^{2}+\|Y\|_{2}^{2}$
with equality exactly when $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ and $\|X\|_{2}=\|Y\|_{2}$. This happens in particular when $(X,Y)=(U\Sigma^{1/2},\Sigma^{1/2}V^{T})$. If $\|X\|_{2}\neq\|Y\|_{2}$, say $\|X\|_{2}<\|Y\|_{2}$, then $\|tX\|_{2}^{2}+\|t^{-1}Y\|_{2}^{2}<\|X\|_{2}^{2}+\|Y\|_{2}^{2}$ for all $t\in(1,\|Y\|_{2}^{1/2}\|X\|_{2}^{-1/2}]$. Suppose $\|XY\|_{2}<\|X\|_{2}\|Y\|_{2}$, then consider some singular value decompositions $X=U_{X}\Sigma_{X}V_{X}^{T}$ and $Y=U_{Y}\Sigma_{Y}V_{Y}^{T}$ and let $d_{X}$ denote the multiplicity of the maximal singular value of $X$. Consider $D_{t}=\mathrm{diag}(tI_{d_{X}},I_{r-d_{X}})$ where $t>0$. Let $X_{t}=XV_{X}D_{t}V_{X}^{T}$ and $Y_{t}=V_{X}D_{t}^{-1}V_{X}^{T}Y$. For all $t\in(0,1)$,
$\|X_{t}\|_{2}^{2}+\|Y_{t}\|_{2}^{2}$ $=\|XV_{X}D_{t}V_{X}^{T}\|_{2}^{2}+\|V_{X}D_{t}^{-1}V_{X}^{T}Y\|_{2}^{2}$
$=\|U_{X}\Sigma_{X}V_{X}^{T}V_{X}D_{t}V_{X}^{T}\|_{2}^{2}+\|V_{X}D_{t}^{-1}V_{X}^{T}U_{Y}\Sigma_{Y}V_{Y}^{T}\|_{2}^{2}$
$=\|\Sigma_{X}D_{t}\|_{2}^{2}+\|D_{t}^{-1}V_{X}^{T}U_{Y}\Sigma_{Y}\|_{2}^{2}$
$=t^{2}\|\Sigma_{X}\|_{2}^{2}+\|D_{t}^{-1}V_{X}^{T}U_{Y}\Sigma_{Y}\|_{2}^{2}$
$<t^{2}\|\Sigma_{X}\|_{2}^{2}+\|D_{t}^{-1}V_{X}^{T}\|_{2}^{2}\|U_{Y}\Sigma_{Y}\|_{2}^{2}$
$=t^{2}\|X\|_{2}^{2}+t^{-2}\|Y\|_{2}^{2}$
$<\|X\|_{2}^{2}+\|Y\|_{2}^{2}$
by ˜16. ∎
In contrast to the multilayer case [33], it is possible to determine the maximal eigenvalue and eigenspace of the Hessian at any global minimum.
Proposition 9.
Given $M\in\mathbb{R}^{m\times n}$, let $f:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\to\mathbb{R}$ be defined by $f(X,Y)=\|XY-M\|_{F}^{2}/2$. Suppose $XY=M$. Then
$\lambda_{1}(\nabla^{2}f(X,Y))=\|X\|_{2}^{2}+\|Y\|_{2}^{2}~~~\text{and}$
$E_{1}(\nabla^{2}f(X,Y))=\left\{\left(U_{X}\begin{pmatrix}A&0\\ 0&0\end{pmatrix}U_{Y}^{T},V_{X}\begin{pmatrix}B&0\\ 0&0\end{pmatrix}V_{Y}^{T}\right):\|X\|_{2}A=\|Y\|_{2}B\in\mathbb{R}^{d_{X}\times d_{Y}}\right\},$
given any singular value decompositions $X=U_{X}\Sigma_{X}V_{X}^{T}$ and $Y=U_{Y}\Sigma_{Y}V_{Y}^{T}$, where $d_{X}$ (resp. $d_{Y}$) denotes the multiplicity of the maximal singular value of $X$ (resp. $Y$). Also,
$\exists(H,K)\in E_{1}(\nabla^{2}((X,Y)):~\mathrm{Im}\hskip 0.85358ptH^{T}\cap\mathrm{Im}\hskip 0.85358ptK\neq\{0\}~~\Longrightarrow~~\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}.$
Proof.
Observe that $f(X,Y)=\|F(X,Y)\|_{F}^{2}/2$ where F(X,Y)=XYMF(X,Y)=XY-M. By ˜11, $\lambda_{1}(\nabla f(X,Y))=\|F^{\prime}(X,Y)\|_{2}^{2}$ so we compute $F^{\prime}(X,Y)(H,K)=HY+XK$. One has
$\|HY+XK\|_{F}$ $\leq\|HY\|_{F}+\|XK\|_{F}\leq\|H\|_{F}\|Y\|_{2}+\|X\|_{2}\|K\|_{F}$
$\leq\sqrt{\|H\|_{F}^{2}+\|K\|_{F}^{2}}\sqrt{\|X\|_{2}^{2}+\|Y\|_{2}^{2}}$
with equality when $H=U_{X}E^{mr}_{11}U_{Y}^{T}$ and $K=V_{X}E^{rn}_{11}V_{Y}^{T}$ (where $E^{pq}_{11}\in\mathbb{R}^{p\times q}$ with only one nonzero entry, $(1,1)$, equal to 1). To determine exactly when equality holds, first reduce to the diagonal case:
$\|F^{\prime}(X,Y)\|_{2}$ $=\max_{\|H\|_{F}^{2}+\|K\|_{F}^{2}\leq 1}\|HY+XK\|_{F}=\max_{\|H\|_{F}^{2}+\|K\|_{F}^{2}\leq 1}\|HU_{Y}\Sigma_{Y}V_{Y}^{T}+U_{X}\Sigma_{X}V_{X}^{T}K\|_{F}$
$=\max_{\|H\|_{F}^{2}+\|K\|_{F}^{2}\leq 1}\|U_{X}^{T}HU_{Y}\Sigma_{Y}+\Sigma_{X}V_{X}^{T}KV_{Y}\|_{F}=\max_{\|H\|_{F}^{2}+\|K\|_{F}^{2}\leq 1}\|H\Sigma_{Y}+\Sigma_{X}K\|_{F}.$
Equality in the Cauchy-Schwarz inequality holds iff $(\|H\|_{F},\|K\|_{F})$ and $(\|X\|_{2},\|Y\|_{2})$ are positively colinear. Let $h_{j}$ denote the columns of $H$, and $k_{i}^{T}$ the rows of $K$. One has $\|H\Sigma_{Y}\|_{F}=\|H\|_{F}\|\Sigma_{Y}\|_{2}$ iff $\sum_{j}(\Sigma_{Y})_{jj}^{2}|h_{j}|^{2}=\sum_{j}(\Sigma_{Y})_{11}^{2}|h_{j}|^{2}$ iff $\sum_{j}[(\Sigma_{Y})_{jj}^{2}-(\Sigma_{Y})_{11}^{2}]|h_{j}|^{2}=0$ iff $(\Sigma_{Y})_{jj}^{2}-(\Sigma_{Y})_{11}^{2}]|h_{j}|^{2}=0$ iff $h_{j}=0$ for all $j\notin\llbracket 1,d_{Y}\rrbracket$. Likewise, $\|\Sigma_{X}K\|_{F}=\|\Sigma_{X}\|_{2}\|K\|_{F}$ iff $k_{i}=0$ for all $i\notin\llbracket 1,d_{X}\rrbracket$. Next, $\|H\Sigma_{Y}+\Sigma_{X}K\|_{F}=\|H\Sigma_{Y}\|_{F}+\|\Sigma_{X}K\|_{F}$ iff $H\Sigma_{Y}$ and $\Sigma_{X}K$ are positively colinear. Hence all equalities hold iff $\|X\|_{2}H=\|Y\|_{2}K$ and $H_{ij}=K_{ij}=0$ for all $(i,j)\notin\llbracket 1,d_{X}\rrbracket\times\llbracket 1,d_{Y}\rrbracket$.
Let $\widehat{V}_{X}$ (resp. $\widehat{U}_{Y}$) denote the first $d_{X}$ (resp. $d_{Y}$) columns of $V_{X}$ (resp. $U_{Y}$). If $(H,K)\in E_{1}(\nabla^{2}((X,Y))$ and $\mathrm{Im}\hskip 0.85358ptH^{T}\cap\mathrm{Im}\hskip 0.85358ptK\neq\{0\}$, then $\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}A^{T}\cap\mathrm{Im}\hskip 0.85358pt\widehat{V}_{Y}B\neq\{0\}$, $\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}\cap\mathrm{Im}\hskip 0.85358pt\widehat{V}_{Y}\neq\{0\}$, and $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ by ˜16. ∎
It is now possible to completely characterize flat minima in matrix factorization. Recall that $\mathrm{Im}\hskip 0.85358ptAA^{T}=\mathrm{Im}\hskip 0.85358ptA$ for any matrix $A\in\mathbb{R}^{m\times n}$.
Proposition 10.
Given $M\in\mathbb{R}^{m\times n}$, let $f:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\to\mathbb{R}$ be defined by $f(X,Y)=\|XY-M\|_{F}^{2}/2$. Suppose $\overline{X}\hskip 1.13809pt\overline{Y}=M$. The following are equivalent:
  1. (i)
    $(\overline{X},\overline{Y})$ is globally flat;
  2. (ii)
    $(\overline{X},\overline{Y})$ is flat;
  3. (iii)
    $(\overline{X},\overline{Y})\in\arg\mathrm{loc}\min\{\lambda_{1}(\nabla^{2}f(X,Y)):XY=M\}$;
  4. (iv)
    $\|\overline{X}\|_{2}=\|\overline{Y}\|_{2}~\land~\exists(\overline{H},\overline{K})\in E_{1}(\nabla^{2}f(\overline{X},\overline{Y}))\setminus\{0\}:~\overline{H}^{T}\overline{H}=\overline{K}\hskip 1.13809pt\overline{K}^{T}$;
  5. (v)
    $(\overline{X},\overline{Y})\in\arg\min\{\lambda_{1}(\nabla^{2}f(X,Y)):XY=M\}$;
  6. (vi)
    $\|\overline{X}\|_{2}=\|\overline{Y}\|_{2}=\sqrt{\|M\|_{2}}$.
Proof.
(i) $\Longrightarrow$ (ii) follows by Definition˜3.
(ii) $\Longrightarrow$ (iii) is due to Corollary˜2.
(iii) $\Longrightarrow$ (iv) If $\|\overline{X}\|_{2}\neq\|\overline{Y}\|_{2}$, say $\|\overline{X}\|_{2}<\|\overline{Y}\|_{2}$, then $\|t\overline{X}\|_{2}^{2}+\|t^{-1}\overline{Y}\|_{2}^{2}<\|\overline{X}\|_{2}^{2}+\|\overline{Y}\|_{2}^{2}$ for all $t\in(1,\|\overline{Y}\|_{2}^{1/2}\|\overline{X}\|_{2}^{-1/2}]$. If $(\overline{X},\overline{Y})=(0,0)$, then take $\overline{H}^{T}$ and $\overline{K}$ to be the rectangular identities. Otherwise, $\nabla^{2}f(\overline{X},\overline{Y})\neq 0$ so Theorem˜3 applies.
(iv) $\Longrightarrow$ (v) By Lemma˜5, $\|\overline{X}\|_{2}=\|\overline{Y}\|_{2}$. Since $\overline{H}^{T}\overline{H}=\overline{K}\hskip 1.13809pt\overline{K}^{T}\neq 0$, $\mathrm{Im}\hskip 0.85358pt\overline{H}^{T}=\mathrm{Im}\hskip 0.85358pt\overline{H}^{T}\overline{H}=\mathrm{Im}\hskip 0.85358pt\overline{K}\hskip 1.13809pt\overline{K}^{T}=\mathrm{Im}\hskip 0.85358pt\overline{K}\neq\{0\}$ and so $\|\overline{X}\hskip 1.13809pt\overline{Y}\|_{2}=\|\overline{X}\|_{2}\|\overline{Y}\|_{2}$ by Proposition˜9.
(v) $\Longrightarrow$ (i) For any $XY=M$ such that $\|X\|_{2}=\|Y\|_{2}$, we have
$\accentset{\circ}{f}(X,Y,r)$ $=\max_{\|H\|_{F}^{2}+\|K\|_{F}^{2}\leq r^{2}}\|HY+XK+HK\|_{F}^{2}/2$
$=\max_{\|H\|_{F}^{2}+\|K\|_{F}^{2}\leq r^{2}}\|HU_{Y}\Sigma_{Y}V_{Y}^{T}+U_{X}\Sigma_{X}V_{X}^{T}K+HK\|_{F}^{2}/2$
$=\max_{\|H\|_{F}^{2}+\|K\|_{F}^{2}\leq r^{2}}\|U_{X}^{T}HU_{Y}\Sigma_{Y}+\Sigma_{X}V_{X}^{T}KV_{Y}+U_{X}^{T}HKV_{Y}\|_{F}^{2}/2$
$=\max_{\|H\|_{F}^{2}+\|K\|_{F}^{2}\leq r^{2}}\|H\Sigma_{Y}+\Sigma_{X}K+HU_{Y}^{T}V_{X}K\|_{F}^{2}/2$
$=\max_{\|H\|_{F}^{2}+\|K\|_{F}^{2}\leq r^{2}}(\|H\|_{F}\|Y\|_{2}+\|X\|_{2}\|K\|_{F}+\|H\|_{F}\|K\|_{F})^{2}/2$
$=(\sqrt{2}\|X\|_{2}r+r^{2}/2)^{2}/2$
$=(2\|X\|_{2}^{2}r^{2}+\sqrt{2}\|X\|_{2}r^{3}+r^{4}/4)/2$
$=(\|X\|_{2}^{2}+\|Y\|_{2}^{2})r^{2}/2+\sqrt{2}(\|X\|_{2}+\|Y\|_{2})r^{3}/4+r^{4}/8$
using the same direction to obtain the equalities as in first part of the proof of Proposition˜9. Let $XY=M$. Since $f$ is $D^{2}$ and $\nabla f(X,Y)=0$, by Corollary˜2, $\accentset{\circ}{f}(X,Y,r)=\lambda_{1}(\nabla^{2}f(X,Y))r^{2}/2+o(r^{2})$. Hence, if $\lambda_{1}(\nabla^{2}f(\overline{X},\overline{Y}))<\lambda_{1}(\nabla^{2}f(X,Y))$, then there exists $\overline{r}>0$ such that $\accentset{\circ}{f}(\overline{X},\overline{Y},r)<\accentset{\circ}{f}(X,Y,r)$ for all $r\in(0,\overline{r}]$. Otherwise, by Lemma˜5, $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}=\|\overline{X}\|_{2}=\|\overline{Y}\|_{2}$ and so $\accentset{\circ}{f}(\overline{X},\overline{Y},r)=\accentset{\circ}{f}(X,Y,r)$ for all $r\geq 0$. In either case, $(\overline{X},\overline{Y})\preceq(X,Y)$, hence $(\overline{X},\overline{Y})$ is globally flat.
(v) $\Longleftrightarrow$ (vi) This was already shown in Lemma˜5. ∎
Proposition˜10 implies that when $(X,Y)$ is a flat global minimum, there must exist a maximal eigenvector $(H,K)$ of the Hessian such that $H^{T}H=KK^{T}$. This appears to be new. On the other hand, it is possible that $X^{T}X\neq YY^{T}$, as noted in [33]. We give an example using a well-known fact.
Fact 17.
Given $M\in\mathbb{R}^{m\times n}$, consider a compact singular value decomposition $M=U\Sigma V^{T}$ and let $r=\mathrm{rank}\hskip 0.85358ptM$. For all $(X,Y)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}$,
$XY=M~~~\Longleftrightarrow~~~\exists A\in\mathrm{GL}(r,\mathbb{R}):(X,Y)=(U\Sigma^{1/2}A,A^{-1}\Sigma^{1/2}V^{T}).$
Proof.
$XY=M$ iff $\Sigma^{-1/2}U^{T}XYV\Sigma^{-1/2}=I_{r}$ iff $\Sigma^{-1/2}U^{T}X=A$ and $YV\Sigma^{-1/2}=A^{-1}$ for some $A\in\mathrm{GL}(r,\mathbb{R})$ iff $X=U\Sigma^{1/2}A$ and $Y=A^{-1}\Sigma^{1/2}V^{T}$. ∎
Example 5.
Given $a>b>0$, the flat minima of $\mathbb{R}^{2\times 2}\times\mathbb{R}^{2\times 2}\ni(X,Y)\mapsto\|XY-M\|_{F}^{2}\in\mathbb{R}$ with $M=\mathrm{diag}(a,b)$ are of the form
$(X,Y)=\left(\begin{pmatrix}\sqrt{a}&0\\ 0&\sqrt{b}t\end{pmatrix}Q,Q^{T}\begin{pmatrix}\sqrt{a}&0\\ 0&\sqrt{b}/t\end{pmatrix}\right)$
where $\sqrt{a/b}\leq t\leq\sqrt{a/b}$ and $Q\in\mathrm{O}(2)$. They are globally but not strictly flat, and satisfy $X^{T}X-YY^{T}=b\hskip 1.13809pt\mathrm{diag}(0,t^{2}-1/t^{2})$.
Proof.
By Proposition˜10 and ˜17, $(X,Y)$ is a flat minimum iff $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}$ and $(X,Y)=(M^{1/2}A,A^{-1}M^{1/2})$ for some $A\in\mathrm{GL}(2,\mathbb{R})$. Consider a decomposition $A=DNQ$ where $D=\mathrm{diag}(\alpha,\beta)$ with $\alpha,\beta>0$, $N$ is upper triangular of order 2 with ones on the diagonal with $N_{12}=\gamma\in\mathbb{R}$, and $Q\in\mathrm{O}(2)$. We have $\|X\|_{2}=\|M^{1/2}A\|_{2}=\|M^{1/2}DNQ\|_{2}=\|M^{1/2}DN\|_{2}$ and $\|Y\|_{2}=\|AM^{1/2}\|_{2}=\|Q^{-1}N^{-1}D^{-1}M^{1/2}\|_{2}=\|N^{-1}D^{-1}M^{1/2}\|_{2}$. Observe that $\|M^{1/2}DN\|_{2}=\|N^{-1}D^{-1}M^{1/2}\|_{2}=\|M^{1/2}\|_{2}=\sqrt{a}$ iff $\alpha=1$, $\sqrt{b/a}\leq\beta\leq\sqrt{a/b}$, and $N=I_{2}$. Indeed,
$M^{1/2}DN=\begin{pmatrix}\sqrt{a}&0\\ 0&\sqrt{b}\end{pmatrix}\begin{pmatrix}\alpha&0\\ 0&\beta\end{pmatrix}\begin{pmatrix}1&\gamma\\ 0&1\end{pmatrix}=\begin{pmatrix}\sqrt{a}\alpha&\sqrt{a}\alpha\gamma\\ 0&\sqrt{b}\beta\end{pmatrix}~~~\text{and}$
$N^{-1}D^{-1}M^{1/2}=\begin{pmatrix}1&-\gamma\\ 0&1\end{pmatrix}\begin{pmatrix}1/\alpha&0\\ 0&1/\beta\end{pmatrix}\begin{pmatrix}\sqrt{a}&0\\ 0&\sqrt{b}\end{pmatrix}=\begin{pmatrix}\sqrt{a}/\alpha&-\gamma\sqrt{b}/\beta\\ 0&\sqrt{b}/\beta\end{pmatrix}$.
Bear in mind that for any scalars $u$ and $v$, one has
$\left\|\begin{pmatrix}1&u\\ 0&v\end{pmatrix}\right\|_{2}=\max_{x^{2}+y^{2}=1}\sqrt{(x+uy)^{2}+v^{2}y^{2}}\geq\max_{x^{2}+y^{2}=1}|x+uy|=\sqrt{1+u^{2}}.\qed$
5 Examples
Some examples are in order.
Example 6.
A point $x$ is a flat minimum of $f(x)=(x_{2}\mathrm{ReLU}(x_{1})+x_{3}-1)^{2}$ iff $(x_{1}<0\land x_{3}=1)\lor(x_{1}=0\land|x_{2}|\leq 1\land x_{3}=1)$.
Proof.
Observe that $f$ fails to be regular at $(0,x_{2},x_{3})$ when $x_{2}(x_{3}-1)<0$, so the calculus rule in Corollary˜1 does not apply. Regardless, by Remark˜1, $\mathrm{lip}f(x)=0$ whenever $f(x)=0$, so it wouldn’t be of much use. We simply resort to the definition. If $(x_{1}<0\land x_{3}=1)\lor(x_{1}=0\land|x_{2}|\leq 1\land x_{3}=1)$, then $\accentset{\circ}{f}(x,r)=r^{2}$ for $r$ near 0. If $x_{1}=0\land|x_{2}|>1\land x_{3}=1$, then $\accentset{\circ}{f}(x,r)=x_{2}^{2}r^{2}>r^{2}$ near 0. Otherwise, $x_{1}>0$ and $f(x)=(x_{2}x_{1}+x_{3}-1)^{2}$, in which case
$\nabla f(x)=2(x_{2}x_{1}+x_{3}-1)\begin{pmatrix}x_{2}\\ x_{1}\\ 1\end{pmatrix}~~~\text{and}~~~\nabla^{2}f(x)=2\begin{pmatrix}x_{2}^{2}&2x_{2}x_{1}+x_{3}-1&x_{2}\\ 2x_{2}x_{1}+x_{3}-1&x_{1}^{2}&x_{1}\\ x_{2}&x_{1}&1\end{pmatrix}$
If $f(x)=0$, then
$\nabla^{2}f(x)=2\begin{pmatrix}x_{2}^{2}&x_{2}x_{1}&x_{2}\\ x_{2}x_{1}&x_{1}^{2}&x_{1}\\ x_{2}&x_{1}&1\end{pmatrix}=\begin{pmatrix}x_{2}\\ x_{1}\\ 1\end{pmatrix}\begin{pmatrix}x_{2}\\ x_{1}\\ 1\end{pmatrix}^{T}$
and so $\lambda_{1}(\nabla^{2}f(x))=x_{2}^{2}+x_{1}^{2}+1$. By Proposition˜4, $\accentset{\circ}{f}(x,r)=(x_{2}^{2}+x_{1}^{2}+1)r^{2}/2+o(r^{2})>r^{2}$. ∎
Example 7.
The flat minima of $f(x)=(x_{1}x_{2}-1)^{4}$ are $\pm(1,1)$.
Proof.
By ˜11, $\|f^{(4)}(x)\|=4|F^{\prime}(x)|_{2}^{4}$ when $F(x)=x_{1}x_{2}-1=0$, i.e., $\|f^{(4)}(x)\|=4(x_{1}^{2}+x_{2}^{2})^{2}$. This quantity is strictly minimized at $\pm(1,1)$ over the solution set, so one concludes by Corollary˜2. Alternatively, one notices that the flat minima are the same as that of $|x_{1}x_{2}-1|$ by using Definition˜3 and the fact that $t\in\mathbb{R}_{+}\mapsto t^{4}$ is strictly increasing. Since the new function is regular and locally Lipschitz, by Corollary˜1 it suffices to compute its Lipschitz modulus when $x_{1}x_{2}=1$, i.e., $x_{1}^{2}+x_{2}^{2}$. ∎
Example 8.
The origin is the sole flat minimum of $f(x)=x_{2}^{2}+x_{1}^{2}x_{2}^{4}$.
Proof.
Compute
$F^{\prime}(x)=2\begin{pmatrix}x_{1}x_{2}^{4}\\ x_{2}+2x_{1}^{2}x_{2}^{3}\end{pmatrix}~~~\text{and}~~~\nabla^{2}f(x)=2\begin{pmatrix}x_{2}^{4}&4x_{1}x_{2}^{3}\\ 4x_{1}x_{2}^{3}&1+6x_{1}^{2}x_{2}^{2}\end{pmatrix}$.
Thus $\lambda_{1}(\nabla^{2}f(x_{1},0))=2$, as claimed in the introduction. By virtue of Examples 2 and 3, it would seem futile to compute higher-order derivatives. But in this special case, the maximal eigenvectors of $f^{(2)}(x_{1},0)$ and $f^{(4)}(x_{1},0)$ align (with $(0,1)$) while $f^{(3)}(x_{1},0)=0$. Since $f^{(4)}(x_{1},0)=24x_{1}^{2}$, we deduce that $\accentset{\circ}{f}(x_{1},0,r)=r^{2}+x_{1}^{2}r^{4}+o(r^{4})$ and conclude by Definition˜3.
Example 9.
A global minimum of $f(x)=(a_{1}x_{1}^{2}+\cdots+a_{n}x_{n}^{2}-1)^{2}$ where $a\in\mathbb{R}^{n}$ is flat iff $a_{i}x_{i}=0$ for all $i\notin I=\arg\min\{a_{i}:a_{i}>0\}$.
Proof.
Without loss of generality, assume $a_{i}\neq 0$ for all $i\in\llbracket 1,n\rrbracket$. If $I=\emptyset$, then $f(x)=(|a_{1}|x_{1}^{2}+\cdots+|a_{n}|x_{n}^{2}+1)^{2}$ and every global minimum is flat. Otherwise, let $m$ be the cardinal of $I$. The objective $f$ is invariant under the natural action of $G=\mathrm{diag}(O(m),I_{n-m})$ once we reorder the indices so that those in $I$ come first. By ˜11, $\lambda_{1}(\nabla^{2}f(x))=2|F^{\prime}(x)|_{2}^{2}$ when $F(x)=a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}-1=0$. Since $F^{\prime}(x)=2(a_{1}x_{1},\dots,a_{n}x_{n})$, we seek to minimize $|F^{\prime}(x)|^{2}=4(a_{1}^{2}x_{1}^{2}+\dots+a_{n}^{2}x_{n}^{2})$ subject to $a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}=1$. Consider the Lagrangian $L(x,\lambda)=\sum_{i}a_{i}^{2}x_{i}^{2}+\lambda(1-\sum_{i}a_{i}x_{i}^{2})$. Since the constraint is qualified, at optimality $a_{i}^{2}x_{i}-\lambda a_{i}x_{i}=0$, namely, $a_{i}x_{i}(a_{i}-\lambda)=0$. If $\lambda\neq a_{i}$ for all $i$, then $0=a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}=1$, a contradiction. If $\lambda=a_{i}<0$ for some $i$, then $0>a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}=1$, a contradiction. If $\lambda=a_{i}>0$ for some $i$, then $a_{1}^{2}x_{1}^{2}+\dots+a_{n}^{2}x_{n}^{2}=\lambda$ and so $i\in I$. Thus, for any global minimum $\overline{x}$ of $f$ such that $\overline{x}_{i}=0$ for all $i\notin I$, $G\overline{x}$ is a strict global minimum of $\lambda_{1}(\nabla^{2}f(x))+\delta_{[f=f(\overline{x})]}$. It follows that every point in $G\overline{x}$ is flat by Corollary˜4. No other global minimum is flat by Corollary˜2. ∎
Example 10.
The flat global minima of $f(x)=|x_{1}x_{3}-a|+|x_{2}x_{3}-b|$ are
$x=\pm\left(a\sqrt{\frac{\sqrt{2}}{|a|+|b|}},b\sqrt{\frac{\sqrt{2}}{|a|+|b|}},\sqrt{\frac{|a|+|b|}{\sqrt{2}}}\right)$
if $(a,b)\neq(0,0)$, else $(0,0,0)$.
Proof.
By ˜10, $f$ is regular,
$\partial f(x)=\left\{\begin{pmatrix}\lambda_{1}x_{3}\\ \lambda_{2}x_{3}\\ \lambda_{1}x_{1}+\lambda_{2}x_{2}\end{pmatrix}~,\hskip 8.53581pt\lambda\in\begin{pmatrix}\mathrm{sign}(x_{1}x_{3}-a)\\ \mathrm{sign}(x_{2}x_{3}-b)\end{pmatrix}\right\}$,
and $\mathrm{lip}f(x)=2|x_{3}|^{2}+(|x_{1}|+|x_{2}|)^{2}$. Thus $(0,0,0)$ is the sole flat global minimum when $(a,b)=(0,0)$ by Corollary˜1. When $(a,b)\neq(0,0)$, $\mathrm{argmin}f=\{(at,bt,1/t):t\neq 0\}$. Accordingly, given $t\neq 0$, let $x_{t}=(at,bt,1/t)$ and compute
$\mathrm{lip}f(x_{t})^{2}$ $=\max\{|v|^{2}:v\in\partial f(x_{t})\}$
$=\max\{(\lambda_{1}/t)^{2}+(\lambda_{2}/t)^{2}+(\lambda_{1}at+\lambda_{2}bt)^{2}:\lambda_{1},\lambda_{2}\in[-1,1]\}$
$=\max\{(\lambda_{1}^{2}+\lambda_{2}^{2})/t^{2}+(\lambda_{1}a+\lambda_{2}b)^{2}t^{2}:\lambda_{1},\lambda_{2}\in[-1,1]\}$
$=2/t^{2}+(|a|+|b|)^{2}t^{2}.$
When $ab\neq 0$, the last max is reached exactly at $\pm(\mathrm{sign}(a),\mathrm{sign}(b))$, whence
$\mathrm{argmax}\{|v|:v\in\partial f(x_{t})\}=\left\{\pm\begin{pmatrix}\mathrm{sign}(a)/t\\ \mathrm{sign}(b)/t\\ (|a|+|b|)t\end{pmatrix}\right\}$.
Since
$\frac{d\mathrm{lip}f(x_{t})^{2}}{dt}=-\frac{4}{t^{3}}+2(|a|+|b|)^{2}t~~~\text{and}~~~\frac{d^{2}\mathrm{lip}f(x_{t})^{2}}{dt^{2}}=\frac{12}{t^{4}}+2(|a|+|b|)^{2}>0,$
the Lipschitz modulus is strictly minimized when
$|t|=\sqrt{\frac{\sqrt{2}}{|a|+|b|}}.$
This yields the expression of the flat global minima by Corollary˜1.
We now verify the claim of Proposition˜8. The objective $f$ is invariant under the natural action of the Lie group $G=\{\mathrm{diag}(t,t,1/t):t\neq 0\}$ whose Lie algebra is $\mathfrak{g}=\mathrm{span}\{\mathrm{diag}(1,1,-1)\}$. By Proposition˜6, a conserved quantity is given by $C(x)=x_{1}^{2}+x_{2}^{2}-x_{3}^{2}$. When $ab\neq 0$, the quantity
$C\left(\pm\begin{pmatrix}\mathrm{sign}(a)/t\\ \mathrm{sign}(b)/t\\ (|a|+|b|)t\end{pmatrix}\right)=(\mathrm{sign}(a)/t)^{2}+(\mathrm{sign}(b)/t)^{2}-[(|a|+|b|)t]^{2}=2/t^{2}-(|a|+|b|)^{2}t^{2}$
indeed cancels out iff $x_{t}$ is flat. (The same holds with $ab=0$, but we omit this case for brevity.) On the other hand,
$C(x_{t})=(a^{2}+b^{2})t^{2}-1/t^{2}\neq 0$
unless $|a|=|b|$. Indeed, $C(x_{t})=0$ for a flat minimum if and only if
$\sqrt{\frac{\sqrt{2}}{|a|+|b|}}=\frac{1}{\sqrt[4]{a^{2}+b^{2}}}~\Longleftrightarrow~2(a^{2}+b^{2})=(|a|+|b|)^{2}~\Longleftrightarrow~a^{2}+b^{2}=2|ab|~\Longleftrightarrow~|a|=|b|.$
Example 11.
The flat global minima of $f(x)=(x^{\upsilon}-1)^{2}$ where $x^{\upsilon}=x_{1}^{\upsilon_{1}}\cdots x_{n}^{\upsilon_{n}},\upsilon\in{\mathbb{N}^{*}}^{n},$ are
$|x_{i}|=\frac{\sqrt{\upsilon_{i}}}{\sqrt{\upsilon_{1}^{\upsilon_{1}}\cdots\upsilon_{n}^{\upsilon_{n}}}^{1/|\upsilon|_{1}}},~~~i=1,\ldots,n,$
for any choice of signs such that $x^{\upsilon}=1$. Any solution to
$\left\{\begin{array}[]{ccl}\dot{x}_{i}&=&-\upsilon_{n}x_{i}(\upsilon_{n}x_{i}^{2}-\upsilon_{i}x_{n}^{2}),~i=1,\ldots,n-1,\\ ~\dot{x}_{n}&=&\sum_{i=1}^{n-1}\upsilon_{i}x_{n}(\upsilon_{n}x_{i}^{2}-\upsilon_{i}x_{n}^{2}),\end{array}\right$.
initialized at a global minimum is globally defined, flattens over time, and converges to a flat global minimum.
Proof.
Let $\overline{\upsilon}=(\upsilon_{1},\ldots,\upsilon_{n-1})$. The objective is invariant under the natural action of the Lie group
$G=\{\mathrm{diag}(t_{1}^{\upsilon_{n}},\ldots,t_{n-1}^{\upsilon_{n}},t^{-\overline{\upsilon}}):t_{1},\ldots,t_{n-1}>0\}$
whose Lie algebra is
$\mathfrak{g}=\mathrm{span}\{\mathrm{diag}(\upsilon_{n},0,\ldots,0,-\upsilon_{1}),\ldots,\mathrm{diag}(0,\ldots,0,\upsilon_{n},-\upsilon_{n-1})\}.$
By Proposition 6, a conserved quantity is given by
$C(x)=(\upsilon_{n}x_{1}^{2}-\upsilon_{1}x_{n}^{2},\ldots,\upsilon_{n}x_{n-1}^{2}-\upsilon_{n-1}x_{n}^{2}).$
The connected components of the global minima of $f$ are homogeneous $G$-spaces. Indeed, given $\overline{x}$ a global minimum of $f$, we have
$G\overline{x}=\{x\in\mathbb{R}^{n}:x^{\upsilon}=1,~\mathrm{sign}(x)=\mathrm{sign}(\overline{x})\}.$
Let $x$ be such that $x^{\upsilon}=1$ and $\mathrm{sign}(x)=\mathrm{sign}(\overline{x})$. Then $x=\mathrm{diag}(t_{1}^{\upsilon_{n}},\ldots,t_{n-1}^{\upsilon_{n}},t^{-\overline{\upsilon}})\overline{x}$ with $t_{i}=(x_{i}/\overline{x}_{i})^{1/\upsilon_{n}}$ since $t_{i}^{\upsilon_{n}}\overline{x}_{i}=x_{i}$ for all $i$ and $t^{-\overline{\upsilon}}\overline{x}_{n}=t^{-\overline{\upsilon}}\overline{x}_{1}^{-\upsilon_{1}/\upsilon_{n}}\cdots\overline{x}_{n-1}^{-\upsilon_{n-1}/\upsilon_{n}}=(t_{1}^{\upsilon_{n}}\overline{x}_{1})^{-\upsilon_{1}/\upsilon_{n}}\cdots(t_{n-1}^{\upsilon_{n}}\overline{x}_{n-1})^{-\upsilon_{n-1}/\upsilon_{n}}=x_{1}^{-\upsilon_{1}/\upsilon_{n}}\cdots x_{n-1}^{-\upsilon_{n-1}\upsilon_{n}}=x_{n}$.
It is now possible to determine flat points. Let $F(x)=x^{\upsilon}-1$ and compute
$F^{\prime}(x)=(\upsilon_{1}x^{\upsilon}/x_{1},\ldots,\upsilon_{n}x^{\upsilon}/x_{n}).$
By ˜11, $\lambda_{1}(\nabla^{2}f(x))=2|F^{\prime}(x)|^{2}$ when $F(x)=0$. Given $t_{1},\ldots,t_{n-1}>0$, let $x_{t}=\mathrm{diag}(t_{1}^{\upsilon_{n}},\ldots,t_{n-1}^{\upsilon_{n}},t^{-\overline{\upsilon}})\overline{x}$ where $\overline{x}$ is global minimum such that $|\overline{x}_{i}|=1$ for all $i$. We have
$|F^{\prime}(x_{t})|^{2}=\upsilon_{1}^{2}t_{1}^{-2\upsilon_{n}}+\cdots+\upsilon_{n-1}^{2}t_{n-1}^{-2\upsilon_{n}}+\upsilon_{n}^{2}t^{2\overline{\upsilon}}$
and
$\frac{\partial|F^{\prime}(x_{t})|}{\partial t_{i}}=2\upsilon_{i}\upsilon_{n}(-\upsilon_{i}t_{i}^{-2\upsilon_{n}}+\upsilon_{n}t^{2\overline{\upsilon}})/t_{i}.$
Thus
$\nabla|F^{\prime}(x_{t})|^{2}=0\Longleftrightarrow\upsilon_{1}t_{1}^{-2\upsilon_{n}}=\cdots=\upsilon_{n-1}t_{n-1}^{-2\upsilon_{n}}=\upsilon_{n}t^{2\overline{\upsilon}}$
and
$\frac{\partial^{2}|F^{\prime}(x_{t})|^{2}}{\partial t_{i}\partial t_{j}}=\left\{\begin{array}[]{cl}2\upsilon_{i}\upsilon_{n}((2\upsilon_{n}+1)\upsilon_{i}t_{i}^{-2\upsilon_{n}}+(2\upsilon_{i}-1)\upsilon_{n}t^{2\overline{\upsilon}})/t_{i}^{2}&\text{if}~i=j,\\ 4\upsilon_{i}\upsilon_{j}\upsilon_{n}^{2}t^{2\overline{\upsilon}}/(t_{i}t_{j})&\text{else}.\end{array}\right$.
Observe that
$\nabla|F^{\prime}(x_{t})|^{2}=0~~~\Longrightarrow~~~\nabla^{2}|F^{\prime}(x_{t})|^{2}=4\upsilon_{n}^{2}t^{2\overline{\upsilon}}\left(\mathrm{diag}(\upsilon\oslash t)^{2}+(\upsilon\oslash t)(\upsilon\oslash t)^{T}\right)\succ 0$
[…]
$\lambda_{1}(\nabla^{2}f(x_{t}))=2(\upsilon_{1}^{2}t_{1}^{-2\upsilon_{n}}+\cdots+\upsilon_{n-1}^{2}t_{n-1}^{-2\upsilon_{n}}+\upsilon_{n}^{2}t^{2\overline{\upsilon}})$
is positive and coercive.
Observe that a global minimum $x$ is flat iff $C(x)=0$. The ‘only if’ part follows from the formula we just found, so it remains to check the ‘if’ part. If $C_{i}(x)=\upsilon_{n}x_{i}^{2}-\upsilon_{i}x_{n}^{2}=0$, then $|x_{i}|=\sqrt{\upsilon_{i}/\upsilon_{n}}|x_{n}|$,
$|x|^{\upsilon}=\prod_{i=1}^{n}\sqrt{\upsilon_{i}/\upsilon_{n}}^{\upsilon_{i}}|x_{n}|^{\upsilon_{i}}=1,~~~\text{and}~~~|x_{n}|^{|\upsilon|_{1}}=\prod_{i=1}^{n}\sqrt{\upsilon_{n}/\upsilon_{i}}^{\upsilon_{i}}=\frac{\sqrt{\upsilon_{n}}^{|\upsilon|_{1}}}{\prod_{i=1}^{n}\sqrt{\upsilon_{i}^{\upsilon_{i}}}}.$
Accordingly, let $c(x)=\|C(x)\|_{F}^{2}/4$. Notice that a global minimum $x$ is flat iff $\nabla c(x)=0$. For all $x\in\mathbb{R}^{n}$, we have
$C_{i}(\nabla f(x))$ $=4(x^{\upsilon}-1)^{2}[\upsilon_{n}(\upsilon_{i}x^{\upsilon}/x_{i})^{2}-\upsilon_{i}(\upsilon_{n}x^{\upsilon}/x_{n})^{2}]$
$=4(x^{\upsilon}-1)^{2}\upsilon_{i}\upsilon_{n}(x^{\upsilon}/(x_{i}x_{n}))^{2}[\upsilon_{i}x_{n}^{2}-\upsilon_{n}x_{i}^{2}]$
$=-4\upsilon_{i}\upsilon_{n}((x^{\upsilon}-1)(x^{\upsilon}/(x_{i}x_{n}))^{2}C_{i}(x).$
Let $\overline{x}\in[f=0]$ be such that $C(\overline{x})\neq 0$. There is an index $i_{0}$ such that $C_{i_{0}}(\overline{x})\neq 0$. For all $x\in\mathbb{R}^{n}$ near $\overline{x}$, we have
$\langle C(x),C(\nabla f(x))\rangle$ $=\sum_{i=1}^{n-1}4C_{i}(x)[-\upsilon_{i}\upsilon_{n}((x^{\upsilon}-1)/(x_{i}x_{n}))^{2}C_{i}(x)]$
$=-4(x^{\upsilon}-1)^{2}\frac{\upsilon_{n}x^{2\upsilon}}{x_{n}^{2}}\sum_{i=1}^{n-1}\frac{\upsilon_{i}C_{i}(x)^{2}}{x_{i}^{2}}\leq-4(x^{\upsilon}-1)^{2}\frac{\upsilon_{n}x^{2\upsilon}}{x_{n}^{2}}\frac{\upsilon_{i_{0}}C_{i_{0}}(\overline{x})^{2}}{x_{i_{0}}^{2}}$
$\leq-4\omega(x^{\upsilon}-1)^{2}x^{2\upsilon}\sum_{i=0}^{n}\frac{\upsilon_{i}^{2}}{x_{i}^{2}}=-\omega|\nabla f(x)|^{2}$
for some constant $\omega>0$. Together with Proposition˜6, this means that ˜2 holds. Applying Corollary˜5, $\lambda_{1}(\nabla^{2}f)$ strictly decreases along the trajectories of $\nabla c$. Since $\lambda_{1}(\nabla^{2}f)+\delta_{[f=0]}$ is coercive, they must be bounded and hence globally defined by [21, Proposition 2]. Since $c$ is semi-algebraic, they must also converge to a point $\overline{x}$ where $\nabla c(\overline{x})=0$ [30], i.e., a flat global minimum. ∎
To avoid overburdening the examples, we have not mentioned that all the flat minima are, in fact, globally flat.
Acknowledgments
I would like to thank Wenqing Ouyang for his careful reading of the manuscript and valuable feedback.

### v2

@@
-assumeAssumption
+On the geometry of flat minima
 Abstract
@@
 Flat minima were informally introduced by Hochreiter and Schmidhuber [22, 23] in the context of deep learning as connected regions of minima where the objective is nearly constant.
-In a desire to formalize this idea, it was later suggested to use the volume of connected components of sublevel sets [13, Definition 1].
-Around the same time, an empirical measure of sharpness [26, Metric 2.1] was proposed to analyze the role of batch size in the stochastic subgradient method for training neural networks.
-For nonnegative functions, it was redefined as the maximal ratio of the function variation over one plus the function value, over a ball of fixed radius [13, Definition 2].
+In a desire to formalize this idea, it was later suggested to use the volume of connected components of sublevel sets [15, Definition 1].
+Around the same time, an empirical measure of sharpness [31, Metric 2.1] was proposed to analyze the role of batch size in the stochastic subgradient method for training neural networks.
+For nonnegative functions, it was redefined as the maximal ratio of the function variation over one plus the function value, over a ball of fixed radius [15, Definition 2].
 More recently, flat minima of deep matrix factorization were defined as global minima which minimize the trace [18] or the maximum eigenvalue [38, 34, 37] of the Hessian of the objective function.
@@
 where $\mathrm{ReLU}(t)=\max\{0,t\}$.
-The objective function is differentiable at its global minima, so its gradient is defined and equal to zero (details in Example˜6).
+The objective function is differentiable at its global minima, so its gradient is defined and equal to zero (details in Example˜5.1).
 But it is not twice differentiable there, so its Hessian is not defined.
-As for the volumes of sublevel sets, they are generally not finite on ReLU networks [13, Theorem 2], nor on the most simple neural network, i.e., $f(x)=(x_{1}x_{2}-1)^{2}$.
+As for the volumes of sublevel sets, they are generally not finite on ReLU networks [15, Theorem 2], nor on the most simple neural network, i.e., $f(x)=(x_{1}x_{2}-1)^{2}$.
 Even when the Hessian is available, flatness is currently ill-defined.
@@
 obeys $\nabla f(x_{1},0)=(0,0)$ and $\lambda_{1}(\nabla^{2}f(x_{1},0))=2$ for all $x_{1}\in\mathbb{R}$, according to which all the global minima $(x_{1},0)$ are allegedly flat.
-But factoring in the fourth-order growth actually suggests that $(0,0)$ is the unique flat minimum (see Example˜8).
+But factoring in the fourth-order growth actually suggests that $(0,0)$ is the unique flat minimum (see Example˜5.3).
 This is not merely a theoretical matter, as gradient descent with sufficiently slowly diminishing step lengths converges to the origin.
@@
 For instance, if $x$ is a flat minimum of a smooth function $f$, then there exists a maximal eigenvector $v$ of $\nabla^{2}f(x)$ such that $C(v)=0$.
-This clarifies the picture in matrix factorization
+Third, we determine flat minima in a series of examples that were out of reach with previous definitions and tools.
+They demonstrate that the definition of flatness proposed in this work is both versatile and workable.
+We also illustrate the connection with symmetry and conservation.
+This work clarifies the picture in overparamatrized matrix factorization
 $\begin{array}[]{cccc}f:&\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}&\longrightarrow&\mathbb{R}\\ &(X,Y)&\longmapsto&\|XY-M\|_{F}^{2}\end{array}$
-where $M\in\mathbb{R}^{m\times n}$, $\|\cdot\|_{F}$ is the Frobenius norm, and $C(X,Y)=X^{T}X=YY^{T}$.
+where $M\in\mathbb{R}^{m\times n}$, $\|\cdot\|_{F}$ is the Frobenius norm, $r\geq\mathrm{rank}\hskip 0.85358pt(M)$, and $C(X,Y)=X^{T}X=YY^{T}$.
 Some authors [14] report the flat minima are “balanced”, meaning that $X^{T}X=YY^{T}$, while others [38] say that they are “nearly balanced”.
-With our definition, we find that while $X^{T}X-YY^{T}$ can be nonzero, there does exist a balanced maximal eigenvector of the Hessian $\nabla^{2}f(X,Y)$.
+With our definition, we find that balanced global minima are flat, but the converse is false.
+While $X^{T}X-YY^{T}$ can be nonzero at flat minima, there exists a balanced maximal eigenvector of the Hessian $\nabla^{2}f(X,Y)$.
 In passing, we note that flat minima admit a simple characterization.
@@
 While it is known the every local minimum is a global minimum [5, 46], it was not known that locally flat minima are globally flat.
-Third, we determine flat minima in a series of examples that were out of reach with previous definitions and tools.
-They demonstrate that the definition of flatness proposed in this work is both versatile and workable.
-We also illustrate the connection with symmetry and conservation.
-In one example,
-$f(x)=(x_{1}^{\upsilon_{1}}\cdots x_{n}^{\upsilon_{n}}-1)^{2}$
-where $\upsilon_{1},\ldots,\upsilon_{n}$ are positive integers, any gradient trajectory of the conserved quantity initialized at a global minimum converges to a flat global minimum.
-It does so while remaining in the (n1)(n-1)-dimensional solution manifold.
+A related novelty is that any gradient trajectory of the conserved quantity $\|C\|_{F}^{2}$ initialized at a global minimum converges to a flat global minimum.
 The paper is organized as follows.
@@
 Given a symmetric matrix $M\in\mathbb{R}^{n\times n}$, $\lambda_{1}(M)$ and $E_{1}(M)$ respectively denote its maximal eigenvalue and its associated eigenspace, whose nonzero elements are referred to as maximal eigenvectors.
+We use $M\succcurlyeq 0$ to denote that $M$ is positive semidefinite.
 2.1 Ordered sets
@@
 […]
+Let $S$ be a subset of an ordered set $X$.
+An element $y\in X$ is a least (resp.
+greatest) element of $S$ if $y\in S$ and $x\leq y$ (resp.
+$y\leq x$) for all $x\in S$.
+A lower (resp.
+upper) bound of $S$ is an element $y\in X$ such that $y\leq x$ (resp.
+$x\leq y$) for all $x\in S$.
+A lower bound of $S$ is an infimum of $S$ if it is a greatest lower bound.
+If an infimum exists, then it is unique.
+An infimum is a minimum if it belongs to $S$.
+The set of infima and minima are respectively denoted $\inf S$ and $\min S$.
+Suprema and maxima are defined analogously, and denoted $\sup S$ and $\max S$ respectively.
+A totally ordered set is complete if every nonempty subset that has a lower bound, has a greatest lower bound.
+The real numbers $\mathbb{R}$ is such an example.
+Let $S\subseteq\mathbb{R}$.
+If $S$ has no lower bound, then $\inf S=$-$\infty$.
+If $S=\emptyset$, then $\inf S=\infty$.
 2.2 Differential calculus
@@
 When it is real-valued, i.e., $W=\mathbb{R}$, $F$ is called a multilinear form.
-Given a symmetric multilinear map $F:V\times\cdots\times V\to W$ and norms $\|\cdot\|_{V},\|\cdot\|_{W}$ on $V,W$, consider the norm [7, Theorem A]
+Given a symmetric multilinear map $F:V\times\cdots\times V\to W$ and norms $\|\cdot\|_{V},\|\cdot\|_{W}$ on $V,W$, consider the norm [8, Theorem A]
 $\|F\|=\sup_{\|v\|_{V}=1}\|F(v,\ldots,v)\|_{W}.$
@@
 $L$-Lipschitz continuous).
-If $f$ is $D^{k}$ at $\overline{x}$, then $f^{(k)}$ is a symmetric multilinear map [17, Proposition C.16].
+If $f$ is $D^{k}$ at $\overline{x}$, then $f^{(k)}$ is a symmetric multilinear map [19, Proposition C.16].
 When evaluated on the diagonal, we use the shorthand $f^{(k)}(x)v^{k}=f^{(k)}(x)(v,\ldots,v)$.
@@
 2.3 Differential geometry
-An action of a group $G$ with identity $e$ on a set $M$ [27, p. 161] is a map $\theta:G\times M\rightarrow M$ such that
+An action of a group $G$ with identity $e$ on a set $M$ [32, p. 161] is a map $\theta:G\times M\rightarrow M$ such that
   1. (i)
@@
 It is homogeneous if for all $x,y\in M$, there exists $g\in G$ such that $\theta(g,x)=y$.
-A function $f:M\rightarrow N$ between sets $M$ and $N$ is invariant under an action of a group $G$ on $M$ if $f(gx)=f(x)$ for all $g\in G$ and $x\in M$.
+A function $f:M\rightarrow N$ between sets $M$ and $N$ is invariant under an action of a group $G$ on $M$ if $f(\theta(g,x))=f(x)$ for all $g\in G$ and $x\in M$.
 A Lie group $G$ is a smooth manifold and a group whose operations are smooth.
@@
 A Lie subgroup of a Lie group $G$ is a subgroup of $G$ is endowed with a topology and smooth structure making into a Lie group and an immersed submanifold.
-Topologically closed subgroups of Lie groups are Lie subgroups by the closed subgroup theorem [27, Theorem 20.12].
+Topologically closed subgroups of Lie groups are Lie subgroups by the closed subgroup theorem [32, Theorem 20.12].
 Let $\mathfrak{g}$ denote the Lie algebra of a Lie group $G$, which we identify with its tangent space at $e$.
 Let $I_{n}$ denote the identity matrix of order $n$.
-The set of invertible matrices with real coefficients of order $n$, denoted $\mathrm{GL}(n,\mathbb{R})$, is a Lie group.
-The natural action of a Lie subgroup $G$ of $\mathrm{GL}(n,\mathbb{R})$ on $\mathbb{R}^{n}$ is defined by the matrix vector multiplication $G\times\mathbb{R}^{n}\ni(g,x)\mapsto gx\in\mathbb{R}^{n}$.
-The orthogonal group $\mathrm{O}(n)=\{Q\in\mathbb{R}^{n\times n}:Q^{T}Q=I_{n}\}$ is a Lie subgroup of $\mathrm{GL}(n,\mathbb{R})$.
+The set of invertible matrices with real coefficients of order $n$, denoted $\mathrm{GL}(n)$, is a Lie group.
+The natural action of a Lie subgroup $G$ of $\mathrm{GL}(n)$ on $\mathbb{R}^{n}$ is defined by the matrix vector multiplication $G\times\mathbb{R}^{n}\ni(g,x)\mapsto gx\in\mathbb{R}^{n}$.
+The orthogonal group $\mathrm{O}(n)=\{Q\in\mathbb{R}^{n\times n}:Q^{T}Q=I_{n}\}$ is a Lie subgroup of $\mathrm{GL}(n)$.
 Given a Lipschitz continuous function $F:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $x_{0}\in\mathbb{R}^{n}$, consider the ODE
@@
 which we will also denote by $\arg\min\{f(x):x\in X\}$. A local minimum is spurious if it is not a global minimum.
-It will be convenient to use the generalization of local optiamlity from points to sets proposed in [25, 24].
+It will be convenient to use the generalization of local optimality from points to sets proposed in [29, 28].
 A nonempty set $S\subseteq\mathbb{R}^{n}$ is a local minimum (resp.
@@
 $f=\ell]=\{x\in\mathbb{R}^{n}:f(x)=\ell\}$
-and define other expressions like $f\leq\ell]$ similarly. Let
-$\mathrm{dom}f=\{x\in\mathbb{R}^{n}:f(x)<\infty\}~~~\text{and}~~~\mathrm{gph}f=\{(x,t)\in\mathbb{R}^{n}\times\mathbb{R}:f(x)=t\}.$
-The indicator of $S\subseteq\mathbb{R}^{n}$ is defined by
-$\delta_{S}(x)=\left\{\begin{array}[]{cl}0&\text{if}~x\in S,\\ \infty&\text{if}~x\notin S.\end{array}\right$.
-A function $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is lower semicontinuous at $\overline{x}\in\mathrm{dom}f$ if $\liminf_{x\rightarrow\overline{x}}f(x)\geq f(\overline{x})$ [36, Definition 1.5].
+and define other expressions like $f\leq\ell]$ similarly. Let $\mathrm{dom}f=\{x\in\mathbb{R}^{n}:f(x)<\infty\}$,
+$\mathrm{gph}f=\{(x,t)\in\mathbb{R}^{n}\times\mathbb{R}:f(x)=t\},~~~\text{and}~~~\mathrm{epi}f=\{(x,t)\in\mathbb{R}^{n}\times\mathbb{R}:f(x)\leq t\}.$
+The indicator and the support function of a set $C\subseteq\mathbb{R}^{n}$ are respectively defined by
+$\forall x\in\mathbb{R}^{n},~\delta_{C}(x)=\left\{\begin{array}[]{cl}0&\text{if}~x\in C,\\ \infty&\text{if}~x\notin C.\end{array}\right.~~~\text{and}~~~\forall v\in\mathbb{R}^{n},~\sigma_{C}(v)=\sup_{w\in C}\langle v,w\rangle.$
+Given $C\subseteq\mathbb{R}^{n}$ and $\overline{x}\in C$, the tangent cone [44, Definition 6.1] is defined by
+$T_{C}(\overline{x})=\{v\in\mathbb{R}^{n}:~\exists x_{k}\xrightarrow[C]{}\overline{x}~,~\tau_{k}\searrow 0:(x^{k}-\overline{x})/\tau^{k}\to v\}$
+where $x_{k}\xrightarrow[C]{}\overline{x}$ is a shorthand for $x_{k}\rightarrow\overline{x}$ and $x_{k}\in C$.
+A function $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is lower semicontinuous at $\overline{x}\in\mathrm{dom}f$ if $\liminf_{x\rightarrow\overline{x}}f(x)\geq f(\overline{x})$ [44, Definition 1.5].
 It is lower semicontinuous if it is so at every point in its domain.
-The regular normal cone and normal cone [36, Definition 6.3] are defined by
+The regular normal cone and normal cone [44, Definition 6.3] are defined by
 $\widehat{N}_{C}(\overline{x})=\{v\in\mathbb{R}^{n}:\langle v,x-\overline{x}\rangle\leq o(|x-\overline{x}|)~\text{for}~x\in C~\text{near}~\overline{x}\},$
@@
 $\limsup_{\scriptsize\begin{array}[]{c}x\xrightarrow[C]{}\overline{x}\\ x\neq\overline{x}\end{array}}\frac{\langle v,x-\overline{x}\rangle}{|x-\overline{x}|}\leqslant 0.$
-A set $C\subseteq\mathbb{R}^{n}$ is regular [36, Definition 6.4] at one of its points $\overline{x}$ if it is locally closed and $\widehat{N}_{C}(\overline{x})=N_{C}(\overline{x})$.
-Given $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ and a point $\overline{x}\in\mathbb{R}^{n}$ where $f(\overline{x})$ is finite, the regular subdifferential, subdifferential [36, Definition 8.3], and Clarke subdifferential of $f$ at $\overline{x}$ [14, Definition 4.1] are
+A set $C\subseteq\mathbb{R}^{n}$ is regular [44, Definition 6.4] at one of its points $\overline{x}$ if it is locally closed and $\widehat{N}_{C}(\overline{x})=N_{C}(\overline{x})$.
+The polar set of $C\subseteq\mathbb{R}^{n}$ [6, Definition 6.22] is defined by
+$C^{*}=\{v\in\mathbb{R}^{n}:~\forall w\in C,~\langle v,w\rangle\leq 0\}.$
+By [44, Theorem 6.28], the relationship $\widehat{N}_{C}(\overline{x})=T_{C}(\overline{x})^{*}$ holds.
+Given $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ and a point $\overline{x}\in\mathbb{R}^{n}$ where $f(\overline{x})$ is finite, the one-sided directional derivative (if it exists in $\mathbb{R}\cup\{\pm\infty\}$) and the subderivative for $w\in\mathbb{R}^{n}$ are respectively given by
+$f^{\prime}(\overline{x};w)=\lim_{\tau\searrow 0}\frac{f(\overline{x}+\tau w)-f(\overline{x})}{\tau}~~~\text{and}~~~df(\overline{x})(w)=\liminf_{\scriptsize\begin{array}[]{c}\tau\searrow 0\\ w^{\prime}\to w\end{array}}\frac{f(\overline{x}+\tau w^{\prime})-f(\overline{x})}{\tau}.$
+According to [44, Definition 7.20], $f$ is semidifferential at $\overline{x}$ for $w\in\mathbb{R}^{n}$ if
+$\lim_{\scriptsize\begin{array}[]{c}\tau\searrow 0\\ w^{\prime}\to w\end{array}}\frac{f(\overline{x}+\tau w^{\prime})-f(\overline{x})}{\tau}\in\mathbb{R}\cup\{\pm\infty\}.$
+If this holds for all $w\in\mathbb{R}^{n}$, then $f$ is semidifferentiable at $\overline{x}$.
+Given $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ and a point $\overline{x}\in\mathbb{R}^{n}$ where $f(\overline{x})$ is finite, the regular subdifferential, subdifferential, horizon subdifferential [44, Definition 8.3], and Clarke subdifferential of $f$ at $\overline{x}$ [16, Definition 4.1] are respectively given by
 $\widehat{\partial}f(\overline{x})=\{v\in\mathbb{R}^{n}:f(x)\geq f(\overline{x})+\langle v,x-\overline{x}\rangle+o(|x-\overline{x}|)~\text{near}~\overline{x}\},$
 $\partial f(\overline{x})=\{v\in\mathbb{R}^{n}:\exists(x_{k},v_{k})\in\mathrm{gph}\hskip 1.42262pt\widehat{\partial}f:(x_{k},f(x_{k}),v_{k})\rightarrow(\overline{x},f(\overline{x}),v)\},$
+$\partial^{\infty}f(\overline{x})=\{v\in\mathbb{R}^{n}:\exists(x_{k},v_{k})\in\mathrm{gph}\hskip 1.42262pt\widehat{\partial}f:\exists\tau_{k}\searrow 0:(x_{k},f(x_{k}),\tau_{k}v_{k})\rightarrow(\overline{x},f(\overline{x}),v)\},$
 $\overline{\partial}f(\overline{x})=\overline{\mathrm{co}}[\partial f(\overline{x})+\partial^{\infty}f(\overline{x})],$
-where $\mathrm{co}$} denotes the convex hull, and $\overline{\mathrm{co}}$ its closure.
-The $o$ means that $\liminf[f(x)-f(\overline{x})-\langle v,x-\overline{x}\rangle]/|x-\overline{x}|\geq 0$ where $\overline{x}\neq x\to\overline{x}$.
+where $\mathrm{co}$} denotes the convex hull, and $\overline{\mathrm{co}}$ its closure. The $o$ means that
+$\liminf_{\scriptsize\begin{array}[]{c}x\rightarrow\overline{x}\\ x\neq\overline{x}\end{array}}\frac{f(x)-f(\overline{x})-\langle v,x-\overline{x}\rangle}{|x-\overline{x}|}\geq 0.$
 A point $x\in\mathbb{R}^{n}$ is critical (resp.
@@
 $0\in\overline{\partial}f(x)$).
-A function $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is regular [36, Definition 7.25] at $\overline{x}$ if $f(\overline{x})$ is finite and $\mathrm{epi}f$ is regular at $(\overline{x},f(\overline{x}))$ as a subset of $\mathbb{R}^{n+1}$.
-The Lipschitz modulus of a function $f:\mathbb{R}^{n}\to\mathbb{R}$ is defined by [36, p.
+A function $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is regular [44, Definition 7.25] at $\overline{x}$ if $f(\overline{x})$ is finite and $\mathrm{epi}f$ is regular at $(\overline{x},f(\overline{x}))$ as a subset of $\mathbb{R}^{n+1}$.
+The Lipschitz modulus of a function $f:\mathbb{R}^{n}\to\mathbb{R}$ is defined by [44, p.
 354]
@@
 In what amounts to a slight abuse of notation, we may replace $D$ by some $D^{\prime}\subseteq D$ such that $D\setminus D^{\prime}$ has zero Lebesgue measure.
-Still assuming $\mathrm{lip}f(\overline{x})<\infty$, by [36, Theorem 9.13] $\partial f(\overline{x})$ is nonempty and compact, and $\mathrm{lip}f(\overline{x})=\max\{|v|:v\in\partial f(\overline{x})\}$.
-By [36, Theorem 9.61], $\overline{\nabla}f(\overline{x})$ is a nonempty compact subset of $\partial f(\overline{x})$ and $\mathrm{co}\overline{\nabla}f(\overline{x})=\mathrm{co}\hskip 0.85358pt\partial f(\overline{x})=\overline{\partial}f(\overline{x})$.
+Still assuming $\mathrm{lip}f(\overline{x})<\infty$, by [44, Theorem 9.13] $\partial f(\overline{x})$ is nonempty and compact, and $\mathrm{lip}f(\overline{x})=\max\{|v|:v\in\partial f(\overline{x})\}$.
+By [44, Theorem 9.61], $\overline{\nabla}f(\overline{x})$ is a nonempty compact subset of $\partial f(\overline{x})$ and $\mathrm{co}\overline{\nabla}f(\overline{x})=\mathrm{co}\hskip 0.85358pt\partial f(\overline{x})=\overline{\partial}f(\overline{x})$.
 Thus $\arg\max\{|v|:v\in\partial f(\overline{x})\}=\arg\max\{|v|:v\in\overline{\nabla}f(\overline{x})\}$, using the following fact.
-Fact 1.
+Fact 2.1.
 For any set $X\subseteq\mathbb{R}^{n}$, $\arg\max\{|x|:x\in X\}=\arg\max\{|x|:x\in\mathrm{co}\hskip 0.85358ptX\}$.
@@
 $\{x\in\mathbb{R}^{n}:f_{1}(x)>0,\ldots,f_{p}(x)>0,f_{p+1}(x)=0,\ldots,f_{q}(x)=0\}$
-where $f_{1},\ldots,f_{q}$ are polynomials with real coefficients. We adopt [41, Definition p. 503-506] below.
-Definition 1.
+where $f_{1},\ldots,f_{q}$ are polynomials with real coefficients. We adopt [49, Definition p. 503-506] below.
+Definition 2.2.
 An o-minimal structure on the real field is a sequence $S=(S_{k})_{k\in\mathbb{N}}$ such that for all $k\in\mathbb{N}$:
@@
 Throughout this paper, we fix an arbitrary o-minimal structure $(S_{k})_{k\in\mathbb{N}}$.
-A key property of univariate definable functions is that they satisfy the monotonicity theorem [41, 4.1].
+A key property of univariate definable functions is that they satisfy the monotonicity theorem [49, 4.1].
 It states that on bounded open intervals, for any $p\in\mathbb{N}$ there exist finitely many open subintervals where the function is $C^{p}$ and either constant or strictly monotone.
-The extension of the monotonicity theorem to multivariate functions is the cell decomposition theorem [41, 4.2], which is used to prove the definable Morse-Sard theorem [8, Corollary 9].
+The extension of the monotonicity theorem to multivariate functions is the cell decomposition theorem [49, 4.2], which is used to prove the definable Morse-Sard theorem [9, Corollary 9].
 It asserts that that lower semi-continuous definable functions have finitely many Clarke critical values.
@@
 Below are two ways to measure the variation of a function around a point.
-Definition 2.
+Definition 3.1.
 Given $f:\mathbb{R}^{n}\to\mathbb{R}$, let $\accentset{\circ}{f},\overline{f}:\mathbb{R}^{n}\times\mathbb{R}_{+}\to\overline{\mathbb{R}}$ be defined by
@@
 The following definition is the central tenet of this paper.
-Definition 3.
+Definition 3.2.
 Consider the binary relation on $\mathbb{R}^{n}$ defined by
@@
 Given $x$, $f$ is constant iff $\overline{f}(x,\ell)=\infty$ for all $\ell>0$, and $f$ is continuous at $x$ iff $\overline{f}(x,\ell)>0$ for all $\ell>0$.
-A set $X\subset\mathbb{R}^{n}$ is nowhere dense if its closure has empty interior.
+A set $X\subseteq\mathbb{R}^{n}$ is nowhere dense if its closure has empty interior.
 If $f$ is continuous and $\alpha\in\mathbb{R}$, then $f$ is nonconstant near every point in $f=\alpha]$ iff $f=\alpha]$ is nowhere dense.
-A function $\varphi:S\to\overline{\mathbb{R}}$ where $S\subset\mathbb{R}$ is increasing (resp.
+A function $\varphi:S\to\overline{\mathbb{R}}$ where $S\subseteq\mathbb{R}$ is increasing (resp.
 strictly increasing) if $\varphi(s)\leq\varphi(t)$ (resp.
@@
 Given $x\in\mathbb{R}^{n}$, let $\accentset{\circ}{f}_{x}=\accentset{\circ}{f}(x,\cdot)$ and $\overline{f}_{x}=\overline{f}(x,\cdot)$, both of which are increasing.
-Fact 2.
+Fact 3.3.
 $\preceq$ is a preorder. If $f$ is definable, then it is a total preorder and
@@
 […]
-We now gather several useful facts when $f$ is continuous (i.e., Facts 3-8).
-Fact 3.
+We now gather several useful facts when $f$ is continuous (i.e., Facts 3.4-3.9).
+Fact 3.4.
 $\forall(x,\ell)\in\mathbb{R}^{n}\times\mathbb{R}_{+},~P_{[|f-f(x)|\geq\ell]}(x)=P_{[|f-f(x)|=\ell]}(x)~\land~\overline{f}(x,\ell)=d(x,[|f-f(x)|=\ell])$.
@@
 There exists a natural duality between $\accentset{\circ}{f}$ and $\overline{f}$.
-Fact 4.
+Fact 3.5.
 $\forall(x,\ell)\in\mathbb{R}^{n}\times\mathbb{R}^{+}$, $\overline{f}(x,\ell)=\inf\{r\geq 0:\accentset{\circ}{f}(x,r)\geq\ell\}$.
@@
 ∎
-Fact 5.
+Fact 3.6.
 $\forall(x,r)\in\mathbb{R}^{n}\times\mathbb{R}_{+}$, $\accentset{\circ}{f}(x,r)=\sup\{\ell\geq 0:\overline{f}(x,\ell)\leq r\}$.
@@
 ∎
-˜4 and ˜5 actually hold without continuity if one uses strict relations, including in the definition of $\accentset{\circ}{f}$ and $\overline{f}$, with the convention that $\accentset{\circ}{f}(x,0)=\overline{f}(x,0)=0$.
-Fact 6.
+Fact˜3.5 and Fact˜3.6 actually hold without continuity if one uses strict relations, including in the definition of $\accentset{\circ}{f}$ and $\overline{f}$, with the convention that $\accentset{\circ}{f}(x,0)=\overline{f}(x,0)=0$.
+Fact 3.7.
 If $f$ is nonconstant near $x$, then the following are equivalent:
@@
 Proof.
-(i) $\Longrightarrow$ (ii) Let $\overline{\ell}=\accentset{\circ}{f}(x,\overline{r})>0$ and $\ell\in(0,\overline{\ell}]$. By ˜4,
+(i) $\Longrightarrow$ (ii) Let $\overline{\ell}=\accentset{\circ}{f}(x,\overline{r})>0$ and $\ell\in(0,\overline{\ell}]$. By Fact˜3.5,
 $\overline{f}(x,\ell)$ $=\inf\{r\geq 0:\accentset{\circ}{f}(x,r)\geq\ell\}$
@@
 Let $\overline{r}=\accentset{\circ}{f}(y,\overline{\ell})/2>0$ and $r\in(0,\overline{r}]$.
-Since $\overline{f}(x,\ell)\geq\overline{f}(x,\overline{\ell})\geq\overline{f}(y,\overline{\ell})=2\overline{r}>r$ for all $\ell\in(\overline{\ell},\infty)$, by ˜5,
+Since $\overline{f}(x,\ell)\geq\overline{f}(x,\overline{\ell})\geq\overline{f}(y,\overline{\ell})=2\overline{r}>r$ for all $\ell\in(\overline{\ell},\infty)$, by Fact˜3.6,
 $\accentset{\circ}{f}(x,r)$ $=\sup\{\ell\geq 0:\overline{f}(x,\ell)\leq r\}$
@@
 $\leq\sup\{\ell\geq 0:\overline{f}(y,\ell)\leq r\}=\accentset{\circ}{f}(y,r).$
-Fact 7.
+Fact 3.8.
 $\accentset{\circ}{f}$ is continuous.
@@
 ∎
-Fact 8.
+Fact 3.9.
 $\overline{f}$ is lower semicontinuous.
@@
 ∎
-Since $\accentset{\circ}{f}$ is continuous, one has equality in ˜4, namely, $\overline{f}(x,\ell)=\inf\{r\geq 0:\accentset{\circ}{f}(x,r)=\ell\}$.
+Since $\accentset{\circ}{f}$ is continuous, one has equality in Fact˜3.5, namely, $\overline{f}(x,\ell)=\inf\{r\geq 0:\accentset{\circ}{f}(x,r)=\ell\}$.
 This holds by the intermediate value theorem as $\accentset{\circ}{f}(x,0)=0$.
-In contrast, since $\overline{f}$ is merely lower semicontinuous, equality does not necessarily hold in ˜5 (think of the Cantor function).
+In contrast, since $\overline{f}$ is merely lower semicontinuous, equality does not necessarily hold in Fact˜3.6 (think of the Cantor function).
 Likewise, $\accentset{\circ}{f}_{x}$ need not be strictly increasing.
 To obtain such properties and others, more assumptions are needed.
-Assumption 1.
+Assumption 3.10.
 Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz and nonconstant near $x\in\mathbb{R}^{n}$.
@@
 Let $\overline{\ell}=\accentset{\circ}{f}(x,\overline{r})$.
-Proposition 1.
-Under ˜1,
+Proposition 3.11.
+Under Assumption˜3.10,
   1. (i)
@@
 Proof.
-(i) & (ii) Continuity of $\accentset{\circ}{f}_{x}$ follows from ˜7.
+(i) & (ii) Continuity of $\accentset{\circ}{f}_{x}$ follows from Fact˜3.8.
 Clearly $\accentset{\circ}{f}_{x}$ is increasing on $0,\overline{r}]$.
@@
 Otherwise, by continuity there exists $y\in\overline{B}_{r}(x)$ such that $\accentset{\circ}{f}(x,r)=|f(y)-f(x)|$, which must be a local maximum of |ff(x)||f-f(x)|.
-Since $\accentset{\circ}{f}(x,r)>0$, $y$ is either a local minimum or a local maximum of $f$, and hence $0\in\partial f(y)=\mathrm{co}\hskip 0.85358pt\partial f(y)=\overline{\partial}f(y)$ or $0\in\partial(-f)(y)\subseteq\overline{\partial}(-f)(y)=-\overline{\partial}f(y)$ by Fermat’s rule [27, Theorem 10.1].
+Since $\accentset{\circ}{f}(x,r)>0$, $y$ is either a local minimum or a local maximum of $f$, and hence $0\in\partial f(y)=\mathrm{co}\hskip 0.85358pt\partial f(y)=\overline{\partial}f(y)$ or $0\in\partial(-f)(y)\subseteq\overline{\partial}(-f)(y)=-\overline{\partial}f(y)$ by Fermat’s rule [32, Theorem 10.1].
 Since $\mathbb{R}\setminus\{f(x)\}$ contains no Clarke critical value of $f$ reached in $\overline{B}_{\overline{r}}(x)$, we have $f(y)=f(x)$.
 It follows that $\accentset{\circ}{f}(x,r)=0$, a contradiction.
-By ˜4 and the discussion below ˜8, $\accentset{\circ}{f}_{x}(\overline{f}_{x}(\ell))=\ell$ for all $\ell\in[0,\overline{\ell}]$.
+By Fact˜3.5 and the discussion below Fact˜3.9, $\accentset{\circ}{f}_{x}(\overline{f}_{x}(\ell))=\ell$ for all $\ell\in[0,\overline{\ell}]$.
 Thus for all $r\in[0,\overline{r}]$, $\accentset{\circ}{f}_{x}((\overline{f}_{x}\circ\accentset{\circ}{f}_{x})(r))=(\accentset{\circ}{f}_{x}\circ\overline{f}_{x})(\accentset{\circ}{f}_{x}(r))=\accentset{\circ}{f}_{x}(r)$ and $(\overline{f}_{x}\circ\accentset{\circ}{f}_{x})(r)=r$.
@@
 Flatness can be equivalently defined using $\overline{f}$ under minimal assumptions.
-Proposition 2.
+Proposition 3.12.
 Suppose $f:\mathbb{R}^{n}\to\mathbb{R}$ is continuous and $f=\alpha]$ is nowhere dense for some $\alpha\in\mathbb{R}$. Then
@@
 Proof.
-The first point follows from ˜6, while the second follows from Facts 2-7, Proposition˜1, and the definable Morse-Sard theorem [8, Corollary 9]. ∎
+The first point follows from Fact˜3.7, while the second follows from Facts 3.3-3.8, Proposition˜3.11, and the definable Morse-Sard theorem [9, Corollary 9]. ∎
 3.2 Curve selection
+[41, Definition 3.1]
 With the above definitions in place, it is natural to try and select optimal curves. This will be useful later.
-Lemma 1.
-Under ˜1,
+Lemma 3.13.
+Under Assumption˜3.10,
   1. (i)
-    There exist $\gamma:[0,\overline{r}]\to\mathbb{R}^{n}$, $\lambda:[0,\overline{r}]\to\mathbb{R}$, and $v:[0,\overline{r}]\to\mathbb{R}^{n}$ such that for all $r\in[0,\overline{r}]$, $\gamma(r)=x+\lambda(r)v(r)\in\arg\max_{\overline{B}_{r}(x)}|f-f(x)|$ and $v(r)\in\overline{\partial}f(\gamma(r))$.
+    There exist $\gamma:[0,\overline{r}]\to\mathbb{R}^{n}$, $\lambda:[0,\overline{r}]\to\mathbb{R}$, and $v:[0,\overline{r}]\to\mathbb{R}^{n}$ such that for all $r\in[0,\overline{r}]$, $\gamma(r)=x+\lambda(r)v(r)\in\arg\max_{\overline{B}_{r}(x)}|f-f(x)|$, $|\gamma(r)-x|=r$, and $v(r)\in\overline{\partial}f(\gamma(r))$.
   2. (ii)
-    If $x$ is a local minimum of $f$, then $\lambda(r)>0$ and $v(r)\in-\partial(-f)(\gamma(r))$ for all $r\in(0,\overline{r}]$ after possibly reducing $\overline{r}$.
-If in addition, $f$ is regular near $x$, then $v(r)\in\partial f(\gamma(r))$ for all $r\in(0,\overline{r}]$.
+    If $x$ is a local minimum of $f$, then $\lambda(r)>0$ and $v(r)\in-\partial(-f)(\gamma(r))$ for all $r\in(0,\overline{r}]$.
+If in addition, $f$ is regular, then one may choose $v(r)\in\overline{\nabla}f(\gamma(r))$ for all $r\in(0,\overline{r}]$.
   3. (iii)
@@
 Proof.
-(i) Let $(r,\ell)\in\mathrm{gph}\accentset{\circ}{f}_{x}$.
-There exists $\gamma(r)\in\overline{B}_{r}(x)$ such that $\accentset{\circ}{f}(x,r)=|f(\gamma(r))-f(x)|$.
-By [36, Theorem 8.15], $0\in\partial|f-f(x)|(\gamma(r))+N_{\overline{B}_{r}(x)}(\gamma(r))$.
-If $r=0$, then let $\lambda(0)=0$ and $v(0)=0$.
+(i) By Proposition˜3.11 (iii), there exists $\gamma(r)\in S_{r}(x)$ such that $\accentset{\circ}{f}(x,r)=|f(\gamma(r))-f(x)|$.
+The first-order optimality condition [44, Theorem 8.15] reads $0\in\partial(-|f-f(x)|)(\gamma(r))+N_{\overline{B}_{r}(x)}(\gamma(r))$.
+If $r=0$, then let $\lambda(0)=0$ and $v(0)\in\overline{\partial}f(x)\neq\emptyset$ by [44, Theorem 9.13].
 Otherwise, there exist $w(r)\in\partial(s(r)f)(\gamma(r))\subseteq\mathrm{co}\hskip 0.85358pt\partial(s(r)f)(\gamma(r))=\overline{\partial}(s(r)f)(\gamma(r))=s(r)\overline{\partial}f(\gamma(r))$ with $s(r)=\mathrm{sign}(f(x)-f(\gamma(r)))$ and $\mu(r)\geq 0$ such that $w(r)+\mu(r)(\gamma(r)-x)=0$.
@@
 We may hence define $\lambda(r)=-s(r)/\mu(r)$ and $v(r)=s(r)w(r)$.
-(ii) When $x$ is a local minimum of $f$, s(r)=1s(r)=-1. By [36, Theorem 9.16],
-$\lim_{\tau\searrow 0}\frac{f(\gamma(r)+\tau w)-f(\gamma(r))}{\tau}=\lim_{\scriptsize\begin{array}[]{c}\tau\searrow 0\\ w^{\prime}\to w\end{array}}\frac{f(\gamma(r)+\tau w^{\prime})-f(\gamma(r))}{\tau}=\max_{v\in\partial f(\gamma(r))}\langle w,v\rangle.$
-Since the directional derivative is nonpositive for any direction $w$ such that $\langle w,\gamma(r)-x\rangle<0$, this also holds for any direction $w$ such that $\langle w,\gamma(r)-x\rangle=0$.
-Given $v\in\partial f(\gamma(r))$, let $w=v|\gamma(r)-x|^{2}-\langle v,\gamma(r)-x\rangle(\gamma(r)-x)$.
-Since $\langle w,\gamma(r)-x\rangle=0$, it follows that $\langle w,v\rangle=|v|^{2}|\gamma(r)-x|^{2}-\langle v,\gamma(r)-x\rangle^{2}\leq 0$.
-There is equality in the Cauchy-Schwarz inequality, so there exists $\mu\in\mathbb{R}$ such that $v=\mu(\gamma(r)-x)$.
-Since $\langle x-\gamma(r),\gamma(r)-x\rangle<0$, we have $\langle x-\gamma(r),v\rangle=-\mu|\gamma(r)-x|^{2}\leq 0$ and so $\mu\geq 0$.
-Since $\partial f(\gamma(r))\neq\emptyset$ by [36, Theorem 9.13] and $\mathbb{R}\setminus\{f(x)\}$ contains no Clarke critical value of $f$ reached in $\overline{B}_{\overline{r}(x)}$, there exists $v(r)\in\partial f(\gamma(r))\setminus\{0\}$ and it satisfies $\gamma(r)-x=\lambda(r)v(r)$ for some $\lambda(r)>0$.
-(iii) When $f$ is definable, $0,\overline{r}]\ni r\rightrightarrows\arg\max_{\overline{B}_{r}(x)}|f-f(x)|$ is definable, there exists a definable selection $\gamma$ [41, 4.5].
+(ii) When $x$ is a local minimum of $f$, s(r)=1s(r)=-1.
+Since $\gamma(r)$ is a local minimum of $g=-f+\delta_{B_{r}(x)}$, $dg(\gamma(r))\geq 0$ by Fermat’s rule [44, Theorem 10.1].
+As $f$ is locally Lipschitz regular, by [44, Theorem 9.16] $f$ is semidifferentiable and $df(\cdot)=\sigma_{\partial f(\cdot)}$.
+Thus f-f is semidifferentiable and d(f)=dfd(-f)=-df.
+Since f-f is semidifferentiable, $dg=d(-f+\delta_{B_{r}(x)})=d(-f)+d\delta_{B_{r}(x)}$.
+Thus $dg(\gamma(r))=-\sigma_{\partial f(\gamma(r))}+\sigma_{N_{B_{r}(x)}(\gamma(r))}\geq 0$ and $\sigma_{\partial f(\gamma(r))}\leq\sigma_{N_{B_{r}(x)}(\gamma(r))}$, namely $\partial f(\gamma(r))\subseteq N_{B_{r}(x)}(\gamma(r))=\mathbb{R}_{+}(\gamma(r)-x)$.
+But $\partial f$ is nonempty convex compact valued by [44, Theorem 9.61], and so $\partial f(\gamma(r))=[a(r),b(r)](\gamma(r)-x)$ for some $0<a(r)\leq b(r)$ (recall that $0\notin\overline{\partial}f(\gamma(r))$).
+Also $\mathrm{co}\hskip 0.85358pt\overline{\nabla}f(\gamma(r))=\partial f(\gamma(r))$, hence $b(r)(\gamma(r)-x)\in\nabla f(\gamma(r))$.
+Now, simply let $v(r)=b(r)(\gamma(r)-x)$ and $\lambda(r)=1/b(r)>0$.
+(iii) When $f$ is definable, $0,\overline{r}]\ni r\rightrightarrows\arg\max_{\overline{B}_{r}(x)}|f-f(x)|$ is definable, so there exists a definable selection $\gamma$ [49, 4.5].
 Consequently $0,\overline{r}]\ni r\rightrightarrows\{v\in\overline{\partial}f(\gamma(t)):\exists\lambda\in\mathbb{R}:\gamma(t)=x+\lambda v\}$ is definable, yielding a definable selection $v$.
@@
 By the mean value theorem, $\gamma_{i}^{\prime}(0)=\lim_{r\searrow 0}(\gamma_{i}(r)-x_{i})/r=\lim_{r\searrow 0}\gamma_{i}^{\prime}(r)$.
-Finally, one can take $v(0)=\lim_{r\searrow 0}v(r)\in\overline{\partial}f(\gamma(0))$, where the limit exists because $v(r)$ is bounded by [36, Theorem 9.13] and [4, Proposition 3 p.
+Finally, one can take $v(0)=\lim_{r\searrow 0}v(r)\in\overline{\partial}f(\gamma(0))$, where the limit exists because $v(r)$ is bounded by [44, Theorem 9.13] and [4, Proposition 3 p.
 42].
@@
 We begin with the nonsmooth case.
-Lemma 2.
+Lemma 3.14.
 If $f:\mathbb{R}^{n}\to\mathbb{R}$ is Lipschitz continuous and regular near $x\in\mathbb{R}^{n}$, then
@@
 Proof.
-Since $f$ is Lipschitz continuous near $x$, by [36, Theorem 9.13] $\partial f(x)$ is nonempty and compact, and $\mathrm{lip}f(x)=\max\{|v|:v\in\partial f(x)\}$.
-Since $f$ is Lipschitz continuous and regular near $x$, by [36, Theorem 9.16] we have $\lim_{\tau\searrow 0}|f(x+\tau u)-f(x)|/\tau=\max\{\langle u,v\rangle:v\in\partial f(x)\}$.
+Since $f$ is Lipschitz continuous near $x$, by [44, Theorem 9.13] $\partial f(x)$ is nonempty and compact, and $\mathrm{lip}f(x)=\max\{|v|:v\in\partial f(x)\}$.
+Since $f$ is Lipschitz continuous and regular near $x$, by [44, Theorem 9.16] we have $\lim_{\tau\searrow 0}|f(x+\tau u)-f(x)|/\tau=\max\{\langle u,v\rangle:v\in\partial f(x)\}$.
 Thus
 $\max_{|u|=1}\lim_{\tau\searrow 0}\frac{|f(x+\tau u)-f(x)|}{\tau}=\max_{|u|=1}\max_{v\in\partial f(x)}\langle u,v\rangle=\max_{v\in\partial f(x)}\max_{|u|=1}\langle u,v\rangle=\max_{v\in\partial f(x)}|v|=\mathrm{lip}f(x).$
-Proposition 3.
+Proposition 3.15.
 If $f:\mathbb{R}^{n}\to\mathbb{R}$ is regular near $x\in\mathbb{R}^{n}$ and $\mathrm{lip}f(x)<\infty$, then $\accentset{\circ}{f}(x,r)=\mathrm{lip}f(x)r+o(r)$.
@@
 On the other hand, let $\overline{u}\in\arg\max_{|u|=1}\lim_{\tau\searrow 0}|f(x+\tau u)-f(x)|/\tau$.
-By Lemma˜2, we have $\accentset{\circ}{f}(x,r)\geq|f(x+r\overline{u})-f(x)|=\mathrm{lip}f(x)r+o(r)$ for $r\geq 0$ near 0.
+By Lemma˜3.14, we have $\accentset{\circ}{f}(x,r)\geq|f(x+r\overline{u})-f(x)|=\mathrm{lip}f(x)r+o(r)$ for $r\geq 0$ near 0.
 ∎
-Without regularity, Proposition˜3 may fail.
-Example 1.
+Without regularity, Proposition˜3.15 may fail.
+Example 3.16.
 If $f(x,y)=\min\{x^{2},|y|\}$, then $\mathrm{lip}f(0,0)=1$ and yet $\accentset{\circ}{f}((0,0),r)=(\sqrt{1+4r^{2}}-1)/2=r^{2}+o(r^{2})\neq r+o(r)$.
 We next turn to the differentiable case.
-Proposition 4.
+Proposition 3.17.
 If $f:\mathbb{R}^{n}\to\mathbb{R}$ is $D^{k}$ at $x\in\mathbb{R}^{n}$ and $f^{(i)}(x)=0$ for all $i\in\llbracket 1,k-1\rrbracket$, then $\accentset{\circ}{f}(x,r)=\|f^{(k)}(x)\|r^{k}/k!+o(r^{k})$.
@@
 ∎
-Proposition˜3 and Proposition˜4 admit the following corollaries.
-Corollary 1.
+Proposition˜3.15 and Proposition˜3.17 admit the following corollaries.
+Corollary 3.18.
 Suppose $f:\mathbb{R}^{n}\to\mathbb{R}$ is regular near $\overline{x}\in\mathbb{R}^{n}$ and $\mathrm{lip}f(x)<\infty$ for all $x\in[f=f(\overline{x})]$ near $\overline{x}$.
@@
 and in particular $\accentset{\circ}{f}(\overline{x},r)/r\leq\accentset{\circ}{f}(x,r)/r$ for all $r\in(0,\overline{r}]$.
-Passing to the limit as $r\searrow 0$ yields $\mathrm{lip}f(\overline{x})\leq\mathrm{lip}f(x)$ by Proposition˜3.
+Passing to the limit as $r\searrow 0$ yields $\mathrm{lip}f(\overline{x})\leq\mathrm{lip}f(x)$ by Proposition˜3.15.
 (ii) There exists a neighborhood $U$ of $\overline{x}$ in $f=f(\overline{x})]$ such that $\mathrm{lip}f(\overline{x})<\mathrm{lip}f(x)$ for all $x\in U\setminus\{\overline{x}\}$.
-By Proposition˜3, $\accentset{\circ}{f}(x,r)=\mathrm{lip}f(x)r+o(r)$ for all $x\in U$ after possibly reducing $U$.
+By Proposition˜3.15, $\accentset{\circ}{f}(x,r)=\mathrm{lip}f(x)r+o(r)$ for all $x\in U$ after possibly reducing $U$.
 Thus, for all $x\in U\setminus\{\overline{x}\}$, there exists $\overline{r}>0$ such that $\accentset{\circ}{f}(\overline{x},r)<\accentset{\circ}{f}(x,r)$ for all $r\in(0,\overline{r}]$.
 ∎
-Corollary 2.
+Corollary 3.19.
 Suppose $f:\mathbb{R}^{n}\to\mathbb{R}$ is $D^{k}$ near $\overline{x}\in\mathbb{R}^{n}$ and for all $x\in[f=f(\overline{x})]$ near $\overline{x}$, $f^{(i)}(x)=0$ for all $i\in\llbracket 1,k-1\rrbracket$.
@@
 Proof.
-One argues as for Corollary˜2 using Proposition˜4. ∎
+One argues as for Corollary˜3.19 using Proposition˜3.17. ∎
 When $\|f^{(i)}\|$ are merely constant on $f=f(\overline{x})]$, it is neither necessary nor sufficient for $\overline{x}$ to be flat that it be a local or strict local minimum of $\|f^{(k)}\|+\delta_{[f=f(\overline{x})]}$.
-Example 2.
+Example 3.20.
 All the minima of $f(x)=x_{1}^{2}+x_{2}^{4}(1+x_{3}^{2})$ are flat even though $\|f^{(1)}(0,0,x_{3})\|=0$, $\|f^{(2)}(0,0,x_{3})\|=2$, $\|f^{(3)}(0,0,x_{3})\|=0$, and $\|f^{(4)}(0,0,x_{3})\|=24(1+x_{3}^{2})$.
@@
 For all $x_{3}\in\mathbb{R}$ and $r\in[0,1/(1+x_{3}^{2})]$, $\accentset{\circ}{f}(0,0,x_{3},r)=r^{2}$. ∎
-Example 3.
+Example 3.21.
 $(0,0,1)$ is the unique flat minimum of $f(x)=x_{1}^{2}+x_{2}^{4}(1+x_{3}^{2})+x_{1}^{6}(1+(x_{3}-1)^{2})$ yet $\|f^{(1)}(0,0,x_{3})\|=0$, $\|f^{(2)}(0,0,x_{3})\|=2$, $\|f^{(3)}(0,0,x_{3})\|=0$, and $\|f^{(4)}(0,0,x_{3})\|=24(1+x_{3}^{2})$.
@@
 Notwithstanding, there is one situation where local optimality can be sufficient for flatness.
-Fact 9.
+Fact 3.22.
 If $f:\mathbb{R}^{n}\to\mathbb{R}$ is invariant under the natural action of a Lie subgroup $G$ of $\mathrm{O}(n)$, then $\accentset{\circ}{f}(x,r)=\accentset{\circ}{f}(gx,r)$ for $(x,r,g)\in\mathbb{R}^{n}\times\mathbb{R}_{+}\times G$.
@@
 $\accentset{\circ}{f}(x,r)=\sup_{y\in\overline{B}_{r}(x)}|f(y)-f(x)|=\sup_{gy\in\overline{B}_{r}(gx)}|f(gy)-f(gx)|=\sup_{z\in\overline{B}_{r}(gx)}|f(z)-f(gx)|=\accentset{\circ}{f}(gx,r)$. ∎
-Corollary 3.
+Corollary 3.23.
 Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz regular and invariant under the natural action of a Lie subgroup $G$ of $\mathrm{O}(n)$.
@@
 Proof.
-˜9 implies that $x\preceq y$ and $y\preceq x$ for all $x,y\in G\overline{x}$.
-By Proposition˜3, $\accentset{\circ}{f}(x,r)=\mathrm{lip}f(x)r+o(r)$ for all $x\in\mathbb{R}^{n}$.
+Fact˜3.22 implies that $x\preceq y$ and $y\preceq x$ for all $x,y\in G\overline{x}$.
+By Proposition˜3.15, $\accentset{\circ}{f}(x,r)=\mathrm{lip}f(x)r+o(r)$ for all $x\in\mathbb{R}^{n}$.
 There exists a neighborhood $U$ of $G\overline{x}$ in $f=f(\overline{x})]$ such that $\mathrm{lip}f(x)<\mathrm{lip}f(y)$ for all $x\in G\overline{x}$ and $y\in U\setminus G\overline{x}$.
@@
 An analogous result holds in the differentiable case.
-Corollary 4.
+Corollary 3.24.
 Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be $D^{k}$ and invariant under the natural action of a Lie subgroup $G$ of $\mathrm{O}(n)$.
@@
 When applying the calculus rules, a composite structure can help.
-Fact 10.
-[36, Theorem 10.6] Let $f=g\circ F$ where $g:\mathbb{R}^{m}\to\mathbb{R}$ is locally Lipschitz regular and $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is $C^{1}$. Then $f$ is regular and
+Fact 3.25.
+[44, Theorem 10.6] Let $f=g\circ F$ where $g:\mathbb{R}^{m}\to\mathbb{R}$ is locally Lipschitz regular and $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is $C^{1}$. Then $f$ is regular and
 $\forall x\in\mathbb{R}^{n},~~~\partial f(x)=F^{\prime}(x)^{*}\partial g(F(x)).$
 In particular, if $F(x)=0$ and $g=|\cdot|_{1}$, then $\mathrm{lip}f(x)=\|F^{\prime}(x)^{*}\|_{\infty,2}$.
-Fact 11.
+Fact 3.26.
 Let $f=g\circ F$ where $g:\mathbb{R}^{m}\to\mathbb{R}$ and $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$ are $D^{k}$, and $g^{(i)}(0)=0$ for all $i\in\llbracket 1,k-1\rrbracket$.
@@
 Sometimes the composite structure helps with subdifferentiation, but not with determining flat minima.
-Remark 1.
+Remark 3.27.
 Let $f=g\circ F$ where $g:\mathbb{R}^{m}\to\mathbb{R}$ is $C^{1}$ such that $\nabla g(0)=0$ and $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is locally Lipschitz.
-The Jacobian chain rule [11, Theorem 2.6.6] yields $\overline{\partial}f(x)=\overline{\partial}F(x)^{T}\nabla g(F(x))$.
+The Jacobian chain rule [12, Theorem 2.6.6] yields $\overline{\partial}f(x)=\overline{\partial}F(x)^{T}\nabla g(F(x))$.
 Hence, if $F(x)=0$, then $\overline{\partial}f(x)=\{0\}$.
-Since the Clarke subdifferential is a singleton, by [36, Theorem 9.18] $f$ is strictly differentiable [36, Definition 9.17] at $x$, $\nabla f(x)=0$, and $\mathrm{lip}f(x)=0$.
+Since the Clarke subdifferential is a singleton, by [44, Theorem 9.18] $f$ is strictly differentiable [44, Definition 9.17] at $x$, $\nabla f(x)=0$, and $\mathrm{lip}f(x)=0$.
 Two additional rules are useful in the presence of symmetries.
-Fact 12.
-Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be lower semicontinuous and invariant under the natural action of a Lie subgroup of $\mathrm{GL}(n,\mathbb{R})$.
+Fact 3.28.
+Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be lower semicontinuous and invariant under the natural action of a Lie subgroup of $\mathrm{GL}(n)$.
 For all $\overline{x}\in\mathrm{dom}f$ and $g\in G$, $\partial f(g^{-1}\overline{x})=g^{T}\partial f(\overline{x})$.
 Proof.
-Applying the chain rule [36, Exercise 10.7] to $f(x)=f(gx)$ at $g^{-1}\overline{x}$ for any $g\in G\subseteq\mathrm{GL}(n,\mathbb{R})$ yields the result. ∎
-Fact 13.
-Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be $D^{k}$ and invariant under the natural action of a Lie subgroup of $\mathrm{GL}(n,\mathbb{R})$.
+Applying the chain rule [44, Exercise 10.7] to $f(x)=f(gx)$ at $g^{-1}\overline{x}$ for any $g\in G\subseteq\mathrm{GL}(n,\mathbb{R})$ yields the result. ∎
+Fact 3.29.
+Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be $D^{k}$ and invariant under the natural action of a Lie subgroup of $\mathrm{GL}(n)$.
 For all $\overline{x},v_{1},\ldots,v_{k}\in\mathbb{R}^{n}$, $f^{(k)}(g^{-1}\overline{x})(v_{1},\ldots,v_{k})=f^{(k)}(\overline{x})(gv_{1},\ldots,gv_{k})$.
@@
 A curve is a function $x:I\to\mathbb{R}$ where $I$ is an interval of $\mathbb{R}$.
-We refer to trajectories as curves that solve an ODE.
-We say that a curve $x:I\to\mathbb{R}$ is flattening, or flattens over time, if $x(t)\prec x(s)$ for all $s,t\in I$ such that $s<t$.
-A curve $x:I\to\mathbb{R}$ sharpens over time if $I\ni t\mapsto x(-t)\in\mathbb{R}^{n}$ flattens over time.
-We seek to construct such curves using the following assumption.
-Assumption 2.
+Definition 4.1.
+A trajectory of $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ is an absolutely continuous curve $x:I\to\mathbb{R}^{n}$ where $I$ is an interval of $\mathbb{R}$ such that
+[…]
+If $F$ is singleton-valued and continuous, then its trajectories are $C^{1}$.
+We refer to solutions to the differential inclusion $\dot{x}\in F(x)$ as trajectories of $F$ over an interval $I=[0,T)$ for some $T\in(0,\infty]$ or $I=[0,T]$ for some $T\in(0,\infty)$.
+A solution $x:I\to\mathbb{R}^{n}$ is maximal if for any other solution $y:J\to\mathbb{R}^{n}$ such that $I\subseteq J$ and $x(t)=y(t)$ for all $t\in I$, we have $I=J$.
+A solution is globally defined if $I=[0,\infty)$.
+Definition 4.2.
+Given $f:\mathbb{R}^{n}\to\mathbb{R}$, a curve $x:I\to\mathbb{R}^{n}$ is flattening, or flattens over time (resp.
+strictly), if $x(t)\preceq x(s)$ (resp.
+$x(t)\prec x(s)$) for all $s,t\in I$ such that $s<t$.
+A curve $x:I\to\mathbb{R}$ sharpens (resp.
+strictly) over time if $I\ni t\mapsto x(-t)\in\mathbb{R}^{n}$ flattens over time (resp.
+strictly).
+We seek to construct such flattening curves using the following assumption. It is satisfied in Example˜5.6, [25, Examples 4.4, 4.5], and [27, Examples 4.2, 4.3, …, 4.8].
+Assumption 4.3.
 Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz definable and $c:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be $C^{4}$ on a bounded convex open set $U\subseteq\mathbb{R}^{n}$ such that
-$\exists\omega>0:~\forall x\in U,~\forall v\in\overline{\partial}f(x),~~~\langle\nabla c(x),v\rangle=0~~\text{and}~~~\langle\nabla^{2}c(x)v,v\rangle\leq-\omega|v|^{2}.$
-The first-order condition in ˜2 induces conservation laws.
-Proposition 5.
+$\exists\omega>0:~\forall x\in U,~\forall v\in\overline{\nabla}f(x),~~~\langle\nabla c(x),v\rangle=0~~\text{and}~~~\langle\nabla^{2}c(x)v,v\rangle\leq-\omega|v|^{2}.$
+The first-order condition in Assumption˜4.3 induces conservation laws.
+Proposition 4.4.
 Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz definable and $c:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be $D^{1}$ on an open set $U\subseteq\mathbb{R}^{n}$ such that $\langle\nabla c(x),v\rangle=0$ for all $x\in U$ and $v\in\overline{\partial}f(x)$.
   1. (i)
-    If $x:I\to U$ is a solution to $\dot{x}\in-\overline{\partial}f(x)$ almost everywhere, then $c$ is conserved.
+    If $x:I\to U$ is a trajectory of $\overline{\partial}f$, then $c$ is conserved.
   2. (ii)
-    If $x:I\to U$ is a solution to $\dot{x}=\nabla c(x)$, then $f$ is conserved.
+    If $x:I\to U$ is a trajectory of $\nabla c$, then $f$ is conserved.
 Proof.
@@
 Thus, for almost every $t\in I$, $(c\circ x)^{\prime}(t)=\langle\nabla c(x),x^{\prime}(t)\rangle=0$ and $I\ni t\to c(x(t))$ is constant.
-(ii) Since $f$ is definable and Lipschitz continuous, by [9, Corollary 2, Proposition 2] $f$ is path differentiable. Thus, for almost every $t\in I$,
+(ii) Since $f$ is definable and Lipschitz continuous, by the chain rule in [16, Corollary 5.4], for almost every $t\in I$,
 $\forall v\in\overline{\partial}f(x(t)),~~~(f\circ x)^{\prime}(t)=\langle v,x^{\prime}(t)\rangle=\langle v,\nabla c(x(t))\rangle=0$
 and $I\ni t\to f(x(t))$ is constant. ∎
-Under ˜2, gradient trajectories of the conserved quantity $c$ sharpen over time. This is because the level sets are contracting. To see why, we begin with a simple lemma.
-Lemma 3.
+Under Assumption˜4.3, gradient trajectories of the conserved quantity $c$ strictly sharpen over time. This is because the level sets are contracting. To see why, we begin with a simple lemma.
+Lemma 4.5.
 If $c:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is $C^{4}$ on a bounded convex open set $U\subseteq\mathbb{R}^{n}$, then there exists $M>0$ such that
@@
 Since $c$ is $C^{4}$ on $U$, $g$ is $C^{3}$ on $U$.
-By [34, Lemma 1.2.4], there exists $M>0$ such that for all $x,y\in U$, $u\in U+\{-x\}$, and $v\in U+\{-y\}$ we have
+By [39, Lemma 1.2.4], there exists $M>0$ such that for all $x,y\in U$, $u\in U+\{-x\}$, and $v\in U+\{-y\}$ we have
 $\left|g(x+u,y+v)-g(x,y)-\langle\nabla g(x,y),(u,v)\rangle-\langle\nabla^{2}g(x,y)(u,v),(u,v)\rangle/2\right|\leq M(|u|^{3}+|v|^{3}).$
@@
 The following result is one of the main findings of this paper.
-Theorem 1.
-Under ˜2, for all $x_{0}\in U$ near which $f$ is nonconstant, there exists a local solution $x:[0,\overline{t})\to\mathbb{R}^{n}$ to
+Theorem 4.6.
+Under Assumption˜4.3, for all $x_{0}\in U$ near which $f$ is nonconstant, there exists a solution $x:[0,\overline{t})\to\mathbb{R}^{n}$ to
 $\left\{\begin{array}[]{ccc}\dot{x}&=&\nabla c(x)\\ x(0)&=&x_{0}\end{array}\right$.
@@
 Proof.
-Since $\nabla c$ is Lipschitz continuous on $U$, by [27, Theorem D.1], there exist $\overline{t},\overline{r}>0$ such that the ODE
+Since $\nabla c$ is Lipschitz continuous on $U$, by [32, Theorem D.1], there exist $\overline{t},\overline{r}>0$ such that the ODE
 $\left\{\begin{array}[]{ccc}\dot{z}&=&\nabla c(z)\\ z(0)&=&z_{0}\end{array}\right$.
 admits a unique $C^{1}$ solution $z:[0,\overline{t})\to U$ for all $z_{0}\in B_{\overline{r}}(x_{0})$.
-Since $\nabla c$ is $C^{2}$, by [27, Theorem D.1], the semiflow $\theta:[0,\overline{t})\times B_{\overline{r}}(x_{0})\to\mathbb{R}^{n}$ defined by $\theta(t,z_{0})=z(t)$ is $C^{2}$.
-By [34, Lemma 1.2.4], there exists $C>0$ such that for all $t,\widetilde{t}\in[0,\overline{t})$ and $z_{0},\widetilde{z}_{0}\in B_{\overline{r}}(x_{0})$, we have
+Since $\nabla c$ is $C^{2}$, by [32, Theorem D.1], the semiflow $\theta:[0,\overline{t})\times B_{\overline{r}}(x_{0})\to\mathbb{R}^{n}$ defined by $\theta(t,z_{0})=z(t)$ is $C^{2}$.
+By [39, Lemma 1.2.4], there exists $C>0$ such that for all $t,\widetilde{t}\in[0,\overline{t})$ and $z_{0},\widetilde{z}_{0}\in B_{\overline{r}}(x_{0})$, we have
 $|\theta(t,z_{0})-\theta(\widetilde{t},\widetilde{z}_{0})-\theta^{\prime}(\widetilde{t},\widetilde{z}_{0})(t-\widetilde{t},z_{0}-\widetilde{z}_{0})|\leq C\left[(t-\widetilde{t})^{2}+|z_{0}-\widetilde{z}_{0}|^{2}\right].$ (1)
-Since $f$ is locally Lipschitz definable and nonconstant near $x_{0}$, by Lemma˜1 there exist definable curves $\gamma:[0,\overline{r}]\to\mathbb{R}^{n}$, $\lambda:[0,\overline{r}]\to\mathbb{R}$, and $v:[0,\overline{r}]\to\mathbb{R}^{n}$ such that $\gamma$ is $C^{1}$ on $0,\overline{r}]$ and
-$\forall r\in[0,\overline{r}],~~~\gamma(r)=x_{0}+\lambda(r)v(r)\in\arg\max_{\overline{B}_{r}(x_{0})}|f-f(x_{0})|~~~\text{and}~~~v(r)\in\overline{\partial}f(\gamma(r)),$
+Since $f$ is locally Lipschitz, regular, definable, and nonconstant near $x_{0}$, by Lemma˜3.13 there exist definable curves $\gamma:[0,\overline{r}]\to\mathbb{R}^{n}$, $\lambda:[0,\overline{r}]\to\mathbb{R}$, and $v:[0,\overline{r}]\to\mathbb{R}^{n}$ such that $\gamma$ is $C^{1}$ on $0,\overline{r}]$ and
+$\forall r\in[0,\overline{r}],~~\gamma(r)=x_{0}+\lambda(r)v(r)\in\arg\max_{\overline{B}_{r}(x_{0})}|f-f(x_{0})|,~~|\gamma(r)-x_{0}|=r,~~v(r)\in\overline{\nabla}f(\gamma(r)),$
 after possibly reducing $\overline{r}$. As $c$ is $C^{4}$ on $U$, there exists $L>0$ such that
 $\forall x,y\in U,~~~|\nabla^{2}c(x)-\nabla^{2}c(y)|\leq L|x-y|.$
-Let $M>0$ be given by Lemma˜3.
+Let $M>0$ be given by Lemma˜4.5.
 For all $(t,r)\in[0,\overline{t})\times[0,\overline{r})$, let $x_{r}(t)=\theta(t,\gamma(r))$ and $x(t)=x_{0}(t)=\theta(t,\overline{x})$.
@@
 It follows that $|x_{r}(t)-\gamma(r)|\leq|x_{r}(t)-x_{0}|+|x_{0}-\gamma(r)|\leq\omega/(8L)+\omega/(8L)=\omega/(4L)$ and $|x_{r}(t)-x(t)|\leq|x_{r}(t)-x_{0}|+|x_{0}-x_{0}(t)|\leq\omega/(4M)$.
-The main idea of the proof is now as follows. Using Proposition˜5, we have
+The main idea of the proof is now as follows. Using Proposition˜4.4, we have
 $\accentset{\circ}{f}(x_{0},r)=\max_{\overline{B}_{r}(x_{0})}|f-f(x_{0})|=|f(x_{r}(0))-f(x(0))|=|f(x_{r}(t))-f(x(t))|\leq\accentset{\circ}{f}(x(t),e^{-\omega t/2}r)$
@@
 Initially $y_{r}(0)=x_{r}(0)-x(0)=\gamma(r)-x_{0}=\lambda(r)v(r)\neq 0$ for all $r\in(0,\overline{r})$.
-Since $v(r)\in\overline{\partial}f(\gamma(r))$, we have $\langle\nabla^{2}c(\gamma(r))v(r),v(r)\rangle\leq-\omega|v(r)|^{2}$.
+Since $v(r)\in\overline{\nabla}f(\gamma(r))$, by Assumption˜4.3 we have $\langle\nabla^{2}c(\gamma(r))v(r),v(r)\rangle\leq-\omega|v(r)|^{2}$.
 Together with $\lambda(r)\neq 0$, we find $\langle\nabla^{2}c(\gamma(r))\lambda(r)v(r),\lambda(r)v(r)\rangle\leq-\omega|\lambda(r)v(r)|^{2}$.
@@
 Each entry of $(\gamma-x_{0})/|\gamma-x_{0}|$ is definable and bounded.
-By the monotonicity theorem [40, 4.1], it is monotone near 0 and thus convergent.
+By the monotonicity theorem [48, 4.1], it is monotone near 0 and thus convergent.
 Thus there exists $u\in S^{n-1}$ such that $(\gamma(r)-x_{0})/|\gamma(r)-x_{0}|\to u$ as $r\searrow 0$.
@@
 $\leq-\omega|y_{r}|^{2}.$
-The ODE comparison theorem [27, Theorem D.2] implies that $|y_{r}(t)|^{2}\leq e^{-\omega t}|y_{r}(0)|^{2}$, namely, $|x_{r}(t)-x(t)|\leq e^{-\omega t/2}|x_{r}(0)-x(0)|$ for all $t\in[0,\overline{t})$ and $r\in[0,\overline{r})$.
-As for the equivalence, it is obtained via Lemma˜1 and the change of variables $\ell=\accentset{\circ}{f}(x_{0},r)$: $\accentset{\circ}{f}(x_{0},r)\leq\accentset{\circ}{f}(x(t),e^{-\omega t/2}r)$ iff $\overline{f}(x(t),\accentset{\circ}{f}(x_{0},r))\leq\overline{f}(x(t),\accentset{\circ}{f}(x(t),e^{-\omega t/2}r))$ iff $\overline{f}(x(t),\ell)\leq e^{-\omega t/2}\overline{f}(x_{0},\ell)$.
+The ODE comparison theorem [32, Theorem D.2] implies that $|y_{r}(t)|^{2}\leq e^{-\omega t}|y_{r}(0)|^{2}$, namely, $|x_{r}(t)-x(t)|\leq e^{-\omega t/2}|x_{r}(0)-x(0)|$ for all $t\in[0,\overline{t})$ and $r\in[0,\overline{r})$.
+As for the equivalence, it is obtained via Lemma˜3.13 and the change of variables $\ell=\accentset{\circ}{f}(x_{0},r)$: $\accentset{\circ}{f}(x_{0},r)\leq\accentset{\circ}{f}(x(t),e^{-\omega t/2}r)$ iff $\overline{f}(x(t),\accentset{\circ}{f}(x_{0},r))\leq\overline{f}(x(t),\accentset{\circ}{f}(x(t),e^{-\omega t/2}r))$ iff $\overline{f}(x(t),\ell)\leq e^{-\omega t/2}\overline{f}(x_{0},\ell)$.
 ∎
-Due to the local nature of Theorem˜1, it does not guarantee that reversing time yields flattening trajectories from a given initial point.
+Due to the local nature of Theorem˜4.6, it does not guarantee that reversing time yields strictly flattening trajectories from a given initial point.
 In order to do so, we will add some assumptions and rely on the following fact.
-Fact 14.
+Fact 4.7.
 Let $g:(a,b)\to\mathbb{R}$ be continuous such that for all $s\in(a,b)\subseteq\mathbb{R}$, there exists $s^{\prime}\in(s,b)$ such that $g(t)\geq e^{t-s}g(s)$ for all $t\in(s,s^{\prime})$.
@@
 ∎
-Corollary 5.
-Under ˜2, let $x:I\to U$ be such that $\dot{x}=-\nabla c(x)$ on a compact interval $I$ of $\mathbb{R}$ containing $t_{0}$. Let $s,t\in I$ be such that $s<t$.
+Corollary 4.8.
+Under Assumption˜4.3, let $x:I\to U$ be such that $\dot{x}=-\nabla c(x)$ on a compact interval $I$ of $\mathbb{R}$ containing $t_{0}$. Let $s,t\in I$ be such that $s<t$.
   1. (i)
@@
 Proof.
-(i) Let $t\in I$.
-By Theorem˜1 and Proposition˜3, $\mathrm{lip}f(x(t))+o(1)=\accentset{\circ}{f}(x(t),r)/r\leq\accentset{\circ}{f}(x(s),e^{-\omega(t-s)/2}r)/r=e^{-\omega(t-s)/2}\mathrm{lip}f(x(s))+o(1)$ for all $s\in I\cap(-\infty,t]$ sufficiently close to $t$.
-˜14 enables one to extend it to any $s\in I\cap(-\infty,t]$.
-(ii) Assuming the existence of a solution $x:I\to U$ to the ODE passing through $x_{0}=x(t_{0})$ at time $t_{0}$ implies the existence of solutions on an open interval $J\supset I$ for initial points in a neighborhood $U_{0}$ of $x_{0}$.
-This holds because $\nabla c$ is Lipschitz continuous on $U$, which also implies uniqueness of solutions (using the ODE comparison theorem [27, Theorem D.2]).
+(i) Let $t\in I$. By Theorem˜4.6 and Proposition˜3.15,
+$\mathrm{lip}f(x(t))+o(1)=\accentset{\circ}{f}(x(t),r)/r\leq\accentset{\circ}{f}(x(s),e^{-\omega(t-s)/2}r)/r=e^{-\omega(t-s)/2}\mathrm{lip}f(x(s))+o(1)$
+for all $s\in I\cap(-\infty,t]$ sufficiently close to $t$. Fact˜4.7 enables one to extend it to any $s\in I\cap(-\infty,t]$.
+(ii) Assuming the existence of a solution $x:I\to U$ to the ODE passing through $x_{0}=x(t_{0})$ at time $t_{0}$ implies the existence of solutions on an open interval $J\supseteq I$ for initial points in a neighborhood $U_{0}$ of $x_{0}$.
+This holds because $\nabla c$ is Lipschitz continuous on $U$, which also implies uniqueness of solutions (using the ODE comparison theorem [32, Theorem D.2]).
 Thus the flow $\theta:J\times U_{0}\to U$ is well defined and $C^{k}$.
-For any $t\in J$, $\theta_{t}=\theta(t,\cdot)$ defines a diffeomorphism from $U_{0}$ to $\theta_{t}(U_{0})$ [27, p.
+For any $t\in J$, $\theta_{t}=\theta(t,\cdot)$ defines a diffeomorphism from $U_{0}$ to $\theta_{t}(U_{0})$ [32, p.
 209], which is a neighborhood of $x(t)$.
@@
 By the chain rule, $(f\circ\theta_{t})^{\prime}(x)=f^{\prime}(\theta_{t}(x))\circ\theta_{t}^{\prime}(x)$.
-In particular, $0=f^{\prime}(x_{0})(v)=(f\circ\theta_{t})^{\prime}(x_{0})(v)=f^{\prime}(\theta_{t}(x_{0}))(\theta_{t}^{\prime}(x_{0})v)$ for all $v\in\mathbb{R}^{n}$, so that $f^{\prime}(\theta_{t}(x_{0}))=0$.
+In particular,
+$\forall v\in R^{n},~~~0=f^{\prime}(x_{0})(v)=(f\circ\theta_{t})^{\prime}(x_{0})(v)=f^{\prime}(\theta_{t}(x_{0}))(\theta_{t}^{\prime}(x_{0})v),$
+so that $f^{\prime}(\theta_{t}(x_{0}))=0$.
 By induction, $(f\circ\theta_{t})^{(i)}(x_{0})v^{i}=f^{(i)}(\theta_{t}(x_{0}))(\theta_{t}^{\prime}(x_{0})v)^{i}$ for all $i\in\llbracket 1,k\rrbracket$ and thus $f^{(i)}(x(t))=0$ for all $i\in\llbracket 1,k-1\rrbracket$.
-It follows that $\|f^{(k)}(x(t))\|/k!+o(1)=\accentset{\circ}{f}(x(t),r)/r^{k}\leq\accentset{\circ}{f}(x(s),e^{-\omega(t-s)/2}r)/r^{k}=e^{-k\omega(t-s)/2}\|f^{(k)}(x(s))\|/k!+o(1)$ for all $s\in I\cap(-\infty,t]$ sufficiently close to $t$ by Theorem˜1 and Proposition˜4.
-One again concludes by ˜14.
-∎
-Remark 2.
-One can actually weaken the equality $\langle\nabla c(x),v\rangle=0$ for all $v\in\overline{\partial}f(x)$ in ˜2 to an inequality $\langle\nabla c(x),v\rangle\geq 0$.
-In exchange, one must assume that $x(\cdot)$ is defined on $0,\overline{t}]$ (at no cost) and $x(\overline{t})\in\arg\min\{f(x):x\in U\}$ in Theorem˜1.
+It follows that
+$\|f^{(k)}(x(t))\|/k!+o(1)$ $=\accentset{\circ}{f}(x(t),r)/r^{k}\leq\accentset{\circ}{f}(x(s),e^{-\omega(t-s)/2}r)/r^{k}$
+$=e^{-k\omega(t-s)/2}\|f^{(k)}(x(s))\|/k!+o(1)$
+for all $s\in I\cap(-\infty,t]$ sufficiently close to $t$ by Theorem˜4.6 and Proposition˜3.17. One again concludes by Fact˜4.7. ∎
+Remark 4.9.
+One can actually weaken the equality $\langle\nabla c(x),v\rangle=0$ for all $v\in\overline{\nabla}f(x)$ in Assumption˜4.3 to an inequality $\langle\nabla c(x),v\rangle\geq 0$.
+In exchange, one must assume that $x(\cdot)$ is defined on $0,\overline{t}]$ (at no cost) and $x(\overline{t})\in\arg\min\{f(x):x\in U\}$ in Theorem˜4.6.
 In the proof, $f$ instead increases along the trajectories of $\nabla c$, but is still constant on $x(\cdot)$, so that $|f(x_{r}(0))-f(x(0))|\leq|f(x_{r}(t))-f(x(t))|$.
-Similarly, one must assume that $x(t_{0})\in\arg\min\{f(x):x\in U\}$ in Corollary˜5.
+Similarly, one must assume that $x(t_{0})\in\arg\min\{f(x):x\in U\}$ in Corollary˜4.8.
 In the proof, instead of the chain rule, one uses Taylor expansions: $f(y)=f(x_{0})+\langle\nabla f(x_{0}),y-x_{0}\rangle+o(|y-x_{0}|)$ and $f(\theta_{t}(y))=f(\theta_{t}(x_{0}))+\langle\nabla f(\theta_{t}(x_{0})),\theta_{t}(y)-\theta_{t}(x_{0})\rangle+o(|\theta_{t}(y)-\theta_{t}(x_{0})|)$.
@@
 The rest follows by induction.
-With these new assumptions, it is sufficient to satisfy the second inequality in ˜2, i.e., $\langle\nabla^{2}c(x)v,v\rangle\leq-\omega|v|^{2}$ for all $v\in\overline{\partial}f(x)$ and $x\in U$, merely for all $x\in U\setminus\arg\min\{f(x):x\in U\}$.
+With these new assumptions, it is sufficient to satisfy the second inequality in Assumption˜4.3, i.e., $\langle\nabla^{2}c(x)v,v\rangle\leq-\omega|v|^{2}$ for all $v\in\overline{\nabla}f(x)$ and $x\in U$, merely for all $x\in U\setminus\arg\min\{f(x):x\in U\}$.
 4.2 Linear symmetries
 Linear symmetries give rise to a conservation law in subgradient dynamics, as shown in [26].
-We next recall it and compute the Hessian of the conserved quantity in some directions, since it appears in ˜2.
-Note that a conservation law in gradient dynamics was proposed earlier in [43, Proposition 5.1]
-Assumption 3.
+We next recall it and compute the Hessian of the conserved quantity in some directions, since it appears in Assumption˜4.3.
+Note that a conservation law in gradient dynamics was proposed earlier in [51, Proposition 5.1]
+Assumption 4.10.
 Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be locally Lipschitz and invariant under the natural action of a Lie subgroup $G$ of $\mathrm{GL}(n,\mathbb{R})$, $C(x)=P_{\mathrm{s}(\mathfrak{g})}(xx^{T})$, and $\overline{x}\in\mathbb{R}^{n}$.
-Proposition 6.
-Let ˜3 hold and $c(x)=\|C(x)-C(\overline{x})\|_{F}^{2}/4$. Then
+Proposition 4.11.
+Let Assumption˜4.10 hold and $c(x)=\|C(x)-C(\overline{x})\|_{F}^{2}/4$ for some $\overline{x}\in\mathbb{R}^{n}$. Then
 $\forall x\in\mathbb{R}^{n},~\forall v\in\overline{\partial}f(x),~~~\langle\nabla c(x),v\rangle=0~~~\text{and}~~~\langle\nabla^{2}c(x)v,v\rangle=\langle C(x)-C(\overline{x}),C(v)\rangle_{F}.$
 Proof.
-By [23, Corollary 1], for all $x\in\mathbb{R}^{n}$, $v\in\overline{\partial}f(x)$, and $\alpha\in\mathbb{R}$, we have $C(x+\alpha v)=C(x)+\alpha^{2}C(v)$ and thus
+By [26, Corollary 5.1], for all $x\in\mathbb{R}^{n}$, $v\in\overline{\partial}f(x)$, and $\alpha\in\mathbb{R}$, we have $C(x+\alpha v)=C(x)+\alpha^{2}C(v)$ and thus
 $c(x+\alpha v)$ $=\|C(x+\alpha v)-C(\overline{x})\|_{F}^{2}/4=\|C(x)+\alpha^{2}C(v)-C(\overline{x})\|_{F}^{2}/4$
@@
 Our focus is now on deriving necessary conditions for flatness using the conserved quantity. The following result can be seen as a warm up. It will be useful later.
-Proposition 7.
-Under ˜3, $\overline{x}\in\arg\mathrm{loc}\min\{|x|:x\in G\overline{x}\}\Longrightarrow C(\overline{x})=0$.
+Proposition 4.12.
+Under Assumption˜4.10, $\overline{x}\in\arg\mathrm{loc}\min\{|x|:x\in G\overline{x}\}\Longrightarrow C(\overline{x})=0$.
 Proof.
@@
 We have arrived at our second main result.
-Theorem 2.
-Under ˜3,
+Theorem 4.13.
+Under Assumption˜4.10,
 $\overline{x}\in\arg\mathrm{loc}\min\{\mathrm{lip}f(x):x\in G\overline{x}\}~\Longrightarrow~\exists\overline{v}\in\arg\max\{|v|:v\in\partial f(\overline{x})\}:~C(\overline{v})=0.$
@@
 $\inf_{g\in G}~\mathrm{lip}f(g^{-1}\overline{x})^{2}=\inf_{g\in G}\sup_{v\in\partial f(g^{-1}\overline{x})}|v|^{2}=\inf_{g\in G}\sup_{v\in\partial f(\overline{x})}|g^{T}v|^{2}=\inf_{g\in G}-\inf_{v\in\mathbb{R}^{n}}\phi(v,g)=-\sup_{g\in G}\varphi(g)$
-where the second equality holds by ˜12. Also, $\phi:\mathbb{R}^{n}\times G\rightarrow\overline{\mathbb{R}}$ and $\varphi:G\rightarrow\mathbb{R}_{-}$ are defined by
+where the second equality holds by Fact˜3.28. Also, $\phi:\mathbb{R}^{n}\times G\rightarrow\overline{\mathbb{R}}$ and $\varphi:G\rightarrow\mathbb{R}_{-}$ are defined by
 $\phi(v,g)=\delta_{\partial f(\overline{x})}(v)-|g^{T}v|^{2}~~~\text{and}~~~\varphi(g)=\inf_{v\in\mathbb{R}^{n}}\phi(v,g).$
@@
 Thus $\varphi\circ\gamma$ is real-valued on $\epsilon/2,\epsilon/2]$.
-Since 0 is a local maximum of $\varphi\circ\gamma$, we have $0\in\partial(\varphi\circ\gamma)(0)$ by Fermat’s rule [36, Theorem 10.1].
-As the sum of the indicator of a closed convex set and a smooth function, $\tau$ is regular [36, Example 7.28].
-[36, Corollary 10.11] then implies that for all $(v,t)\in\mathrm{dom}\tau$, one has
+Since 0 is a local maximum of $\varphi\circ\gamma$, we have $0\in\partial(\varphi\circ\gamma)(0)$ by Fermat’s rule [44, Theorem 10.1].
+As the sum of the indicator of a closed convex set and a smooth function, $\tau$ is regular [44, Example 7.28].
+[44, Corollary 10.11] then implies that for all $(v,t)\in\mathrm{dom}\tau$, one has
 $\partial\tau(v,t)\subseteq\begin{pmatrix}N_{\partial f(\overline{x})}(v)-2\gamma(t)\gamma(t)^{T}v\\ N_{[-\epsilon/2,\epsilon/2]}(t)-2\langle\gamma(t)^{T}v,\gamma^{\prime}(t)^{T}v\rangle\end{pmatrix}$.
-Due to its bounded domain, the function $\tau(v,t)$ is level-bounded in $v$ locally uniformly in $t$ [36, 1.16 Definition].
-We may thus apply [36, Theorem 10.13] on parametric subdifferentiation.
+Due to its bounded domain, the function $\tau(v,t)$ is level-bounded in $v$ locally uniformly in $t$ [44, 1.16 Definition].
+We may thus apply [44, Theorem 10.13] on parametric subdifferentiation.
 For all $\overline{t}\in[-\epsilon/2,\epsilon/2]$, it yields
@@
 ∎
-In Theorem˜2, one can also take a maximal Bouligand subdifferential by ˜1. The converse of Theorem˜2 do not always hold.
-Example 4.
-The function $f(x)=|2x_{1}^{2}-x_{2}^{2}|$ is invariant under the natural action of $G=\{A\in\mathrm{GL}(2,\mathbb{R}):A^{T}DA=D\}$ where $D=\mathrm{diag}(2,-1)$.
+In Theorem˜4.13, one can also take a maximal Bouligand subdifferential by Fact˜2.1. The converse of Theorem˜4.13 do not always hold.
+Example 4.14.
+The function $f(x)=|2x_{1}^{2}-x_{2}^{2}|$ is invariant under the natural action of $G=\{A\in\mathrm{GL}(2):A^{T}DA=D\}$ where $D=\mathrm{diag}(2,-1)$.
 The only flat minimum is $(0,0)$ even though $C(x)=P_{\mathrm{s}(\mathfrak{g})}(xx^{T})=0$ for all $x\in\mathbb{R}^{2}$.
@@
 Observe that $f(x)=\langle x,Dx\rangle$.
-For all $A\in\mathrm{GL}(2,\mathbb{R})$, $f(Ax)=\langle Ax,DAx\rangle=\langle x,A^{T}DAx\rangle=\langle x,Dx\rangle=f(x)$.
+For all $A\in\mathrm{GL}(2)$, $f(Ax)=\langle Ax,DAx\rangle=\langle x,A^{T}DAx\rangle=\langle x,Dx\rangle=f(x)$.
 Thus $f$ is invariant under the natural action of $G$, whose Lie algebra is
@@
 Thus $\mathrm{s}(\mathfrak{g})=\{0\}$ and $C(x)=P_{\mathrm{s}(\mathfrak{g})}(xx^{T})=0$ for all $x\in\mathbb{R}^{2}$.
-By ˜10, $\mathrm{lip}f(x)=2\sqrt{4x_{1}^{2}+x_{2}^{2}}$ and thus $(0,0)$ is the only flat minimum by Corollary˜1.
+By Fact˜3.25, $\mathrm{lip}f(x)=2\sqrt{4x_{1}^{2}+x_{2}^{2}}$ and thus $(0,0)$ is the only flat minimum by Corollary˜3.18.
 ∎
-The converse of Theorem˜2 does hold in $\ell_{1}$-matrix factorization, as we show in Section˜4.3.
-Corollary 6.
-Under ˜3, if $\overline{x}$ is flat and $f$ is regular near $\overline{x}$, then there exists $\overline{v}\in\arg\max\{|v|:v\in\partial f(\overline{x})\}$ such that $C(\overline{v})=0$.
+The converse of Theorem˜4.13 does hold in $\ell_{1}$-matrix factorization, as we show in Section˜4.3.
+Corollary 4.15.
+Under Assumption˜4.10, if $\overline{x}$ is flat and $f$ is regular near $\overline{x}$, then there exists $\overline{v}\in\arg\max\{|v|:v\in\partial f(\overline{x})\}$ such that $C(\overline{v})=0$.
 Proof.
-This a consequence of ˜12, Corollary˜1, and Theorem˜2. ∎
-Theorem˜2 can be generalized to higher orders, as follows.
-Theorem 3.
+This a consequence of Fact˜3.28, Corollary˜3.18, and Theorem˜4.13. ∎
+Theorem˜4.13 can be generalized to higher orders, as follows.
+Theorem 4.16.
 Suppose $f$ is $D^{k}$ near $\overline{x}\in\mathbb{R}^{n}$ and $f^{(k)}(\overline{x})\neq 0$. Then
@@
 Proof.
-The proof naturally mirrors that of Theorem˜2. By assumption, $I_{n}$ is a local solution to
+The proof naturally mirrors that of Theorem˜4.13. By assumption, $I_{n}$ is a local solution to
 $\inf_{g\in G}\|f^{(k)}(g^{-1}\overline{x})\|^{2}$ $=\inf_{g\in G}\sup_{|v|=1}(f^{(k)}(g^{-1}\overline{x})(v,\ldots,v))^{2}=\inf_{g\in G}\sup_{|v|=1}(f^{(k)}(\overline{x})(gv,\ldots,gv))^{2}$
 $=\inf_{g\in G}-\inf_{v\in\mathbb{R}^{n}}\phi(v,g)=-\sup_{g\in G}\varphi(g)$
-where the second equality holds by ˜13. Also, $\phi:\mathbb{R}^{n}\times G\rightarrow\overline{\mathbb{R}}$ and $\varphi:G\rightarrow\mathbb{R}_{-}$ are defined by
+where the second equality holds by Fact˜3.29. Also, $\phi:\mathbb{R}^{n}\times G\rightarrow\overline{\mathbb{R}}$ and $\varphi:G\rightarrow\mathbb{R}_{-}$ are defined by
 $\phi(v,g)=\delta_{B^{n}}(v)-(f^{(k)}(\overline{x})(gv,\ldots,gv))^{2}~~~\text{and}~~~\varphi(g)=\inf_{v\in\mathbb{R}^{n}}\phi(v,g).$
@@
 $L(v,\lambda)=(f^{(k)}(\overline{x})(v,\ldots,v))^{2}-\lambda(|x|^{2k}-1),$
-following Lim [28, Section 4]. Since the constraint is qualified by gradient independence, the first-order optimality condition ensures the existence of $\overline{\lambda}\in\mathbb{R}$ such that
+following Lim [33, Section 4]. Since the constraint is qualified by gradient independence, the first-order optimality condition ensures the existence of $\overline{\lambda}\in\mathbb{R}$ such that
 $\nabla_{v}L(\overline{v},\overline{\lambda})=2kf^{(k)}(\overline{x})(\overline{v},\ldots,\overline{v})\nabla^{k}f(\overline{x})(\overline{v},\ldots,\overline{v})-2k\overline{\lambda}\overline{v}=0,$
@@
 $=\langle u\overline{v},\overline{\lambda}\overline{v}\rangle=\overline{\lambda}\langle u,\overline{v}\hskip 1.13809pt\overline{v}^{T}\rangle.$
-As in the proof of Theorem˜2, we conclude that $C(\overline{v})=0$. ∎
-Corollary 7.
+As in the proof of Theorem˜4.13, we conclude that $C(\overline{v})=0$. ∎
+Corollary 4.17.
 Suppose $f$ is $D^{k}$ near $\overline{x}\in\mathbb{R}^{n}$ with $k\in\mathbb{N}^{*}$, $f^{(i)}(\overline{x})=0$ for all $i\in\llbracket 1,k-1\rrbracket$, and $f^{(k)}(\overline{x})\neq 0$.
@@
 Proof.
-This a consequence of ˜13, Corollary˜2, and Theorem˜3. ∎
+This a consequence of Fact˜3.29, Corollary˜3.19, and Theorem˜4.16. ∎
 4.3 Matrix factorization
 We seek to establish converse results to those developed in the previous section, in the case of matrix factorization, in a desire to characterize flat minima.
-Given $m,n,r\in\mathbb{N}^{*}$ and $M\in\mathbb{R}^{m\times n}$, the map $F:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\ni(X,Y)\mapsto XY-M\in\mathbb{R}^{m\times n}$ is invariant under the action of $\mathrm{GL}(r,\mathbb{R})$ on $\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}$ defined by $(X,Y,A)\mapsto(XA,A^{-1}Y)$.
-As shown in [23, Example 3], for any locally Lipschitz function $g:\mathbb{R}^{m\times n}\to\mathbb{R}$, $g\circ F$ then admits the conserved quantity $C(X,Y)=X^{T}X-YY^{T}$ (see [3, 15, 21, 31] for other derivations).
-We successively treat the cases where $g=\|\cdot\|_{1}$ and $g=\|\cdot\|_{F}$.
-We begin with a simple fact.
-Fact 15.
+Given $m,n,r\in\mathbb{N}^{*}$ and $M\in\mathbb{R}^{m\times n}$, the map
+$\begin{array}[]{cccc}F:&\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}&\longrightarrow&\mathbb{R}\\ &(X,Y)&\longmapsto&XY-M\end{array}$
+is invariant under the action of $\mathrm{GL}(r)$ on $\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}$ defined by
+$(X,Y,A)\mapsto(XA,A^{-1}Y).$
+As shown in [26, Example 5.3], for any locally Lipschitz function $g:\mathbb{R}^{m\times n}\to\mathbb{R}$, $g\circ F$ then admits the conserved quantity
+$C(X,Y)=X^{T}X-YY^{T}$
+(see [3, 17, 24, 36] for other derivations).
+A pair $(X,Y)$ is balanced if $X^{T}X=YY^{T}$ [21, Chapter 6].
+We successively treat the cases where $g=\|\cdot\|_{1}$ and $g=\|\cdot\|_{F}^{2}/2$.
+We begin with a simple fact, which is a special case of [13, Proposition 1].
+Fact 4.18.
 Let $D=\mathrm{diag}(d_{1}I_{k_{1}},\dots,d_{\ell}I_{k_{\ell}})$ with $d_{1}>\cdots>d_{\ell}$ and $k_{1}+\cdots+k_{\ell}=n$, and let $P\in\mathrm{O}(n)$. Then
@@
 ∎
-˜15 implies that for any diagonal matrix $D$ of order $n$ with nonnegative entries and $P\in\mathrm{O}(n)$, one has $DP=PD$ iff $D^{1/2}P=PD^{1/2}$.
-The optimal value in the lemma below is already known [37, Lemma 1], but the local-global property appears to be new.
-Also, the characterization of global minima via the balance condition on $X$ and $Y$ was stated in [12, Lemma 2.2], although the proof seems invalid.
-Lemma 4.
-$2\|M\|_{*}=\min\{\|X\|_{F}^{2}+\|Y\|_{F}^{2}:XY=M\}$ and a feasible point $(X,Y)$ is a global minimum iff it is a local minimum iff $X^{T}X=YY^{T}$.
+Fact˜4.18 implies that for any diagonal matrix $D$ of order $n$ with nonnegative entries and $P\in\mathrm{O}(n)$, one has $DP=PD$ iff $D^{1/2}P=PD^{1/2}$.
+In particular, given $A\succcurlyeq 0$, there is a unique $B\succcurlyeq 0$ such that $A=B^{2}$, called the square root of $A$.
+(Indeed if $B^{2}=\widetilde{B}^{2}$ with $B,\widetilde{B}\succcurlyeq 0$, then given some eigendecompositions $B=PDP^{T}$ and $\widetilde{B}=\widetilde{P}\widetilde{D}\widetilde{P}^{T}$, we have $D=\widetilde{D}$ and $\widetilde{P}^{T}PD^{2}=D^{2}\widetilde{P}^{T}P$.
+Thus $\widetilde{P}^{T}PD=D\widetilde{P}^{T}P$ and $B=\widetilde{B}$.)
+The optimal value in the lemma below is already known [45, Lemma 1], but the local-global property appears to be new.
+Also, the characterization of global minima via the balance condition on $X$ and $Y$ was stated in [14, Lemma 2.2], although the proof seems invalid.
+A similar result constrained to a single orbit was obtained in [21, Theorem 3.7].
+Lemma 4.19.
+If $r\geq\mathrm{rank}\hskip 0.85358pt(M)$, then $2\|M\|_{*}=\min\{\|X\|_{F}^{2}+\|Y\|_{F}^{2}:XY=M\}$ and a feasible point $(X,Y)$ is a global minimum iff it is a local minimum iff $X^{T}X=YY^{T}$.
 Proof.
@@
 $\|\Sigma\|_{*}=\sum_{i=1}^{p}\langle x_{i},y_{i}\rangle\leq\sum_{i=1}^{p}|x_{i}||y_{i}|\leq\sqrt{\sum_{i=1}^{p}|x_{i}|^{2}}\sqrt{\sum_{i=1}^{p}|y_{i}|^{2}}\leq\|X\|_{F}\|Y\|_{F}\leq(\|X\|_{F}^{2}+\|Y\|_{F}^{2})/2.$
-If $XY=M$, the $U^{T}XYV=\Sigma$ and $2\|M\|_{*}=2\|\Sigma\|_{*}\leq\|U^{T}X\|_{F}^{2}+\|YV\|_{F}^{2}=\|X\|_{F}^{2}+\|Y\|_{F}^{2}$, with equality when $(X,Y)=(U\Sigma^{1/2},\Sigma^{1/2}V^{T})$.
-If $(X,Y)$ is a local minimum then $XY=M$ and $X^{T}X=YY^{T}$ by Proposition˜7.
+If $XY=M$, then $U^{T}XYV=\Sigma$ and $2\|M\|_{*}=2\|\Sigma\|_{*}\leq\|U^{T}X\|_{F}^{2}+\|YV\|_{F}^{2}=\|X\|_{F}^{2}+\|Y\|_{F}^{2}$, with equality when $(X,Y)$ equals $(U\Sigma^{1/2},\Sigma^{1/2}V^{T})$ up to some zero rows/columns.
+If $(X,Y)$ is a local minimum then $XY=M$ and $X^{T}X=YY^{T}$ by Proposition˜4.12.
 Conversely, consider some compact singular value decompositions $X=U_{X}\Sigma_{X}V_{X}^{T}$ and $Y=U_{Y}\Sigma_{Y}V_{Y}^{T}$.
-Then $V_{X}\Sigma_{X}^{2}V_{X}^{T}=U_{Y}\Sigma_{Y}^{2}U_{Y}^{T}$, so that $\Sigma_{X}=\Sigma_{Y}$ and $\Sigma_{X}^{2}=(V_{X}^{T}U_{Y})\Sigma_{X}^{2}(V_{X}^{T}U_{Y})^{T}$.
-In the proof of [12, Lemma 2.2], one concludes that $V_{X}^{T}U_{Y}=I_{r}$, but that seems untrue (think of $\Sigma_{X}=I_{r}$).
-In fact, $\Sigma_{X}^{2}(V_{X}^{T}U_{Y})=(V_{X}^{T}U_{Y})\Sigma_{X}^{2}$ and thus $\Sigma_{X}(V_{X}^{T}U_{Y})=(V_{X}^{T}U_{Y})\Sigma_{X}$ by ˜15.
-Using a compact singular value decomposition $M=U\Sigma V^{T}$, this yields $XY=U_{X}\Sigma_{X}V_{X}^{T}U_{Y}\Sigma_{X}V_{Y}^{T}=U_{X}V_{X}^{T}U_{Y}\Sigma_{X}^{2}V_{Y}^{T}=U\Sigma V^{T}$, and so $\Sigma_{X}=\Sigma_{Y}=\sqrt{\Sigma}$.
+By Lemma˜4.20 below, $\Sigma_{X}=\Sigma_{Y}=\sqrt{\Sigma}$.
 Thus $\|X\|_{F}^{2}+\|Y\|_{F}^{2}=2\|M\|_{*}$ and $(X,Y)$ is globally optimal.
 ∎
-Using Lemma˜4, we can show that the converse of Theorem˜2 holds in $\ell_{1}$-matrix factorization.
-Proposition 8.
+The following result can also be deduced from [40, Lemma 3.1].
+Lemma 4.20.
+If $XY=M$ and $X^{T}X=YY^{T}$, then $\Sigma_{X}=\Sigma_{Y}=\sqrt{\Sigma}$ given some compact singular value decompositions $X=U_{X}\Sigma_{X}V_{X}^{T}$ and $Y=U_{Y}\Sigma_{Y}V_{Y}^{T}$
+Proof.
+Plugging in the decompositions yields $V_{X}\Sigma_{X}^{2}V_{X}^{T}=U_{Y}\Sigma_{Y}^{2}U_{Y}^{T}$, so that $\Sigma_{X}=\Sigma_{Y}$ and $\Sigma_{X}^{2}=(V_{X}^{T}U_{Y})\Sigma_{X}^{2}(V_{X}^{T}U_{Y})^{T}$.
+In the proof of [14, Lemma 2.2], one concludes that $V_{X}^{T}U_{Y}=I_{r}$, but that seems untrue (think of $\Sigma_{X}=I_{r}$).
+In fact, $\Sigma_{X}^{2}(V_{X}^{T}U_{Y})=(V_{X}^{T}U_{Y})\Sigma_{X}^{2}$ and thus $\Sigma_{X}(V_{X}^{T}U_{Y})=(V_{X}^{T}U_{Y})\Sigma_{X}$ by Fact˜4.18.
+Using a compact singular value decomposition $M=U\Sigma V^{T}$, this yields $XY=U_{X}\Sigma_{X}V_{X}^{T}U_{Y}\Sigma_{X}V_{Y}^{T}=U_{X}V_{X}^{T}U_{Y}\Sigma_{X}^{2}V_{Y}^{T}=U\Sigma V^{T}$, and so $\Sigma_{X}=\Sigma_{Y}=\sqrt{\Sigma}$.
+∎
+Using Lemma˜4.19, we can show that the converse of Theorem˜4.13 holds in $\ell_{1}$-matrix factorization.
+Proposition 4.21.
 Given $M\in\mathbb{R}^{m\times n}$, let $f:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\to\mathbb{R}$ be defined by $f(X,Y)=\|XY-M\|_{1}$.
@@
 (i) $\Longrightarrow$ (ii) is obvious.
-(ii) $\Longrightarrow$ (iii) is due to Theorem˜2.
+(ii) $\Longrightarrow$ (iii) is due to Theorem˜4.13.
 (iii) $\Longrightarrow$ (i) Observe that $f(X,Y)=\|F(X,Y)\|_{1}$ where F(X,Y)=XYMF(X,Y)=XY-M.
 We thus compute $F(X,Y)(H,K)=XK+HY$ and $F^{\prime}(X,Y)^{*}(\Lambda)=(X^{T}\Lambda,\Lambda Y^{T})$.
-By ˜10,
+By Fact˜3.25,
 $\partial f(X,Y)=\left\{\begin{pmatrix}\Lambda Y^{T}\\ X^{T}\Lambda\end{pmatrix}:\Lambda\in\mathrm{sign}(XY-M)\right\}.$
@@
 Hence $\overline{H}\hskip 1.13809pt\overline{K}=\overline{\Lambda}\hskip 1.13809pt\overline{Y}^{T}\overline{X}^{T}\overline{\Lambda}=\overline{\Lambda}\hskip 1.13809ptM^{T}\overline{\Lambda}$.
-Since $\overline{H}^{T}\overline{H}=\overline{K}\hskip 1.13809pt\overline{K}^{T}$, by Lemma˜4,
+Since $\overline{H}^{T}\overline{H}=\overline{K}\hskip 1.13809pt\overline{K}^{T}$, by Lemma˜4.19,
 $(\overline{H},\overline{K})\in\arg\min\{\|H\|_{F}^{2}+\|K\|_{F}^{2}:HK=\overline{\Lambda}\hskip 1.13809ptM^{T}\overline{\Lambda}\}.$
@@
 To see why, it is useful to recall when equality holds in $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ where $(X,Y)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}$.
-Given a matrix $A\in\mathbb{R}^{m\times n}$ with singular value decomposition $A=U\Sigma V^{T}$, let $S_{A}=\{v\in\mathbb{R}^{n}:|Av|=\|A\|_{2}|v|\}$ and $d_{A}=\dim S_{A}$, i.e., the multiplicity of the maximal singular value of $A$.
-Let $\widehat{U}$ (resp.
-$\widehat{V}$) denote the first $d_{A}$ columns of $U$ (resp.
-$V$), in other words, the maximal left (resp.
+Given a real-valued matrix $A$ with a singular value decomposition $A=U_{A}\Sigma_{A}V_{A}^{T}$, let $S_{A}=\{v\in\mathbb{R}^{n}:|Av|=\|A\|_{2}|v|\}$ and $d_{A}=\dim S_{A}$, i.e., the multiplicity of the maximal singular value of $A$.
+Let $\widehat{U}_{A}$ (resp.
+$\widehat{V}_{A}$) denote the first $d_{A}$ columns of $U_{A}$ (resp.
+$V_{A}$), in other words, the maximal left (resp.
 right) singular vectors of $A$.
-Fact 16.
+Fact 4.22.
 $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}\Longleftrightarrow S_{X}\cap YS_{Y}\neq\{0\}\Longleftrightarrow\mathrm{Im}\hskip 0.85358pt\widehat{V}_{X}\cap\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}\neq\{0\}$.
@@
 For all $v\in\mathbb{R}^{n}$, $|XYv|\leq\|XY\|_{2}|v|\leq\|X\|_{2}\|Y\|_{2}|v|$ and $|XYv|\leq\|X\|_{2}|Yv|\leq\|X\|_{2}\|Y\|_{2}|v|$.
-Thus $S_{XY}\supset S_{X}\cap YS_{Y}$ and, if $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$, then $S_{XY}\subset S_{X}\cap YS_{Y}$.
+Thus $S_{XY}\supseteq S_{X}\cap YS_{Y}$ and, if $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$, then $S_{XY}\subseteq S_{X}\cap YS_{Y}$.
 Lastly, $S_{X}=\mathrm{Im}\hskip 0.85358pt\widehat{V}_{X}$ and $YS_{Y}=\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}$.
@@
 The optimal value in the next result is known in the multilayer case [38], a fortiori in the two layer case. On the other hand, the local-global property again appears to be new.
-Lemma 5.
-$2\|M\|_{2}=\min\{\|X\|_{2}^{2}+\|Y\|_{2}^{2}:XY=M\}$ and a feasible point $(X,Y)$ is a global minimum iff it is a local minimum iff $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ and $\|X\|_{2}=\|Y\|_{2}$ iff $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}$.
+Lemma 4.23.
+If $r\geq\mathrm{rank}\hskip 0.85358pt(M)$, then $2\|M\|_{2}=\min\{\|X\|_{2}^{2}+\|Y\|_{2}^{2}:XY=M\}$ and a feasible point $(X,Y)$ is a global minimum iff it is a local minimum iff $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ and $\|X\|_{2}=\|Y\|_{2}$ iff $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}$.
 Proof.
@@
 with equality exactly when $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ and $\|X\|_{2}=\|Y\|_{2}$.
-This happens in particular when $(X,Y)=(U\Sigma^{1/2},\Sigma^{1/2}V^{T})$.
+This happens in particular when $(X,Y)$ equals to $(U\Sigma^{1/2},\Sigma^{1/2}V^{T})$ up to some zero rows/columns.
 If $\|X\|_{2}\neq\|Y\|_{2}$, say $\|X\|_{2}<\|Y\|_{2}$, then $\|tX\|_{2}^{2}+\|t^{-1}Y\|_{2}^{2}<\|X\|_{2}^{2}+\|Y\|_{2}^{2}$ for all $t\in(1,\|Y\|_{2}^{1/2}\|X\|_{2}^{-1/2}]$.
@@
 $<\|X\|_{2}^{2}+\|Y\|_{2}^{2}$
-by ˜16. ∎
+by Fact˜4.22. ∎
+Lemma˜4.19, Lemma˜4.20, and Lemma˜4.23 immediately imply the following.
+Corollary 4.24.
+$\arg\min\{\|X\|_{F}^{2}+\|Y\|_{F}^{2}:XY=M\}\subseteq\arg\min\{\|X\|_{2}^{2}+\|Y\|_{2}^{2}:XY=M\}$.
+The inclusion can be strict in Corollary˜4.24, as we will see in Example˜4.29.
 In contrast to the multilayer case [38], it is possible to determine the maximal eigenvalue and eigenspace of the Hessian at any global minimum.
-Proposition 9.
+In fact, shortly after submitting this work on arXiv, an expression for the maximal eigenvalue in the multilayer case was proposed using Kronecker products [30, Theorem 5], implying the expression below [30, Corollary 6].
+Proposition 4.25.
 Given $M\in\mathbb{R}^{m\times n}$, let $f:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\to\mathbb{R}$ be defined by $f(X,Y)=\|XY-M\|_{F}^{2}/2$. Suppose $XY=M$. Then
@@
 Proof.
-Observe that $f(X,Y)=\|F(X,Y)\|_{F}^{2}/2$ where F(X,Y)=XYMF(X,Y)=XY-M. By ˜11, $\lambda_{1}(\nabla f(X,Y))=\|F^{\prime}(X,Y)\|_{2}^{2}$ so we compute $F^{\prime}(X,Y)(H,K)=HY+XK$. One has
+Observe that $f(X,Y)=\|F(X,Y)\|_{F}^{2}/2$ where F(X,Y)=XYMF(X,Y)=XY-M. By Fact˜3.26, $\lambda_{1}(\nabla f(X,Y))=\|F^{\prime}(X,Y)\|_{2}^{2}$ so we compute $F^{\prime}(X,Y)(H,K)=HY+XK$. One has
 $\|HY+XK\|_{F}$ $\leq\|HY\|_{F}+\|XK\|_{F}\leq\|H\|_{F}\|Y\|_{2}+\|X\|_{2}\|K\|_{F}$
@@
 $U_{Y}$).
-If $(H,K)\in E_{1}(\nabla^{2}((X,Y))$ and $\mathrm{Im}\hskip 0.85358ptH^{T}\cap\mathrm{Im}\hskip 0.85358ptK\neq\{0\}$, then $\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}A^{T}\cap\mathrm{Im}\hskip 0.85358pt\widehat{V}_{Y}B\neq\{0\}$, $\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}\cap\mathrm{Im}\hskip 0.85358pt\widehat{V}_{Y}\neq\{0\}$, and $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ by ˜16.
+If $(H,K)\in E_{1}(\nabla^{2}((X,Y))$ and $\mathrm{Im}\hskip 0.85358ptH^{T}\cap\mathrm{Im}\hskip 0.85358ptK\neq\{0\}$, then $\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}A^{T}\cap\mathrm{Im}\hskip 0.85358pt\widehat{V}_{Y}B\neq\{0\}$, $\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}\cap\mathrm{Im}\hskip 0.85358pt\widehat{V}_{Y}\neq\{0\}$, and $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ by Fact˜4.22.
 ∎
@@
 Recall that $\mathrm{Im}\hskip 0.85358ptAA^{T}=\mathrm{Im}\hskip 0.85358ptA$ for any matrix $A\in\mathbb{R}^{m\times n}$.
-Proposition 10.
+Proposition 4.26.
 Given $M\in\mathbb{R}^{m\times n}$, let $f:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\to\mathbb{R}$ be defined by $f(X,Y)=\|XY-M\|_{F}^{2}/2$.
@@
   6. (vi)
+    $\|\overline{X}\|_{2}^{2}+\|\overline{Y}\|_{2}^{2}=2\|M\|_{2}$;
+  7. (vii)
     $\|\overline{X}\|_{2}=\|\overline{Y}\|_{2}=\sqrt{\|M\|_{2}}$.
 Proof.
-(i) $\Longrightarrow$ (ii) follows by Definition˜3.
-(ii) $\Longrightarrow$ (iii) is due to Corollary˜2.
+(i) $\Longrightarrow$ (ii) follows by Definition˜3.2.
+(ii) $\Longrightarrow$ (iii) is due to Corollary˜3.19.
 (iii) $\Longrightarrow$ (iv) If $\|\overline{X}\|_{2}\neq\|\overline{Y}\|_{2}$, say $\|\overline{X}\|_{2}<\|\overline{Y}\|_{2}$, then $\|t\overline{X}\|_{2}^{2}+\|t^{-1}\overline{Y}\|_{2}^{2}<\|\overline{X}\|_{2}^{2}+\|\overline{Y}\|_{2}^{2}$ for all $t\in(1,\|\overline{Y}\|_{2}^{1/2}\|\overline{X}\|_{2}^{-1/2}]$.
 If $(\overline{X},\overline{Y})=(0,0)$, then take $\overline{H}^{T}$ and $\overline{K}$ to be the rectangular identities.
-Otherwise, $\nabla^{2}f(\overline{X},\overline{Y})\neq 0$ so Theorem˜3 applies.
-(iv) $\Longrightarrow$ (v) By Lemma˜5, $\|\overline{X}\|_{2}=\|\overline{Y}\|_{2}$.
-Since $\overline{H}^{T}\overline{H}=\overline{K}\hskip 1.13809pt\overline{K}^{T}\neq 0$, $\mathrm{Im}\hskip 0.85358pt\overline{H}^{T}=\mathrm{Im}\hskip 0.85358pt\overline{H}^{T}\overline{H}=\mathrm{Im}\hskip 0.85358pt\overline{K}\hskip 1.13809pt\overline{K}^{T}=\mathrm{Im}\hskip 0.85358pt\overline{K}\neq\{0\}$ and so $\|\overline{X}\hskip 1.13809pt\overline{Y}\|_{2}=\|\overline{X}\|_{2}\|\overline{Y}\|_{2}$ by Proposition˜9.
+Otherwise, $\nabla^{2}f(\overline{X},\overline{Y})\neq 0$ so Theorem˜4.16 applies.
+(iv) $\Longrightarrow$ (v) By Lemma˜4.23, $\|\overline{X}\|_{2}=\|\overline{Y}\|_{2}$.
+Since $\overline{H}^{T}\overline{H}=\overline{K}\hskip 1.13809pt\overline{K}^{T}\neq 0$, $\mathrm{Im}\hskip 0.85358pt\overline{H}^{T}=\mathrm{Im}\hskip 0.85358pt\overline{H}^{T}\overline{H}=\mathrm{Im}\hskip 0.85358pt\overline{K}\hskip 1.13809pt\overline{K}^{T}=\mathrm{Im}\hskip 0.85358pt\overline{K}\neq\{0\}$ and so $\|\overline{X}\hskip 1.13809pt\overline{Y}\|_{2}=\|\overline{X}\|_{2}\|\overline{Y}\|_{2}$ by Proposition˜4.25.
 (v) $\Longrightarrow$ (i) For any $XY=M$ such that $\|X\|_{2}=\|Y\|_{2}$, we have
@@
 $=(\|X\|_{2}^{2}+\|Y\|_{2}^{2})r^{2}/2+\sqrt{2}(\|X\|_{2}+\|Y\|_{2})r^{3}/4+r^{4}/8$
-using the same direction to obtain the equalities as in first part of the proof of Proposition˜9.
+using the same direction to obtain the equalities as in first part of the proof of Proposition˜4.25.
 Let $XY=M$.
-Since $f$ is $D^{2}$ and $\nabla f(X,Y)=0$, by Corollary˜2, $\accentset{\circ}{f}(X,Y,r)=\lambda_{1}(\nabla^{2}f(X,Y))r^{2}/2+o(r^{2})$.
+Since $f$ is $D^{2}$ and $\nabla f(X,Y)=0$, by Corollary˜3.19, $\accentset{\circ}{f}(X,Y,r)=\lambda_{1}(\nabla^{2}f(X,Y))r^{2}/2+o(r^{2})$.
 Hence, if $\lambda_{1}(\nabla^{2}f(\overline{X},\overline{Y}))<\lambda_{1}(\nabla^{2}f(X,Y))$, then there exists $\overline{r}>0$ such that $\accentset{\circ}{f}(\overline{X},\overline{Y},r)<\accentset{\circ}{f}(X,Y,r)$ for all $r\in(0,\overline{r}]$.
-Otherwise, by Lemma˜5, $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}=\|\overline{X}\|_{2}=\|\overline{Y}\|_{2}$ and so $\accentset{\circ}{f}(\overline{X},\overline{Y},r)=\accentset{\circ}{f}(X,Y,r)$ for all $r\geq 0$.
+Otherwise, by Lemma˜4.23, $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}=\|\overline{X}\|_{2}=\|\overline{Y}\|_{2}$ and so $\accentset{\circ}{f}(\overline{X},\overline{Y},r)=\accentset{\circ}{f}(X,Y,r)$ for all $r\geq 0$.
 In either case, $(\overline{X},\overline{Y})\preceq(X,Y)$, hence $(\overline{X},\overline{Y})$ is globally flat.
-(v) $\Longleftrightarrow$ (vi) This was already shown in Lemma˜5. ∎
-Proposition˜10 implies that when $(X,Y)$ is a flat global minimum, there must exist a maximal eigenvector $(H,K)$ of the Hessian such that $H^{T}H=KK^{T}$.
+(v) $\Longleftrightarrow$ (vi) $\Longleftrightarrow$ (vii) This was already shown in Lemma˜4.23. ∎
+One immediately obtains the following corollary, as announced in the introduction.
+Corollary 4.27.
+Given $M\in\mathbb{R}^{m\times n}$, let $f:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\to\mathbb{R}$ be defined by $f(X,Y)=\|XY-M\|_{F}^{2}/2$.
+If $XY=M$ and $X^{T}X=YY^{T}$, then $(X,Y)$ is flat.
+Proof.
+While the converse of Corollary˜4.27 is false, as noted in [38], Proposition˜4.26 implies that when $(X,Y)$ is a flat global minimum, there must exist a maximal eigenvector $(H,K)$ of the Hessian such that $H^{T}H=KK^{T}$.
 This appears to be new.
-On the other hand, it is possible that $X^{T}X\neq YY^{T}$, as noted in [33].
-We give an example using a well-known fact.
-Fact 17.
+We give a counterexample to the converse of Corollary˜4.27 using a well-known fact.
+Fact 4.28.
 Given $M\in\mathbb{R}^{m\times n}$, consider a compact singular value decomposition $M=U\Sigma V^{T}$ and let $r=\mathrm{rank}\hskip 0.85358ptM$.
 For all $(X,Y)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}$,
-$XY=M~~~\Longleftrightarrow~~~\exists A\in\mathrm{GL}(r,\mathbb{R}):(X,Y)=(U\Sigma^{1/2}A,A^{-1}\Sigma^{1/2}V^{T}).$
+$XY=M~~~\Longleftrightarrow~~~\exists A\in\mathrm{GL}(r):(X,Y)=(U\Sigma^{1/2}A,A^{-1}\Sigma^{1/2}V^{T}).$
 Proof.
-$XY=M$ iff $\Sigma^{-1/2}U^{T}XYV\Sigma^{-1/2}=I_{r}$ iff $\Sigma^{-1/2}U^{T}X=A$ and $YV\Sigma^{-1/2}=A^{-1}$ for some $A\in\mathrm{GL}(r,\mathbb{R})$ iff $X=U\Sigma^{1/2}A$ and $Y=A^{-1}\Sigma^{1/2}V^{T}$.
-∎
-Example 5.
+First consider the diagonal case
+$M=\begin{pmatrix}\Sigma&0\\ 0&0\end{pmatrix},~~~X=\begin{pmatrix}X_{1}\\ X_{2}\end{pmatrix},~~~\text{and}~~~Y=\begin{pmatrix}Y_{1}&Y_{2}\end{pmatrix}$.
+Then $XY=M$ iff $X_{1}Y_{1}=\Sigma$, $X_{1}Y_{2}=0$, $X_{2}Y_{1}=0$, and $X_{2}Y_{2}=0$.
+But $X_{1}Y_{1}=\Sigma$ iff $\Sigma^{-1/2}X_{1}Y_{1}\Sigma^{-1/2}=I_{r}$ iff there is $A\in\mathrm{GL}(r)$ such that $\Sigma^{-1/2}X_{1}=A$ and $Y_{1}\Sigma^{-1/2}=A^{-1}$ iff $X_{1}=\Sigma^{1/2}A$ and $Y_{1}=A^{-1}\Sigma^{1/2}$.
+Now $X_{1}Y_{2}=0$ implies $Y_{2}=0$ and $X_{2}Y_{1}=0$ implies $X_{2}=0$.
+Thus $XY=M$ iff $X_{1}=\Sigma^{1/2}A$ and $Y_{1}=A^{-1}\Sigma^{1/2}$ and $(X_{2},Y_{2})=(0,0)$.
+Now consider the general case
+$M=\begin{pmatrix}U&\widetilde{U}\end{pmatrix}\begin{pmatrix}\Sigma&0\\ 0&0\end{pmatrix}\begin{pmatrix}V&\widetilde{V}\end{pmatrix}^{T}.$
+Then $XY=M$ iff
+$\begin{pmatrix}U&\widetilde{U}\end{pmatrix}^{T}XY\begin{pmatrix}V&\widetilde{V}\end{pmatrix}=\begin{pmatrix}\Sigma&0\\ 0&0\end{pmatrix}$
+iff
+$X=\begin{pmatrix}U&\widetilde{U}\end{pmatrix}\begin{pmatrix}\Sigma^{1/2}A\\ 0\end{pmatrix}=U\Sigma^{1/2}A~~~\text{and}~~~Y=\begin{pmatrix}A^{-1}\Sigma^{1/2}&0\end{pmatrix}\begin{pmatrix}V&\widetilde{V}\end{pmatrix}^{T}=A^{-1}\Sigma^{1/2}V^{T}.\qed$
+Example 4.29.
 Given $a>b>0$, the flat minima of $\mathbb{R}^{2\times 2}\times\mathbb{R}^{2\times 2}\ni(X,Y)\mapsto\|XY-M\|_{F}^{2}\in\mathbb{R}$ with $M=\mathrm{diag}(a,b)$ are of the form
@@
 Proof.
-By Proposition˜10 and ˜17, $(X,Y)$ is a flat minimum iff $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}$ and $(X,Y)=(M^{1/2}A,A^{-1}M^{1/2})$ for some $A\in\mathrm{GL}(2,\mathbb{R})$.
+By Proposition˜4.26 and Fact˜4.28, $(X,Y)$ is a flat minimum iff $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}$ and $(X,Y)=(M^{1/2}A,A^{-1}M^{1/2})$ for some $A\in\mathrm{GL}(2)$.
 Consider a decomposition $A=DNQ$ where $D=\mathrm{diag}(\alpha,\beta)$ with $\alpha,\beta>0$, $N$ is upper triangular of order 2 with ones on the diagonal with $N_{12}=\gamma\in\mathbb{R}$, and $Q\in\mathrm{O}(2)$.
@@
 $\left\|\begin{pmatrix}1&u\\ 0&v\end{pmatrix}\right\|_{2}=\max_{x^{2}+y^{2}=1}\sqrt{(x+uy)^{2}+v^{2}y^{2}}\geq\max_{x^{2}+y^{2}=1}|x+uy|=\sqrt{1+u^{2}}.\qed$
+Corollary˜4.27 means that one can compute a flat global minimum easily using linear algebra: just take a compact singular value decomposition $M=U\Sigma V^{T}$ and let $(X,Y)$ equal to $(U\Sigma^{1/2},\Sigma^{1/2}V^{T})$ up to some zero rows/columns.
+Below, we consider two complementary dynamical systems converging to balanced solutions, which are therefore flat.
+The first concerns gradient trajectories of the objective function.
+Proposition 4.30.
+Given $M\in\mathbb{R}^{m\times n}$ and $r\geq\mathrm{rank}\hskip 0.85358pt(M)$, let $f:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\to\mathbb{R}$ be defined by $f(X,Y)=\|XY-M\|_{F}^{2}/2$.
+There exist differentiable functions $(X,Y):\mathbb{R}\to\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}$ such that
+$(\dot{X},\dot{Y})=-\nabla f(X,Y)$
+and
+$(0,0)\xleftarrow[-\infty\leftarrow t]{}(X(t),Y(t))\xrightarrow[t\to\infty]{}(X_{\infty},Y_{\infty})$
+where $(X_{\infty},Y_{\infty})$ is a balanced global minimum.
+Proof.
+Without loss of generality, we may restrict ourselves to the diagonal case
+$M=\begin{pmatrix}\Sigma&0\\ 0&0\end{pmatrix}$.
+Let $(X_{0},Y_{0})=(\mathrm{diag}(\Sigma,0),\mathrm{diag}(\Sigma,0))/2\in\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}$ and consider the initial value problem
+$\left\{\begin{array}[]{ccc}(\dot{X},\dot{Y})&=&-\nabla f(X,Y),\\ (X(0),Y(0))&=&(X_{0},Y_{0}).\end{array}\right$.
+It yields $\mathrm{rank}\hskip 0.85358pt(M)$ decoupled systems
+$\forall i\in\llbracket 1,\mathrm{rank}\hskip 0.85358pt(M)\rrbracket,~~~\left\{\begin{array}[]{ccc}(\dot{X}_{ii},\dot{Y}_{ii})&=&-\nabla\varphi_{i}(X_{ii},Y_{ii}),\\ (X_{ii}(0),Y_{ii}(0))&=&(\Sigma_{ii}^{1/2},\Sigma_{ii}^{1/2})/2,\end{array}\right$.
+where $\varphi_{i}(x,y)=(xy-\Sigma_{ii})^{2}/2$ and all other entries are fixed to zero throughout time.
+This reduces the problem to the scalar case, i.e., $f(x,y)=(xy-1)^{2}$, and to its dynamics on the diagonal $x=y$, i.e., on the function $g(x)=(x^{2}-1)^{2}$.
+We deduce that
+$(0,0)\xleftarrow[-\infty\leftarrow t]{}(X_{ii}(t),Y_{ii}(t))\xrightarrow[t\to\infty]{}(\Sigma_{ii}^{1/2},\Sigma_{ii}^{1/2})$
+and conclude by Corollary˜4.27. ∎
+The second concerns gradient trajectories of the conserved quantity $\|C\|_{F}^{2}$. We recall a simple fact.
+Fact 4.31.
+Let $\phi:\mathbb{R}^{m\times n}\to\mathbb{R}$ be defined by $\phi(A)=\|A\|_{2}$.
+Given a singular value decomposition $A=U\Sigma V^{T}$ where $u_{i}$ and $v_{i}$ resp.
+denote the columns of $U$ and $V$, $\partial\phi(A)=\mathrm{co}\hskip 0.85358pt\{u_{i}v_{i}^{T}:i\in\llbracket 1,d_{A}\rrbracket\}$.
+Proof.
+We have $\partial\phi(A)=U\partial\phi(\Sigma)V^{T}$ and $\phi(\Sigma+tH)=\sup\{|(\Sigma+tH)w|:w\in S^{n-1}\}=\|\Sigma\|_{2}+t\max\{|H_{ii}|:i\in\llbracket 1,d_{A}\rrbracket\}=\phi(\Sigma)+t\sigma_{\{E_{ii}\}_{i\in\llbracket 1,d_{A}\rrbracket}}(H)$ where $E_{ij}$ denotes the canonical basis of $\mathbb{R}^{m\times n}$.
+By the max formula [43, Theorem 23.4], $\phi^{\prime}(\Sigma,H)=\sigma_{\partial\phi(\Sigma)}(H)=\sigma_{\{E_{ii}\}_{i\in\llbracket 1,d_{A}\rrbracket}}(H)$ and so $\partial\phi(\Sigma)=\mathrm{co}\hskip 0.85358pt\{E_{ii}:i\in\llbracket 1,d_{A}\rrbracket\}$.
+∎
+Proposition 4.32.
+Given $M\in\mathbb{R}^{m\times n}$, let $f,c,\varphi:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\to\mathbb{R}$ be defined by
+$f(X,Y)$ $=\|XY-M\|_{F}^{2}/2,$
+$c(X,Y)$ $=\|X^{T}X-YY^{T}\|_{F}^{2}/4,$
+$\varphi(X,Y)$ $=\|X\|_{2}^{2}-2\|XY\|_{2}+\|Y\|_{2}^{2}.$
+The following hold:
+  1. (i)
+    $f+c$ is coercive;
+  2. (ii)
+[…]
+  3. (iii)
+[…]
+  4. (iv)
+    For any $X_{0}Y_{0}=M$,
+    $\left\{\begin{array}[]{ccc}(\dot{X},\dot{Y})&=&-\nabla c(X,Y),\\ (X(0),Y(0))&=&(X_{0},Y_{0}),\end{array}\right$.
+    has a unique flattening global solution converging to a balanced global minimum of $f$ such that
+    $\frac{d}{dt}\lambda_{1}(\nabla^{2}f(X,Y))\leq-(\lambda_{1}(\nabla^{2}f(X,Y))-2\|M\|_{2})^{2},$
+    $\forall t>0,~~~\lambda_{1}(\nabla^{2}f(X(t),Y(t)))\leq 2\|M\|_{2}+1/t.$
+Proof.
+(i) Following the derivation in [24, Example 1],
+$\|X\|_{2}^{4}+\|Y\|_{2}^{4}$ $\leq\|X^{T}X\|_{F}^{2}+\|YY^{T}\|_{F}^{2}$
+$=\|X^{T}X-YY^{T}\|_{F}^{2}+2\langle X^{T}X,YY^{T}\rangle$
+$=\|X^{T}X-YY^{T}\|_{F}^{2}+2\|XY\|_{F}^{2}$
+$\leq\|X^{T}X-YY^{T}\|_{F}^{2}+2(\|XY-M\|_{F}+\|M\|_{F})^{2}$
+$=2c(X,Y)+2\left(\sqrt{2f(X,Y)}+\|M\|_{F}\right)^{2}$
+$\leq 2(f+c)(X,Y)+2\left(\sqrt{2(f+c)(X,Y)}+\|M\|_{F}\right)^{2}.$
+(ii) For any point $(X,Y)$, we have
+$\nabla c(X,Y)=\begin{pmatrix}X(X^{T}X-YY^{T})\\ -(X^{T}X-YY^{T})Y\end{pmatrix}$. (5)
+If $XY=M$ and $\nabla c(X,Y)=0$, then $XX^{T}X=XYY^{T}=MY^{T}$ and $YY^{T}Y=X^{T}XY=X^{T}M$. Thus $X^{T}XX^{T}X=X^{T}MY^{T}=Y^{T}YY^{T}Y$ and $X^{T}X=YY^{T}$ by Fact˜4.18.
+(iii) By Fact˜4.31, $\partial\varphi(X,Y)$ is the convex hull of the points of the form $2(s_{x}u_{x}v_{x}^{T},s_{y}u_{y}v_{y}^{T})$ where $s_{x}=\|X\|_{2}$ and $s_{y}=\|Y\|_{2}$ and $u_{x},v_{x},u_{y},v_{y}$ are some singular vectors of $X$ and $Y$ resp.
+The max formula [43, Theorem 23.4] states that $\varphi^{\prime}(X,Y;D)=\max\{\langle H,D\rangle:H\in\partial\varphi(X,Y)\}$.
+We thus compute
+$~\langle(s_{x}u_{x}v_{x}^{T},s_{y}u_{y}v_{y}^{T}),\nabla c(X,Y)\rangle$
+$=$ $~\langle s_{x}u_{x}v_{x}^{T},X(X^{T}X-YY^{T})\rangle-\langle(X^{T}X-YY^{T})Y,s_{y}u_{y}v_{y}^{T}\rangle$
+$=$ $~s_{x}u_{x}^{T}X(X^{T}X-YY^{T})v_{x}-s_{y}u_{y}^{T}(X^{T}X-YY^{T})Yv_{y}$
+$=$ $~s_{x}^{2}v_{x}^{T}(X^{T}X-YY^{T})v_{x}-s_{y}^{2}u_{y}^{T}(X^{T}X-YY^{T})u_{y}$
+$=$ $~s_{x}^{4}-s_{x}^{2}|Y^{T}v_{x}|^{2}-s_{y}^{2}|Xu_{y}|^{2}+s_{y}^{4}$
+$=$ $~s_{x}^{4}-|Y^{T}X^{T}u_{x}|^{2}-|XYv_{y}|^{2}+s_{y}^{4}$
+$\geq$ $~s_{x}^{4}-2s_{xy}^{2}+s_{y}^{4}$
+$\geq$ $~(s_{x}^{2}-2s_{xy}+s_{y}^{2})^{2}/2$
+where the last inequality is due to $s_{xy}=\|XY\|_{2}\leq\|X\|_{2}\|Y\|_{2}=s_{x}s_{y}$. Indeed,
+$2s_{x}^{4}-4s_{xy}^{2}+2s_{y}^{4}-(s_{x}^{2}-2s_{xy}+s_{y}^{2})^{2}$ $=s_{x}^{4}+s_{y}^{4}-8s_{xy}^{2}+4s_{x}^{2}s_{xy}+4s_{xy}s_{y}^{2}-2s_{x}^{2}s_{y}^{2}$
+$=(s_{x}^{2}-s_{y}^{2})^{2}+4s_{xy}(s_{x}^{2}+s_{y}^{2}-2s_{xy})$
+$\geq(s_{x}^{2}-s_{y}^{2})^{2}+4s_{xy}(s_{x}^{2}+s_{y}^{2}-2s_{x}s_{y})$
+$=(s_{x}^{2}-s_{y}^{2})^{2}+4s_{xy}(s_{x}-s_{y})^{2}.$
+(iv) Since $XY$ and $f$ are conserved along trajectories of $\nabla c$ by Proposition˜4.4, $f+c$ decreases along them.
+By (i), they are bounded and hence globally defined according to [24, Proposition 2].
+They must converge to balanced points by virtue of (ii).
+Let $\mu=\varphi(X,Y)=\lambda_{1}(\nabla^{2}f(X,Y))-2\|M\|_{2}$, where the second equality is due to Proposition˜4.25.
+Since $\varphi$ agrees with a convex function along any fixed trajectory, by [10, Proposition 2] the chain rule holds almost surely:
+$\dot{\mu}=\sup_{H\in\partial\varphi(X,Y)}\langle H,(\dot{X},\dot{Y})\rangle=\sup_{H\in\partial\varphi(X,Y)}\langle H,-\nabla c(X,Y)\rangle=\varphi^{\prime}(X,Y;H)\leq-\mu^{2}$
+where the last inequality is due to (iii). Integrating $\dot{\mu}\mu^{-2}\geq 1$ yields $1/\mu(t)\geq 1/\mu(t)-1/\mu(0)\geq t$ if $\mu(0)\neq 0$, otherwise $\mu(t)=0$ for all $t\geq 0$. ∎
+Proposition˜4.32 complements the analysis of Riemannian gradient dynamics in [21, Theorem 5.1].
+Given $X_{0}Y_{0}=M$ with $\mathrm{rank}\hskip 0.85358pt(X_{0})=\mathrm{rank}\hskip 0.85358pt(Y_{0})=r$, it considers
+$\left\{\begin{array}[]{ccc}(\dot{X},\dot{Y})&=&-\mathrm{grad}\hskip 0.85358pt\Phi,\\ (X(0),Y(0))&=&(X_{0},Y_{0}),\end{array}\right$.
+where $\Phi:\mathrm{GL}(r)(X_{0},Y_{0})\to\mathbb{R}$ is defined by $\Phi(X,Y)=\|X\|_{F}^{2}+\|Y\|_{F}^{2}$, and the orbit is equipped with the normal Riemannian metric [21, Equation (5.3) p.
+186].
+[21, Theorem 5.1 (a)] shows that $\mathrm{grad}\hskip 0.85358pt\Phi=(X(X^{T}X-YY^{T}),-(X^{T}X-YY^{T})Y)$, but no connection is made with the conserved quantity.
+According to (5), we in fact have $\mathrm{grad}\hskip 0.85358pt\Phi=\nabla c$, which considerably simplifies the analysis and does not require the rank condition.
+If that condition holds, then each trajectory of $\nabla c$ remains in the orbit of the initial point.
 5 Examples
 Some examples are in order.
-Example 6.
+Example 5.1.
 A point $x$ is a flat minimum of $f(x)=(x_{2}\mathrm{ReLU}(x_{1})+x_{3}-1)^{2}$ iff $(x_{1}<0\land x_{3}=1)\lor(x_{1}=0\land|x_{2}|\leq 1\land x_{3}=1)$.
 Proof.
-Observe that $f$ fails to be regular at $(0,x_{2},x_{3})$ when $x_{2}(x_{3}-1)<0$, so the calculus rule in Corollary˜1 does not apply.
-Regardless, by Remark˜1, $\mathrm{lip}f(x)=0$ whenever $f(x)=0$, so it wouldn’t be of much use.
+Observe that $f$ fails to be regular at $(0,x_{2},x_{3})$ when $x_{2}(x_{3}-1)<0$, so the calculus rule in Corollary˜3.18 does not apply.
+Regardless, by Remark˜3.27, $\mathrm{lip}f(x)=0$ whenever $f(x)=0$, so it wouldn’t be of much use.
 We simply resort to the definition.
@@
 $\nabla^{2}f(x)=2\begin{pmatrix}x_{2}^{2}&x_{2}x_{1}&x_{2}\\ x_{2}x_{1}&x_{1}^{2}&x_{1}\\ x_{2}&x_{1}&1\end{pmatrix}=\begin{pmatrix}x_{2}\\ x_{1}\\ 1\end{pmatrix}\begin{pmatrix}x_{2}\\ x_{1}\\ 1\end{pmatrix}^{T}$
-and so $\lambda_{1}(\nabla^{2}f(x))=x_{2}^{2}+x_{1}^{2}+1$. By Proposition˜4, $\accentset{\circ}{f}(x,r)=(x_{2}^{2}+x_{1}^{2}+1)r^{2}/2+o(r^{2})>r^{2}$. ∎
-Example 7.
+and so $\lambda_{1}(\nabla^{2}f(x))=x_{2}^{2}+x_{1}^{2}+1$. By Proposition˜3.17, $\accentset{\circ}{f}(x,r)=(x_{2}^{2}+x_{1}^{2}+1)r^{2}/2+o(r^{2})>r^{2}$. ∎
+Example 5.2.
 The flat minima of $f(x)=(x_{1}x_{2}-1)^{4}$ are $\pm(1,1)$.
 Proof.
-By ˜11, $\|f^{(4)}(x)\|=4|F^{\prime}(x)|_{2}^{4}$ when $F(x)=x_{1}x_{2}-1=0$, i.e., $\|f^{(4)}(x)\|=4(x_{1}^{2}+x_{2}^{2})^{2}$.
-This quantity is strictly minimized at $\pm(1,1)$ over the solution set, so one concludes by Corollary˜2.
-Alternatively, one notices that the flat minima are the same as that of $|x_{1}x_{2}-1|$ by using Definition˜3 and the fact that $t\in\mathbb{R}_{+}\mapsto t^{4}$ is strictly increasing.
-Since the new function is regular and locally Lipschitz, by Corollary˜1 it suffices to compute its Lipschitz modulus when $x_{1}x_{2}=1$, i.e., $x_{1}^{2}+x_{2}^{2}$.
+By Fact˜3.26, $\|f^{(4)}(x)\|=4|F^{\prime}(x)|_{2}^{4}$ when $F(x)=x_{1}x_{2}-1=0$, i.e., $\|f^{(4)}(x)\|=4(x_{1}^{2}+x_{2}^{2})^{2}$.
+This quantity is strictly minimized at $\pm(1,1)$ over the solution set, so one concludes by Corollary˜3.19.
+Alternatively, one notices that the flat minima are the same as that of $|x_{1}x_{2}-1|$ by using Definition˜3.2 and the fact that $t\in\mathbb{R}_{+}\mapsto t^{4}$ is strictly increasing.
+Since the new function is regular and locally Lipschitz, by Corollary˜3.18 it suffices to compute its Lipschitz modulus when $x_{1}x_{2}=1$, i.e., $x_{1}^{2}+x_{2}^{2}$.
 ∎
-Example 8.
+Example 5.3.
 The origin is the sole flat minimum of $f(x)=x_{2}^{2}+x_{1}^{2}x_{2}^{4}$.
@@
 Thus $\lambda_{1}(\nabla^{2}f(x_{1},0))=2$, as claimed in the introduction.
-By virtue of Examples 2 and 3, it would seem futile to compute higher-order derivatives.
+By virtue of Examples 3.20 and 3.21, it would seem futile to compute higher-order derivatives.
 But in this special case, the maximal eigenvectors of $f^{(2)}(x_{1},0)$ and $f^{(4)}(x_{1},0)$ align (with $(0,1)$) while $f^{(3)}(x_{1},0)=0$.
-Since $f^{(4)}(x_{1},0)=24x_{1}^{2}$, we deduce that $\accentset{\circ}{f}(x_{1},0,r)=r^{2}+x_{1}^{2}r^{4}+o(r^{4})$ and conclude by Definition˜3.
-Example 9.
+Since $f^{(4)}(x_{1},0)=24x_{1}^{2}$, we deduce that $\accentset{\circ}{f}(x_{1},0,r)=r^{2}+x_{1}^{2}r^{4}+o(r^{4})$ and conclude by Definition˜3.2.
+Example 5.4.
 A global minimum of $f(x)=(a_{1}x_{1}^{2}+\cdots+a_{n}x_{n}^{2}-1)^{2}$ where $a\in\mathbb{R}^{n}$ is flat iff $a_{i}x_{i}=0$ for all $i\notin I=\arg\min\{a_{i}:a_{i}>0\}$.
@@
 The objective $f$ is invariant under the natural action of $G=\mathrm{diag}(O(m),I_{n-m})$ once we reorder the indices so that those in $I$ come first.
-By ˜11, $\lambda_{1}(\nabla^{2}f(x))=2|F^{\prime}(x)|_{2}^{2}$ when $F(x)=a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}-1=0$.
+By Fact˜3.26, $\lambda_{1}(\nabla^{2}f(x))=2|F^{\prime}(x)|_{2}^{2}$ when $F(x)=a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}-1=0$.
 Since $F^{\prime}(x)=2(a_{1}x_{1},\dots,a_{n}x_{n})$, we seek to minimize $|F^{\prime}(x)|^{2}=4(a_{1}^{2}x_{1}^{2}+\dots+a_{n}^{2}x_{n}^{2})$ subject to $a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}=1$.
-Consider the Lagrangian $L(x,\lambda)=\sum_{i}a_{i}^{2}x_{i}^{2}+\lambda(1-\sum_{i}a_{i}x_{i}^{2})$.
+Consider the Lagrangian
+$L(x,\lambda)=\sum_{i=1}^{n}a_{i}^{2}x_{i}^{2}+\lambda\left(1-\sum_{i=1}^{n}a_{i}x_{i}^{2}\right).$
 Since the constraint is qualified, at optimality $a_{i}^{2}x_{i}-\lambda a_{i}x_{i}=0$, namely, $a_{i}x_{i}(a_{i}-\lambda)=0$.
@@
 Thus, for any global minimum $\overline{x}$ of $f$ such that $\overline{x}_{i}=0$ for all $i\notin I$, $G\overline{x}$ is a strict global minimum of $\lambda_{1}(\nabla^{2}f(x))+\delta_{[f=f(\overline{x})]}$.
-It follows that every point in $G\overline{x}$ is flat by Corollary˜4.
-No other global minimum is flat by Corollary˜2.
+It follows that every point in $G\overline{x}$ is flat by Corollary˜3.24.
+No other global minimum is flat by Corollary˜3.19.
 ∎
-Example 10.
+Example 5.5.
 The flat global minima of $f(x)=|x_{1}x_{3}-a|+|x_{2}x_{3}-b|$ are
@@
 Proof.
-By ˜10, $f$ is regular,
+By Fact˜3.25, $f$ is regular,
 $\partial f(x)=\left\{\begin{pmatrix}\lambda_{1}x_{3}\\ \lambda_{2}x_{3}\\ \lambda_{1}x_{1}+\lambda_{2}x_{2}\end{pmatrix}~,\hskip 8.53581pt\lambda\in\begin{pmatrix}\mathrm{sign}(x_{1}x_{3}-a)\\ \mathrm{sign}(x_{2}x_{3}-b)\end{pmatrix}\right\}$,
-and $\mathrm{lip}f(x)=2|x_{3}|^{2}+(|x_{1}|+|x_{2}|)^{2}$.
-Thus $(0,0,0)$ is the sole flat global minimum when $(a,b)=(0,0)$ by Corollary˜1.
+and $\mathrm{lip}f(x)=\sqrt{2|x_{3}|^{2}+(|x_{1}|+|x_{2}|)^{2}}$.
+Thus $(0,0,0)$ is the sole flat global minimum when $(a,b)=(0,0)$ by Corollary˜3.18.
 When $(a,b)\neq(0,0)$, $\mathrm{argmin}f=\{(at,bt,1/t):t\neq 0\}$.
@@
 $|t|=\sqrt{\frac{\sqrt{2}}{|a|+|b|}}.$
-This yields the expression of the flat global minima by Corollary˜1.
-We now verify the claim of Proposition˜8.
+This yields the expression of the flat global minima by Corollary˜3.18.
+We now verify the claim of Proposition˜4.21.
 The objective $f$ is invariant under the natural action of the Lie group $G=\{\mathrm{diag}(t,t,1/t):t\neq 0\}$ whose Lie algebra is $\mathfrak{g}=\mathrm{span}\{\mathrm{diag}(1,1,-1)\}$.
-By Proposition˜6, a conserved quantity is given by $C(x)=x_{1}^{2}+x_{2}^{2}-x_{3}^{2}$.
+By Proposition˜4.11, a conserved quantity is given by $C(x)=x_{1}^{2}+x_{2}^{2}-x_{3}^{2}$.
 When $ab\neq 0$, the quantity
@@
 $\sqrt{\frac{\sqrt{2}}{|a|+|b|}}=\frac{1}{\sqrt[4]{a^{2}+b^{2}}}~\Longleftrightarrow~2(a^{2}+b^{2})=(|a|+|b|)^{2}~\Longleftrightarrow~a^{2}+b^{2}=2|ab|~\Longleftrightarrow~|a|=|b|.$
-Example 11.
+Example 5.6.
 The flat global minima of $f(x)=(x^{\upsilon}-1)^{2}$ where $x^{\upsilon}=x_{1}^{\upsilon_{1}}\cdots x_{n}^{\upsilon_{n}},\upsilon\in{\mathbb{N}^{*}}^{n},$ are
@@
 $\mathfrak{g}=\mathrm{span}\{\mathrm{diag}(\upsilon_{n},0,\ldots,0,-\upsilon_{1}),\ldots,\mathrm{diag}(0,\ldots,0,\upsilon_{n},-\upsilon_{n-1})\}.$
-By Proposition 6, a conserved quantity is given by
+By Proposition 4.11, a conserved quantity is given by
 $C(x)=(\upsilon_{n}x_{1}^{2}-\upsilon_{1}x_{n}^{2},\ldots,\upsilon_{n}x_{n-1}^{2}-\upsilon_{n-1}x_{n}^{2}).$
@@
 $F^{\prime}(x)=(\upsilon_{1}x^{\upsilon}/x_{1},\ldots,\upsilon_{n}x^{\upsilon}/x_{n}).$
-By ˜11, $\lambda_{1}(\nabla^{2}f(x))=2|F^{\prime}(x)|^{2}$ when $F(x)=0$.
+By Fact˜3.26, $\lambda_{1}(\nabla^{2}f(x))=2|F^{\prime}(x)|^{2}$ when $F(x)=0$.
 Given $t_{1},\ldots,t_{n-1}>0$, let $x_{t}=\mathrm{diag}(t_{1}^{\upsilon_{n}},\ldots,t_{n-1}^{\upsilon_{n}},t^{-\overline{\upsilon}})\overline{x}$ where $\overline{x}$ is global minimum such that $|\overline{x}_{i}|=1$ for all $i$.
@@
 for some constant $\omega>0$.
-Together with Proposition˜6, this means that ˜2 holds.
-Applying Corollary˜5, $\lambda_{1}(\nabla^{2}f)$ strictly decreases along the trajectories of $\nabla c$.
-Since $\lambda_{1}(\nabla^{2}f)+\delta_{[f=0]}$ is coercive, they must be bounded and hence globally defined by [21, Proposition 2].
+Together with Proposition˜4.11, this means that Assumption˜4.3 holds.
+Applying Corollary˜4.8, $\lambda_{1}(\nabla^{2}f)$ strictly decreases along the trajectories of $\nabla c$.
+Since $\lambda_{1}(\nabla^{2}f)+\delta_{[f=0]}$ is coercive, they must be bounded and hence globally defined by [24, Proposition 2].
 Since $c$ is semi-algebraic, they must also converge to a point $\overline{x}$ where $\nabla c(\overline{x})=0$ [35], i.e., a flat global minimum.
@@
 Acknowledgments
-I would like to thank Wenqing Ouyang for his careful reading of the manuscript and valuable feedback.
+I would like to thank Wenqing Ouyang for his careful reading of the manuscript and valuable feedback. I also wish to thank Xiaopeng Li for useful criticism.